A Complete Bibliography of Publications in the Journal of Statistical Planning and Inference:

1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.01

Title word cross-reference

(+1, −1) [CKM87]. (2n, n, λ, 1) [EFJ82]. (7, b, r, 3, λ|16) [FH88]. (n, k, s, 2) [MLF81]. (θ_{1} − dθ_{2})θ_{3} [LGB89]. (v, k, t) [Hwa86]. 2 [KS83, Kur88]. 2(l + 1) [SO85]. 21 × 21 [CKM87]. 2^{m} [Gho81, GK85, Kuw82, NS86, OS85, SO85]. 2_{1}^{m} [Kuw86]. 2_{1}^{m} + m_{2} [KK88, Kuw88, Kuw87]. 2^{m} × 3^{n} [BI84]. 2^{n} [SFM84]. 2l + 1 [KK88]. 3 [Mik89, SS88b]. 3^{k} [KM89]. 3^{m} [Kuw81]. 4 [VD89]. 4 ≤ m_{1>} + m_{2} ≤ 6 [Kuw88]. 4q [Daw85]. $68.25 [Mie82]. 8 [Rok84]. ^{∗} [Bjø86, GY85]. ^{−} [Kuw87]. ^{2} [Hab81, HG86]. _{1} [DK89, Wu89]. _{2} [SD85a]. _{i} [RMW85]. A [Kag88b, Kuw88, MR89, Mas88]. b ≥ t [CM86a]. BIB(9, 18t, 8t, 4, 3t) [KM88]. C [CK84a, GW80, MB85]. χ^{2} [GCG88]. D [Gaf82, Jac84b, Lau88, Man89, MK82, PR88, CKM87, KY88, KC83, KF84]. E [BI84, Jac86, LJ87]. F [Bar82, BS82, FHM84, Hay89, MLF81, SFM84, dL87]. f^{2}(x) [AH84]. F_{9} [Mie87]. k [Chm81, GM85, JS81a, Mac83]. k > υ [JM89]. L [DK89, EL86, HR88, Sho89, Wu89]. L_{1} [Mie87]. L_{2} [Mie87]. L_{∞} [Mie87]. L_{p} [CM86c]. M [HJ81, JS81b, Loh85]. m_{2} [Kuw86]. m ≤ 6 [NS86]. μ [Sin81]. MV [Jac86, LJ87]. n [DN88, MLF81, ZdK81]. n□3 mod 4 [KF84]. N ≡ 1 mod 9 [KM89]. N ≡ 2 mod 16 [CCM89]. N ≡ 3 mod 4 [Mas88]. n ≡ (mod 4) [KC83]. ν_{μ} + 1 + ϵ, ν_{μ}; t, q [HD89]. OL(s, t) [MLF81]. P [Aic83a, Bjø86, Pra85]. P (b_{i} ≤ R_{i} ≤ a_{i}, 1 ≤ i ≤ m) [HM82]. p^{n} [AT80]. PG(4, q) [Cas89]. PG(5, q) [Cas89]. π [Ram82]. q [Daw85, Lau88]. q + 1 [Kes86]. R [HJ81, MB85]. s [Lon84]. s^{m} [KN88]. S_{e}PMR_{i} [RMW85]. T [SD85a, GM85, HK80, KH83, Kur88, PC81, SMK88]. t + 2 [Kur88]. τ [WB80]. U [AGLS84, Bab89, JSV87]. V [But85, CCM89, Kuw82, Kuw86, Kuw88, NS86, Gho80, GK85, Kuw81]. V_{p,q} [KN88]. X [Hab81, HG86]. z [SYT81].

-associate [VD89]. -ball [Lau88]. -concurrence [SS88b]. -condition [Bjø86]. -coverings [KS83]. -density [Sin81]. -designs [EFJ82, KH83, KM89, CK84a]. -designs-part [HK80]. -efficiency [Kag88b]. -estimates [Loh85]. -estimators [HJ81, JS81b]. -linear [CM86c]. -matrices [CKM87, GW80]. -matrix [Hay89]. -min.hypers [HD89]. -NN [Mac83]. -optimal [CKM87, KY88, KC83, KF84, Roy82, BI84, Gaf82, Kuw88, Lau88, LJ87, Man89, MR89, Mas88, MK82, PR88]. -optimality [JS80, Jac84b, Jac86]. -rectangle [FHM84]. -sample [JS81a]. -squares [MLF81, SFM84, dL87]. -statistics [AGLS84, Bab89, EL86, HR88, JSV87, Sho89]. -symbol [Kur88]. -test [PC81, Bar82, BS82]. -tests [GM85]. -th [Pra85]. -transformation [SYT81]. -tuples [Rok84]. -values [Aic83a]. -way [Mik89].

1-10 [Ano84f]. 1-104 [Ano82e]. 1-106 [Ano81c]. 1-108 [Ano82f]. 1-114 [Ano80e]. 1-128 [Ano83f]. 1-130 [Ano84g, Ano88i]. 1-132 [Ano89h]. 1-134 [Ano85g, Ano88j]. 1-136 [Ano88k, Ano89i]. 1-138 [Ano89j]. 1-146 [Ano84h]. 1-154 [Ano86g]. 1-406 [Ano85h]. 1-408 [Ano86h]. 1-412 [Ano87h]. 105-204 [Ano82g]. 107-212 [Ano81d]. 109-192 [Ano82h]. 11 [Oma85a]. 11-20 [Ano88h]. 115-212 [Ano80f]. 129-243 [Ano83g]. 131-238 [Ano88l]. 131-264 [Ano84i]. 133-266 [Ano89k]. 135-266 [Ano85i, Ano88m]. 137-268 [Ano88n]. 137-270 [Ano89l]. 139-276 [Ano89m]. 147-268 [Ano84j]. 155-414 [Ano86i]. 16 [Sch89a]. 193-294 [Ano83h]. 1982 [Bha85]. 1983 [Ano82d, Bha85]. 1987/1989 [Ano86c]. 1988 [Flu89a, Flu89b].

205-308 [Ano82i]. 213-308 [Ano81e]. 213-326 [Ano80g]. 239-472 [Ano88o]. 24061 [Roh88]. 245-396 [Ano83i]. 265-415 [Ano84k]. 267-406 [Ano88p, Ano89n]. 267-412 [Ano85j]. 269-399 [Ano84l]. 269-402 [Ano88q]. 271-406 [Ano89o]. 277-412 [Ano89p]. 295-421 [Ano83j].

309-420 [Ano81f]. 309-424 [Ano82j]. 327-404 [Ano80h].

7 [CH83b].

A-optimal [Wie88b]. ABLUE [Che80]. absolute [iY84]. abstract [Cue88]. accelerated [BE87, MP88, PSSZ87, SS83b]. accidental [SW85]. Acknowledgement [Ano85a, Ano88a, Pur86, Pur89, Sri81, Sri85a]. Acknowledgment [Ano87a]. ad [SO85]. ad-optimal [SO85]. adaptation [SS82b]. adaptive [BN86, BM86b, DV89, HHL84, Sch87b, WR87]. Addendum [Kag86]. additional [CG83, SSP87]. additive [Wie88b, Wie89]. adjacent [Nag88]. admissibility [BLT89, CA85, Sob86]. Admissible [BM88, CLS87, VM83, BM86a, BPS87, GKZ85]. Advanced [Flu89a]. affine [KS83]. after [BA82, Tay89]. against [BT87, Gho82, LaR86, PS87, Sen84b, Whi89]. age [Hjo89]. aging [De 87]. aids [Doc88, GH89]. Algebra [EM85]. algorithm [Dyk81]. algorithms [Fed89, Tuz82, Flu89a]. aligned [aS89]. Allocating [Bar82]. Allocation [Day85, Aic83b, CW83, Rob85, SW85]. allocatory [McC86]. allowed [Bil85]. allowing [De 87]. alternative [BS85]. alternatives [GG86b, HM82, LaR86, Mie86, PS87, PS84b, Sen84b, SP89]. among [RFS86]. analyses [Mie86]. Analysis [CHT84, SNS85, BZ87, BdBP82, Bha84a, Bou81, CC89, CB88, De 87, DK88, GS87, Gup83, Gup85a, Joh81, LKM80, Pre87, SM86, SS86b, Sin82b, SS85b, Ven81, Ylv81, Zab88, Zid88]. analytic [Gil86b]. ancillary [Bha89]. Anil [Flu89a]. Announcement [Ano80a, Ano83a, Ano87b, Sri84, Sri89, Sri86, Sri87a]. anonymity [GH89]. anonymously [FK80]. anova [SS82b, MH84, Wie89]. antibodies [GH89]. Antipodally [Pre82]. application [Bak89, GH89, Sha86, Wie89, Zel88, Zid88]. Applications [Bel84, BJ88, BFPR84, CR82, Cha86, Csá86, DN86b, DKPS88, Gra89, Gup84, Gup87, HD89, Jac89, KK89, Lin89, Nie86, PR88, RM87, WB80, Flu89b]. applied [CL89, Ano80c]. approach [GL89, Jos88, dAV80]. approaches [Doc88, SS85b]. Approximate [AWW87a, RM88, Rob85, Fun82, Puk83]. approximating [HG85]. approximation [AWW87b, Ahm83, PW85, RJDT83]. approximations [DKPS88, HG86]. April [Ano85j]. arbitrary [MP88]. arc [Goo87c]. arc-sine [Goo87c]. argument [Jos89, Law88]. arguments [Hay82a]. arising [Ton81]. ARMA [HG85]. array [MLF81]. Arrays [Lon84, Gil86a, Kag88a, Kur88, SMK88, SO85, SN83]. artificial [DK88]. aspects [PR85, SSP87, Sin82b, SS88b, Tan87]. assays [Gup89, NB85, PG88]. assessing [DeG88, WS88]. assessment [Gas82, RM88]. associate [Kag82, Kag86, Moh81, VD89]. associated [Cas89, Mil88]. association [Gup88, LM88]. assortment [Sid82a]. asymmetric [CM86b]. asymmetrical [GN85, Lew82]. Asymptotic [Amo86, BD88, Bjø86, Cha89b, EH84, GK80, HR88, JT85, JSV87, LP88a, LP88b, Mes87, Now89, OV85, Sha89, aS89, CMS82, GV88, Hab81, Hay82b, HJ81, Kha81, Kla89, LPY86, Lo89, MB85, Pra85, Tah89, Tak86]. Asymptotically [Hor89, MR85, GG86b, Hor88, Sch87a, Sch89a]. attaining [Kag88a]. August [Ano82d, Ano84i]. Author [Ano80b, Ano81a, Ano82a, Ano83b, Ano83c, Ano84b, Ano84a, Ano85b, Ano85c, Ano86a, Ano86b, Ano87c, Ano87d, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c]. auto [Cha89b]. auto- [Cha89b]. autocorrelated [KW89]. Automating [Law88]. automorphisms [SD85a]. autoregressive [Gas88]. Auxiliary [San84, Day85, SS85a]. availability [BK83]. average [BM89b, Wan87].

Back [Hor85]. Back-to-back [Hor85]. bacterial [PK83]. Bahadur [AK88, BS82, CS88]. balance [CH82, JWM83, Jac86, Joh81, Sch87b, SS85c]. Balanced [BI84, Gil86a, Kes86, Kuw86, KK88, SN83, Whi89, Agr87, BDK80, BP88, Cal86, CS89, DD89, Don87, Gho80, GK85, IF88, KT80, Kha82, Kur88, Kuw81, Kuw87, KN88, Kuw88, MR89, NS86, RFS86, Roy88, SMK88, SO85, SS88b]. ball [Lau88]. ballot [Wat86]. BAN [BB82]. bandwidth [MSV89]. Barlow [May87]. Bartels [Ylv81]. based [Ang82, Bau89, BB82, Bus81, Chi86, EL86, JSV87, JS81b, KS83, MB85, NS87a, Nag88, SAU83, aS89, SG84, Tik80]. basic [Gup83]. Bayes [CS84a, DV89, GH83, GY85, GL89, Lia89, Sti82, ZB81, dAV80]. Bayes-P [GY85]. Bayes-type [CS84a]. Bayesian [AWW87a, BZ87, BdBP82, CL89, DeG88, Kha88, PSSZ87, Sin87, Tan85]. be [Nab86]. behavior [Tuz82]. Bernoulli [ZB81]. best [BK81, Bau89, GY85, GL89, Muk87]. beta [Wie88a]. better [BE88, Muk84, Ris85]. between [AC89, Bar82, CK86, GLP84, Gil86b, Hud85, KPY83, Kos89, Lin88, Maj86, Yan88]. Bhaskar [LS84, Seb84, Seb85a]. Bhattacharya [Kha84]. bias [Fun82, Not89, SW85]. bib [Vec84, FH88, HP85, HLT89]. BIBD [Gho82]. binary [GW80, KY88, Yan88, Yeh88]. binomial [BJ88, Gou83, GL89, Lia89, Ris85, Sac82]. bioassay [JJ83]. biomedical [RM88]. Birkhäuser [Bha85]. bivariate [AC89, BE87, SS83c]. Blacksburg [Roh88]. Block [BGH84, Agr87, BDK80, BT85, BT87, BP88, Bak89, BK86c, Cas89, Cat84, CM86a, CC84, DD89, Don87, Fei85, Gaf82, Gno87, Gup83, Gup85b, Gup86, Gup89, Hed84, Jac84a, Jac85, Jac88, JM89, KT80, Kai89, KKL88, Kes86, LJ87, MR89, Mas88, NB85, Puk83, PG88, RFS86, Roe87, Roy82, Sas88, Seb84, Seb85a, aS89, Sin82b, SS85b, SB86, Stu88, Ton81, WO82, Whi89, Yeh88]. blocks [CH83a, CH83b, KM88, RFS86, WO82]. blood [GH89]. blue [Kai89, MB83]. Board [Ano80d, Ano81b, Ano82b, Ano82c, Ano83d, Ano84c, Ano84d, Ano85e, Ano85f, Ano86d, Ano86e, Ano87f, Ano87g, Ano88e, Ano88f, Ano88g, Ano89e, Ano89f, Ano89g]. Bongers [Sid82a]. Book [Ano80c, Bha85, Flu89a, Flu89b, Mie82, Roh88, Sid82a, Wat82, Ylv81]. bootstrapped [AGLS84]. Boston [Bha85, Ylv81]. bound [Kha80, Sim80]. bounded [CR82, Fed89, GSvR88, Seb82]. Bounds [SMK88, BK83, KT80, Kag88a, Kha84, May87, Sah80, Stu88, Wu89]. Bradley [CS84a]. Bradley-Terry [CS84a]. Breach [Ano80c]. broad [SP89]. broken [GR87]. Brown [Bha87]. budget [Ris85].

C [Flu89a, Flu89b, Sid82a]. C3 [Doc88]. calibration [Oma85a, Oma85b, SD85b]. Call [Ano87e]. Campus [Ano82d]. capture [Zel88]. capture-recapture [Zel88]. care [LKM80]. case [BM89a, Gra89, Mór83, Muk84, Tuz82]. case-control [Gra89]. Catalan [DN86a, DN86b]. categorical [BdBP82, Bou81, CHT84]. categorisation [RB85]. Cauchy [Loh85]. causes [RM88]. cells [Khu89]. censored [AWW87a, Bel89b, Che80, GW88, LPY86, Tik80, WB80, Wu87]. censoring [Ara87, Ara89, BN86, CS80, GV89, MB87, Sen84a, Sri87b]. censorship [Che89, GSV85, MP88, PR89]. center [Hal86]. centered [BJV89]. Central [SP89, Wu89]. Certain [Bos86, Car82, DL84b, EW89, Pfa88, Sin82a, iY84]. changepoint [CH87c, WS84]. Chao [Geo87]. Chaology [Goo89b]. Chaotics [Goo89b]. Characterisation [Dur89]. Characteristic [Kuw81, Day85, KN88, Pou87]. Characteristics [RFS86, ACR89, BS85]. Characterization [Gov80, Bak89, CRS87, HD89, Hed84, SNS85, SG84]. Characterizations [BM86a, Lin88, Lin89, NS87a, Nag88]. Charlier [MB86]. Chebyshev [Hal86]. chemical [JWM83, Sch87b]. chi [HG86, MB85, Moo84]. chi-square [MB85]. chi-squared [HG86, Moo84]. choice [MSV89]. choices [FK80]. choosing [Bau89]. Chopra [GK85, NS86]. Chung [Bha85]. cinlar [Bha85]. class [Agr87, AE85, BS89, Cal86, CS80, Fel89, GH80, GN85, HM82, Hed84, Moh81, Oma83, PS87, Pei88, RS84, Sas88, SP89, SR82, yS89a, Wie85, Yeh88]. classes [Sin82a, Zah85]. classical [BM86b, Csá86, SD85b, SB86]. classification [CD83]. Cliffs [Flu89a]. climatology [Mie82]. cluster [PK83]. clustering [Flu89a]. codes [HD89, LS81, Tam81]. coding [FW87]. coefficient [BA82, GG86c, SAM84]. coefficients [BP87, Bor82, CG83, KG89]. coins [Tos89]. collocated [BEH84]. column [AE85, BS89, EM85, IF88, SNS85]. combinations [BM84, BI86]. Combinatorial [Sta89, BFPR84]. combinatorics [Men86]. combined [CMS82, Fei85]. Combining [BD83, BM86b, Zid88]. comment [BP87, Gho89b]. common [CS84b, Nai82]. Company [Mie82]. comparative [Goi88, SG86]. compared [HG86]. comparing [DeG88, GM85, HM88, Jac88, JM89, Stu88]. Comparison [CF82, GK83a, SD85b, TW86, Bab88, Bea83, CB88, CS84a, EH84, Hec87, HPT84, JT85, PC81]. comparisons [CP80, Maj86, Not85, SS82b]. competing [RMW85, RM88]. Competitions [Ano86c]. complements [Mon86]. Complete [CH83a, CH83b, GH80, SFM84, Sin82a, WO82]. complex [CP80]. components [BM84, CD83, Das89, Flu89b, GK83b, GKZ85, Gno87, Pei88]. Compstat’88 [Ano87e]. computability [Weg88]. computable [Oma83]. Computer [Flu89a, Hjo89]. concepts [Bha84b]. concerning [FK80, Kuw86, May87]. concomitant [CC89]. concurrence [RG81, SS88b]. condition [BP88, Bjø86, Nab86, iY84]. conditional [Bab88, Nag86a]. conditionality [Bha84b]. Conditioning [Bha89]. Conditions [BT87, Har88, JS80, Jac85, Jac89, MM87, Rou89]. conference [Ano82d]. Confidence [LGB88, LGB89, Chi80, CS88, GSV85, GSvR88, GV89, KSY82, LB89, Mes87]. confounded [Gho89a]. conjecture [May87]. connected [BT85, Kag88b]. Connectedness [Kag82, Kag86, Moh81]. connection [CK86]. consistencies [Pra85]. Consistency [Cli89, Lia87, Pfa88, Bjø86, CM86c, Jur80, LP86, Sid82b]. Consistent [Whi87]. constants [PS84a]. Constrained [Lee88]. constraint [Ris85]. constraints [Kur88, SMK88, Wie85]. constructed [yS89b]. constructing [BDK80, BK86c, CS89, CK84a]. Construction [Agr87, CCM89, Gho89a, Muk81, Roy82, Roy88, WK87, AE85, Ash81, CH82, Daw85, Ecc80, Kag85, KF84, Law88, LJ87, MLF81, PS88, Pay81, Sch87a, Sch89a, Seb80, Tam81, Kag89]. constructions [BT85, CC84, Str84]. consuming [Sri87b]. contaminacy [Vís86]. contamination [Vís86]. contiguous [HRS88]. contingency [Gil86b, MB85]. continuation [Sko88, Sko89]. continuous [Bar86]. contrast [Cha88a, Cha89a]. Control [GW88, Cox81, DR82, Gra89, GM85, HM88, HPT84, JM89, Not85, SW85, Stu88]. Controlled [BM89a, BM89b]. Controlling [FW87]. Convergence [EW89, Lia89, Bos86, CR82, DK89, Geo87, Har88, KP87b, Mac83, Ruy89, Sin81, Wu87]. coordinates [BD83]. Cornelis [Ylv81]. correcting [HD89]. Correction [Sch89a]. correlated [Cam84, PC81]. correlation [BA82, GG86c, KG89, TL87]. correspondence [Tho84]. corresponding [RMW85]. Corrigendum [Oma85a]. cost [Tay89]. counter [Goo87c]. counter-intuitive [Goo87c]. counterfeit [Tos89]. counting [Mau85]. countries [Ano86c]. coupling [DKPS88]. covariance [CC89, CP80, Hay86, Khu89, Pre87, Ven81]. covariate [Hec87, Tro82a, Tro82b, Wie85]. covariates [Nac89]. covering [Muk84, RS84]. coverings [KS83, Vec84]. Cox [Sen84a, Sen84b]. Cramér [Kha80, Mór83, Sim80]. Cramér-Rao [Kha80, Mór83, Sim80]. Criteria [BP88, BB82, BM86b, GK83a, Pay81]. criterion [Cha88b, Moh81, Pre87, SM82]. crossings [SS83a]. cubic [CK86, Mik89, Not81]. curve [BD88, Che89, JJ83, Kub89, SS86b]. Cybernetics [Ano80c]. cycles [SD86]. Cyclic [Tro82a, Agr87, CC84, DL80, SFM84]. cylinders [FM88].

D [KM89, Mik89]. D-optimality [KM89]. D-optimum [Mik89]. damaged [GR82]. Data [DR82, AWW87a, BT87, BdBP82, Bel89b, BJV89, Bou81, Cam84, CHT84, Cli89, EH83, GW88, GLP84, Gho82, Gil86b, KY88, MSV89, PS87, SS88a, Sin85, WB80, Wu87, Flu89a, Ylv81]. data-driven [MSV89]. Data-swapping [DR82]. dealing [Doc88]. death [RMW85]. December [Ano82h, Ano83i]. decision [DN88, Doc88, GH80, Zid88]. decrement [Gar83]. Dedication [Ano85d]. defective [HX87]. defectives [AS85, WR87]. degrees [Fra80]. demand [LKM80]. Deming [BZ87]. Density [iY84, Fed89, Gup82, HG85, Kha80, KW88, Mac83, MPS84, MSV89, OV85, PK83, Pra85, Ruy89, Sin81]. dependence [Rou89]. dependent [Che89]. derivative [OV85, Pra85]. derivatives [MPS84, Sin81]. derived [SO85]. deriving [Fun82]. Design [CB88, Jai81, CL89, Cha88a, Cha89a, EK81, EM85, Fed89, Gho89b, Hec87, Hed84, Kai89, MB83, Mik89, Moo88, MK82, Nac89, Páz89, Puk83, Roe87, Roy82, Sär82, SW85, SM85, Tho84, WO82, Wie88b]. Designing [San81]. Designs [SD85a, SS85c, Yan88, AP83, Agr87, AC89, AT80, AE85, Ash81, BS89, BGH84, BDK80, BT85, BT87, BP88, Bak89, BK86c, BP80, BI84, Bou81, BI86, Bus81, BFPR84, But85, Cal86, CS89, Cas89, Cat84, CK84a, CM86a, CCM89, CKM87, CH87a, CM86b, CC84, CH82, CH83a, CH83b, CLS87, DD89, DL80, Don87, Ecc80, EH84, EFJ82, FHM84, Fel89, FH88, Gaf82, GK83a, Gho81, GK85, GA85, Gho89a, Gho89b, GW80, GPW87, Gno87, Gup83, GN85, Gup85a, Gup85b, GR85, Gup86, Gup89, HD89, HK80, Hed84, HP85, HM88, HLT89, HS88, Hud83, Iac85, IF88, JWM83, Jac84a, Jac84b, Jac85, Jac86, Jac88, JM89, Jac89, Joh81, KT80, KT81, Kag82, KH83, Kag85, Kag86, Kag88b, Kag89, KKL88, Kes86, KY88]. designs [Kha82, KM88, KM89, KC83, KF84, Kra85, Kuw81, Kuw82, Kuw86, Kuw87, KK88, KN88, Kuw88, LS84, Lau88, LJ87, Lee88, LN82, Lon81, Lon86, Maj86, Man89, MR89, Mas88, MB89, MMR84, Moh81, Moo88, Mor84, Muk81, MR85, MB88, NB85, NS86, Not81, Not82, Not85, Not89, OS85, PS88, PRJ83, PR85, PR88, Pes84, PP87, Puk80, Puk83, PG88, RFS86, Ram82, RS84, Roy88, SS87, Sas88, SP81, SS86a, Sch87b, Seb84, Seb85a, SM87, SO85, SD82, Sin82b, SN83, SS83d, SS85b, SS88b, Spr85, Spr87, SAM84, SB86, Sri87b, Stu88, yS89b, Ton81, Tro82a, Tro82b, Vec84, WR87, Whi89, Wie85, WK87, Yeh88, dL87]. detecting [BK89, ZB81]. detection [Bel84]. determinant [CKM87, PR85]. determination [Gup82, Mar87, TL87, Sid82a]. Determining [AS85, RM87]. determinism [Weg88]. developing [Ano86c]. deviation [Wu87]. diagnostic [Pay81]. Diagnostics [Car88]. dichotomous [TW86]. dictionary [Rot86]. difference [AC89, Hud85, Kos89, LGB88, Mie87, dL87]. differences [AA86, Bos86, yS89b]. different [Fra80]. dimension [Sta89]. dimensional [SB86, YMO87]. dimensions [Hud83]. directed [Bil85, NS85, NS87b]. disclosure [Cox81, DR82]. discontinuous [JS81b]. discounted [BM89a]. discrete [GG81, Hor85, KS82, Nag82, Nag86a, Nag86b, Nag88, Zel88]. discriminant [Amo86, CG83]. discrimination [McC86, Sch84, Yan88, YMO87]. discussed [BZ87]. Discussion [Ano89d, And88, Ano86j, Ano89q, Gho88, Roh88]. disease [GH89]. dispersion [SG84]. disproportionate [MH84]. distinguishing [RFS86]. distributed [Mus87]. Distribution [RG81, AK87, BK86b, BM87, Ber80, CRS87, Ebr86, GSvR88, GG86b, Gov80, HSB86, Kap88, LP88b, MB86, OZ84, Rot86, SS83a, Shi89, Wan87, Wat83, Wie88a, Zah85]. Distribution-free [RG81, GG86b]. distributional [RM87]. Distributions [SK85, AWW87b, BlB85, BE88, Bel89b, BJ88, BK83, Cha86, CS84b, CS80, CW89, Hay89, Hor85, Jos82, Kha84, KK89, KS82, Lin88, Lin89, Mór83, NS87a, Nag88, OV85, Pre82, SS82a, SAU83, Sha89, SM86, Sin87, Weh86, YM89, Zel88]. disturbances [SC85]. divergence [NS85, NS87b, Tan87]. divisible [AP83, BP80, Gup88, Jac84b, KT81, Kag85, LJ87, Lon86, PS88, SP81, Seb84, SN83]. donors [GH89]. Dorfman [Gil85]. Dorfman-type [Gil85]. dose [Kha88]. doses [Gup89]. double [CA83, CA85]. doubly [BK86b]. drifts [Cha89b]. driven [MSV89]. dual [SP81]. Dubes [Flu89a].

EA [Daw85]. easily [Oma83]. Edgeworth [DHK87, PS84a, PS84b]. Edited [Roh88]. Editorial [Ano83e, Sri80, Ano80d, Ano81b, Ano82b, Ano82c, Ano83d, Ano84c, Ano84d, Ano85e, Ano85f, Ano86d, Ano86e, Ano87f, Ano87g, Ano88e, Ano88f, Ano88g, Ano89e, Ano89f, Ano89g]. Editors [Bha85, Mie82, Ylv81]. effect [Mik89, Muk81, PR85, PR88, RB85, SN83, yS89a]. effect-wise [Muk81]. effective [Kha88]. effects [CH82, Kuw86, KK88, MH84, Roe87]. efficiencies [Gup86]. Efficiency [ACR89, AK88, BDK80, BS82, BR86, CC89, EH84, Hab81, Kag88b, Kha81, Stu88]. efficiency-balanced [BDK80]. Efficient [SS88a, AE85, Kal85, Pay81, SS85a, Sri87b]. EGD [CH87a]. EGD-PBIB [CH87a]. Eidgenössische [Ano82d]. eigenvalues [Gas88]. Eisenhart [KKL88]. elements [Bil85, Kes86]. elimination [Bau89]. elliptical [Hay86, Hay89]. elliptically [KP87a]. Elsevier [Mie82]. Embedding [Mon86, SFM84]. Empirical [GH83, PSSZ87, Sti82, DV89, Fel89, GL89, LP88b, Lia89, Pou87, dAV80]. employing [VM83]. empty [Khu89]. Englewood [Flu89a]. English [Rot86]. enumeration [Ges86, Nie86, SS83c]. enumerative [KM86]. enveloping [Goo89a]. enzyme [Sha86]. epidemiologic [RMW85, RM88]. equality [BA82, Cam84, Chm81, CP80, KG89]. equation [Ang82]. Equineighboured [IF88]. equiprobable [HG86]. equireplicate [MB88]. Equivalence [Pra85, BB82, DK89, aS89, SS85b, Wie89]. equivalent [GK83a]. Error [PL88, BLT89, FG85, HD89, MH84, Oma85a, Oma85b, SM86, Vís86, Wie88a, Woo84]. error-correcting [HD89]. errors [CHT84]. Essential [Kal87]. essentially [GH80]. establishing [JS80]. Estimability [OC84, BM84]. estimate [Ahm83, DR85, DK89, Geo87, Hal86, KP87b]. Estimated [LB89, Amo86, Gou83, Hud85, Kos89]. estimates [AWW87a, BL84, CR82, Chi86, Fei85, GCG88, GKZ85, GK80, Loh85, Mac83, MPS84, Pra85, Tse86]. Estimating [Bar86, Ebr86, FS89, Gup84, SAU83, Dur89, GH89, Kal87, Man89, PK83, Tay89]. Estimation [CD83, CW83, CS80, Das89, DN84, Fra80, FG85, GH89, GLP84, HSB86, MB87, SK89, Sin85, SB81, YM89, AC89, Ara87, Ara89, AH84, BK81, BlB85, Bel89a, Bel89b, BD83, Böh85, BM86b, Car82, CK84b, Cha88b, Che89, EW89, Fun82, GD88, GK83b, Gno87, GT89, Gra89, GR80, HHL84, Iac85, Kap88, Kha88, KC81, Lia89, MSV89, MW87, Muk82, PSSZ87, PR89, Pil86, Rao81, Rou88, Rou89, Ruk88, Sah80, Sär82, SS88a, Sin81, SN89, Sin87, Tah89, Weg84, iY84, Zel88, Flu89b]. Estimative [JG85]. estimator [ACR89, BS85, BD88, CS84a, DV89, GV88, Kal85, KC81, LPY86, Mar87, Nai82, Oma85a, Oma85b, PL88, Ram82, SD85b, SC85]. Estimators [Wan87, BM86a, BM88, BLT89, BR86, BB82, BPS87, BH89, CF82, Cli89, CRS87, HJ81, JS81b, Kha84, KP87a, Kla89, KW88, Oma83, OV85, Pfa88, Ruy89, Sch87a, Sch89a, Sim80, SR82, Tik80, VM83]. evaluated [LP88b]. even [Gup89]. event [BB87]. evidence [DK88]. exact [Alb85, CP80, CLS87, MK82, Oma85a, Oma85b, PL88]. exchange [Fed89]. excluding [HRS88]. Existence [BT85, Kur88, SP81, Tse86, dL84a]. Expansion [Shi89, DHK87, PS84a]. expansions [Mes87, PS84b]. expectation [BE88]. expected [LGB88, LGB89, SS82a, SC83]. experiences [Wat82]. experiment [AT80]. experimental [Ecc80, GLP84, GPW87, GR85, Páz89, Wie85, Woo84]. experiments [Bea83, CL89, Cha88a, Cha89a, CB88, EH84, FP82, Gho81, Gho89a, Gup85b, Lew82, Mil88, Nac89, OS85, San81, SNS85, Sin82a, SW85, Zel88]. experimentwise [MH84]. Expert [Goi88, DeG88]. explanatory [Ylv81]. Exploratory [SM85, Ylv81]. exponential [Ara87, Ara89, BB82, CW89, Ebr86, GSV85, Jos82, Kha84, Muk84, Muk87, RB85, Sha89]. exponentiality [Ang82, PS87]. exposure [Gar83]. expressions [Fun82]. extended [Gup88]. extension [GT89, Jos89, Shi82]. extrapolation [HS88]. extreme [HSB86, Lo89]. extreme-value [HSB86].

F17 [Jos89]. F18 [Kag89]. F19 [Gho89b]. F20 [Goo89a]. F21 [Goo89b]. F3 [BP87, Geo87]. F6 [Goo87a]. F7 [Goo87b]. F8 [Goo87c]. factor [GA85, PRJ83, San81]. Factorial [Joh81, AT80, CCM89, EH84, Gho81, GK85, Gho89a, Gho89b, Gup85b, HD89, Kuw81, Kuw82, Kuw86, Kuw87, KK88, KN88, Kuw88, Lew82, Moo88, Mor84, Muk81, MB88, NS86, OS85, PRJ83, PR88, Roe87, SAM84]. Factorial-balance [Joh81]. factorials [CM86b, SO85]. factors [EK81, Kuw86]. failure [PS87, Wan87]. Fallacies [Jos89]. Families [Bus81, BlB85, KT81, Pfa88]. family [BB82, BK89, HK80, KH83, Mór83, SAU83, Sha89]. February [Ano85i, Ano88m, Ano89m]. few [AU89, SAU83, WO82]. fields [Sri86]. Finetti [Jos89]. finite [Bel89a, Gil85, Hor89, KS83, Mon86, Rob85, SSP87, SK89, Tho84, VM83]. First [Gas88, MK82]. First-order [Gas88]. Fisher [BP88, LP86, Lia87, SYT81]. fit [AU89, Ang82, Bel84, GCG88, JT85, KS82, LaR86, Moo84, WK87]. Fitting [AL84, Tay89]. Five [Tos89]. Fixed [GSV85, CS88, GV89, Mes87, Nac89]. fixed-width [CS88, GV89]. flats [SAM84]. fluctuations [Cha86]. fold [PR85]. fold-over [PR85]. forecasting [Wat82]. forecasts [Ven81]. Foreword [Sri86, Sri87a, Sri88]. forms [RM87]. formula [Csá86, HM82, Oma85a, Oma85b]. formulas [Hay82a, Tak86]. forum [And88, Ano86j, Ano89q, Gho88, Roh88]. Forward [Ano84e]. Four [Bro84, Don87, Hud83, MB86]. four-parameter [MB86]. fourth [Kos89, Ano82d]. fractional [CCM89, FS89, Kuw81, Kuw82, Kuw86, Kuw87, KK88, KN88, Kuw88, NS86, PRJ83]. free [GG86b, RG81]. frequency [Hab81]. full [SM82]. fully [Mar87]. function [AWW87a, AL84, Amo86, BH89, CB86, DHK87, Dur89, GG86a, GK80, Kos89, LP86, Lia87, MPS84, MW87, OV85, PR89, SAU83, Sin87, Wan87, Weg84]. functional [Ang82, Chi80]. functionals [Bab89]. functions [BK86a, Bel89a, CK84b, Cha88b, HG85, Hay89, JS81b, Kal87, Kha80, LP88a, LP88b, Oma83, Pou87, Sär82, Zah85]. Further [MMR84, Sob86, Jac88, SG84]. future [Sri86]. fuzzy [GCG88, Ovc88].

Galois [Vec84]. gambler [Gup84]. gamma [Shi89]. Gauss [BM88, BP89]. Gaussian [Amo86, MB87, PW85, Sti82]. GDRS [AP83]. general [BM88, BP89, BI86, HM82, JG85, MB83, RM87, Vís86, Wie88a]. generalised [SS87, SK85, dL84a]. Generalization [SYT81, Wat86]. Generalized [Ash81, Cat84, CS88, LS84, Seb85a, Daw85, DL80, HR88, MMR84, MB86, PR89, Ram82, Sär82, Seb80, SS83a, Wu89]. generating [CB86, LS81]. generator [Hor89]. Generators [Lew82]. geometric [CK86, Gov80, LM88, MM87, NS87a]. geometries [Bus81]. Getoor [Bha85]. GF [KS83]. GH [Daw85]. Ghosal [Ano80c]. given [BT85, BK86a, FK80, GLP84, GK85]. GJP [Wat82]. global [KW88]. Godambe [Ano89d]. Good [Roh88, GH83]. Goodness [Ang82, AU89, Bel84, GCG88, JT85, KS82, LaR86]. Goodness-of-fit [Ang82, AU89, Bel84, LaR86]. Gordon [Ano80c]. Gould [Cha86]. Graig [Goo87a]. graph [CF82, Fra80, Jac89, Now89]. graph-size [CF82]. graphical [FH88]. graphs [Ton81]. Gray [LS81]. Group [GPW87, HX87, Zid88, AP83, BP80, BFPR84, DV89, GH89, Gil85, Gup88, Jac84b, KT81, Kag85, LJ87, Lon86, PS88, SP81, Seb84, SN83, WR87]. group-divisible [SN83]. grouped [Gil86b]. grouping [KS82]. growth [BD88, Kub89, SS86b]. gum [PK83].

H [Ylv81]. Hadamard [Daw85, Seb80, dL84a]. half [Bar82]. Hall [Flu89a]. handling [CA89]. harmonic [BE88]. harmonizable [Pou87]. Hausdorff [Sta89]. Hazard [Rou89, AWW87a, DN84]. hazards [PR89]. Helmert [Alb85]. Hermitian [CP80]. heterogeneity [Not85]. heterogeneous [SS82a]. heteroscedastic [Car82]. heteroskedasticity [Das89]. higher [Sch89b]. higher-order [Sch89b]. Hochschule [Ano82d]. Hoel [Spr85]. hole [Bro84]. Holland [Flu89b]. homogeneity [Sha89]. honeycomb [CK86]. Honggerberg [Ano82d]. Hook [AA86]. Horvitz [Ram82, Rao83]. hybrid [Ebr86]. hypergeometric [WR87]. hyperplanes [KK89]. hypotheses [CS84b, EL86, Hor89]. hypothesis [Hay86, Wie88a].

identically [Mus87]. identification [PL89]. identify [HX87]. identifying [WR87]. II [BN86, KH83, LM88, MB87]. III [BI84, KKL88]. Ikeda [Mie82]. imbalance [Pre87]. imperfect [Bha87, Goi88]. Implications [Sär82]. important [GH88]. Improved [Ruk88, Sah80]. improvements [BLT89, Goo87b]. improving [Sac82]. inaccuracy [GCG88]. inclusion [KGN85]. Incomplete [NB85, PG88, Agr87, Cam84, Gil86a, KKL88, Kes86, MR89, PS87, RFS86, Sas88, Sin85, SB86]. incorrect [MB83]. increase [WK87]. increasing [San84, Wan87]. independent [DORG83, GH88, PC81]. independent-samples [PC81]. Index [Ano83c, Ano89a, Ano80b, Ano81a, Ano82a, Ano83b, Ano84b, Ano84a, Ano84f, Ano85b, Ano85c, Ano86a, Ano86b, Ano87c, Ano87d, Ano88b, Ano88c, Ano88d, Ano88h, Ano89b, Ano89c]. Indirect [TL87]. individual [EK81, FW87]. inequality [CM86a, KK89, Mór83]. Inference [Bel84, CH83b, MP88, Oma85a, Ruk83, Sch89a, SS83b, SG86, Cha89b, Cue88, FM88, Goi88, aS89, Sin82a]. Inferences [FK80, MB85]. inferential [RM88, SSP87]. infinite [Cli89]. influence [CR82, JJ83, JG85]. Information [FP82, Bha89, CG83, FW87, Kuw81, KN88, NS85, NS87b, Páz89, Puk80, SS85a, VM83]. initiate [LKM80]. insight [Tan85]. inspection [BZ87]. integral [AH84]. inter [Fei85]. inter-block [Fei85]. interacting [SNS85]. Interaction [Gup86, KKL88, MH84, SM87]. interactions [San81]. interesting [Kuw86, KK88]. international [Ano82d]. interpersonal [FK80]. interpolation [Spr87]. intersections [Seb82]. interval [Mes87]. intervals [CS88, GSV85, GSvR88, GV89, KSY82, LGB88, LGB89, RMW85]. intra [Fei85, Wie85]. intra- [Fei85]. intra-class [Wie85]. intraclass [KG89]. Introduction [WS88]. intuitive [Goo87c]. Invariance [Bab89, PRJ83, PR88]. Invariant [Chm81, CG83, GPW87, GKZ85, Hay82a, Sin82a, YM89]. Inverse [BP89, Amo86, Oma85a, Oma85b, Sah80, SD85b, SB81]. Inverse-partitioned-matrix [BP89]. inverses [Cat84]. investigation [WB80]. involving [BlB85, CB88, DN88, SNS85]. ISI [Ano86f]. isotonic [Dyk81]. issue [Sri84, Sri86]. issues [Sri86]. item [HX87]. Iterative [Gup85a, Fel89]. IV [GN85]. IV.3 [AT80].

J [Flu89b, Roh88, Sch89a]. Jack [Ano85k, Sri85b]. jackknife [KC81]. Jain [Flu89a]. January [Ano84h, Ano85g, Ano88j, Ano89j]. Jenkins [Wat82]. Journal [CH83b, Oma85a, Sri86]. JSPI [Sri85c]. July [Ano84g, Ano88q, Ano89o]. June [Ano83j, Ano89l].

Kendall [WB80]. Kernel [BH89, DK89, GK80, KP87b, LPY86]. Kernel-type [BH89]. kernels [Sch89b]. Ketellapper [Ylv81]. keys [Pay81]. Kiefer [Sri85b, Ano85k, Wie89]. kinetics [Sha86]. Kleffe [Flu89b]. knowledge [Goi88]. known [Muk82, Sch84]. Kronecker [Gup83, Gup85b, Gup86, MB88]. kurtosis [LaR86].

L [Bha85]. lack [Moo84, WK87]. Lancaster [BR85]. Langevin [Wat83]. Large [BH89, Wat83, Aic83b, AGLS84, Bjø85, Ruk83, Wu87, dL87]. largest [AM89]. latin [SFM84, Bai88]. Lattice [Men86, AA86, CK86, Ges86, HM86, Sha86, Tak86]. lattices [LM88, MM87]. law [Goo87c]. Laws [AGLS84]. layouts [GG86b, Wie88b]. least [BK81, BL84, Cli89, Pes84, Whi87]. lemma [Gup83, Sha86]. length [GSvR88]. lengths [Rot86]. less [Sri87b]. level [Mor84, Vís86]. Levine [Spr85]. Library [Wat82]. life [Ahm83, BE87, BK89, BK83, CS80, Ebr86, MP88, PSSZ87, SS83b, Sin87, Sri87b]. life-time [CS80]. likelihood [Aic83b, GT89, GR87, Gra89, Pfa88, Tse86]. likely [BB87]. limit [GV88, SP89, Wu89]. Limits [Mil88]. line [Gup89, NB85, PG88]. Linear [BK86a, BL84, Cox81, Sch84, AS85, Amo86, BK81, BM84, BM86a, BM88, BP89, BLT89, Bak89, Bea83, Bel89a, BPS87, BI86, Car82, Cha88a, Cha89a, CA89, Chi80, CRS87, CM86c, DN84, DL84b, EW89, GA85, GK83b, GKZ85, GH88, JG85, LP88a, LN82, MB83, McC86, MB87, Men86, Nag88, Oma85a, Oma85b, PL88, Pil86, Puk80, PS84a, PS84b, RF84, Sär82, Sch87a, Sch89a, SD85b, SG86, SM86, Sin82a, SM85, Tam81, Tor84, Wie88a, Wu87]. linearity [OC84]. literature [Goo87b]. Local [AK88, MSV89, SM85]. Locally [Hod88, HPT84, Nab86, Sen84a]. Location [FM88, Hal86, HJ81, JS81a, Jur80, Kap88, PS84b, SAU83]. location-scale [SAU83]. Log [Bea83, AK87, BM87, Cha88a, Cha89a, DN84, Nab86]. log-linear [DN84]. log-logistic [AK87, BM87]. log-rank [Nab86]. logistic [AK87, BK86b, BM87, CL89, Gra89]. Lognormal [Ruk88]. long [BM89b]. long-run [BM89b]. longitudinal [Sin85]. losing [Gup84]. loss [Ang82, Bha89, Whi89]. loss-of-memory [Ang82]. low [YMO87]. low-dimensional [YMO87]. Lower [Fin88, Sim80].

M [JS80, Mie82, Roy82, SS86b, Wat82]. M-methods [SS86b]. main [MH84, PR85, PR88, SN83, yS89a]. main-effect [SN83, yS89a]. majorization [GW80, KK89]. managerial [Doc88]. many [Car82]. mapping [BJ88]. March [Ano83h, Ano84j, Ano88p, Ano89p]. Markov [ABGZ84, BM88, BP89, BM89a, BM89b, GG81, Hor89, Nag86a]. Markovian [Nag82]. Marshall [CW89]. martingale [Bos86]. Martinus [Ylv81]. Master [Ano84f, Ano88h]. matching [Sid82b]. mathematical [Ano82d, Sid82a]. matrices [CKM87, CP80, Daw85, Gas88, GW80, Gup88, Hay86, HP85, Kuw81, Mas88, Páz89, San84, Seb80, SG84, dL84a, dL87]. matrix [BP89, BLT89, Hay82a, Hay89, KW85, Khu89, KN88, MB83, SAM84]. maximal [CH83a, CH83b]. maximize [Bar82, Puk80]. maximum [CKM87, GR87, Gra89, Pfa88, Tse86]. May [Ano84l, Ano86g, Ano89i]. Mean [BLT89, PL88, Ahm83, Ara87, Ara89, BK89, Che80, CW83, CS84b, GSV85, GR80, HM83, Loh85, LGB88, LGB89, Muk82, Nai82, Oma85a, Oma85b, PSSZ87, Pra85, SK89, Sha89, SB81, SR82]. means [AC89, Bab88, Bel89a, BD83, BL84, CG83, DN88, DV89, LB89]. measure [Böh85, Gas82]. measurements [CH82, FG85, Roy88, SM87]. Measures [Moo84, Cox81, JG85, RMW85, Tan87, Wie88b]. measuring [FG85]. median [CM86c, Kha88, Mus87]. medical [LKM80]. mediocre [HX87]. memory [Ang82]. Mendelsohn [Bil85]. Method [SN89, BP89, BK86c, CS89, Fun82, Gas88, Ges86, HG85, PK83, yS89b]. methodology [RM87]. Methods [PS88, BDK80, CK84a, DKPS88, EH83, Gil86b, Sid82a, SS86b]. metric [Mie86]. metrics [DKPS88]. mg [YC80]. mg-procedure [YC80]. min.hypers [HD89]. Minimal [KKL88, KS83, Not82, RS84]. Minimax [Ber80, GM85, Pil86]. Minimizing [SC83]. Minimum [BlB85, Mor84, Aic83b, GCG88]. misclassification [CHT84]. MISER [Pre87]. Mises [Bab89]. missing [Gho89b, Moo88, SS88a]. mixed [Das89, FS89, Gho89a, GK83b, GKZ85, Gno87, Jos82, Khu89, Pei88, Sin81, SC85]. mixing [Gup82, Har88, Rou88]. mixture [Amo86, Sin87]. mixtures [Cha88a, Cha89a, CB88, Hor85, Pfa88, Shi89]. mle [Sac82]. MML [Tik80]. Model [HM88, HS88, KSY82, Spr85, Tho84, BM88, BP89, BK83, Bha87, BI86, Cha88a, Cha89a, CS84a, Das89, DORG83, GA85, GR85, Hjo89, HJ81, JG85, Kai89, KKL88, Khu89, Kub89, MB83, MB89, MB87, PR89, Sen84a, Sen84b, Tro82a, Vís86, Wie85]. model-robustness [MB89]. modelling [Ano82d, Mau85, Wat82]. models [BM86a, Bak89, Bea83, BM89a, BM89b, Car82, DN84, EW89, Fin88, Gas88, GK83b, GKZ85, Goi88, GH88, Mie87, Not89, PL88, Pei88, Roe87, Ruk88, SM86, SM82, Tor84, Tro82b, TW86, Tse86, Wie89, Yan88]. Modified [BJ88, BEH84, LP86, Lia87, Oma83, PK83, WB80]. mols [Bro84, SD86]. moment [CB86, DR85]. Moments [BM87, BK86b, CB86, FS89, Jos82, KPY83, Lin88, Lin89]. Monday [Ano82d]. monotone [BK89, Gut87, PS87]. monotonic [Gup87]. morbidity [LKM80]. mortality [RMW85]. most [BB87, Gaf82, Gut87, Hod88, SS88b]. multi [Jai81, SB86]. multi-dimensional [SB86]. multi-stage [Jai81]. multicomponent [Ahm83]. multidimensional [Hal86, Har88]. multinomial [Ber80]. multinomials [HG86]. multinormal [CG83]. multiple [AL84, ACR89, BS85, Bau89, Gar83, Geo87, GH80, GM85, SS82b]. multiples [BPS87]. multiplicity [CH83a, CH83b]. multiresponse [WK87]. Multivariate [BE88, Bea83, BB82, Bha84a, Bou81, Car88, CHT84, Chi80, CRS87, DORG83, Flu85, GS87, GG86c, JG85, LB89, Not81, OZ84, Sin81, Zah85]. mutually [dL87]. MV [Jac88, SS86a]. MV-optimal [SS86a]. MV-optimality [Jac88]. mvue [OZ84].

Narayana [KM86]. near [PS84b]. nearly [LN82]. nearness [KP87a]. necessarily [SC85]. neighbour [SS85c]. Nested [Lon86, IF88, Lon81, SB86]. Nijhoff [Ylv81]. NJ [Flu89a]. NN [Mac83]. no [Khu89]. noisy [GG86a]. Non [MPS84, MB85, Mie86, MB88, Bos86, CMS82, Mus87, Nag82, San84, SNS85, SM85, WR87, Wie88a, dL84a]. non-adaptive [WR87]. Non-asymptotic [MB85, CMS82]. Non-equireplicate [MB88]. non-existence [dL84a]. non-identically [Mus87]. non-increasing [San84]. non-interacting [SNS85]. non-linear [SM85]. non-local [SM85]. non-Markovian [Nag82]. Non-metric [Mie86]. non-null [Wie88a]. Non-parametric [MPS84]. non-sequential [SM85]. non-uniform [Bos86]. nondeterministic [Fin88]. Nonexistence [Kai89, FH88]. nonlinear [BL84, Páz89]. Nonnegative [BM84]. nonnegatively [PC81].

Nonnegativity [GKZ85]. Nonparametric [Che89, CH87c, DN86a, GG86a, Rou88, SS82b, WS84, Bha84a, CS88, Pfa88, Sch89b, SP89, Sin81, Weg84, Whi87]. nonregular [Kha80, Mór83]. nonresponse [BdBP82]. nonsingularity [KW85]. nonsymmetric [SM86]. nonuniform [Wu89]. normal [Ahm83, AC89, Bar82, BD83, Che80, CS84b, CRS87, DV89, GG86c, GM85, HM83, LB89, Muk82, Nai82, OC84, OZ84, RJDT83, Sch84, SR82, SC85]. normality [Alb85, BR85, Bos86, HR88, JSV87, LaR86, Lo89, Now89, Woo84]. North [Flu89b]. North-Holland [Flu89b]. Note [Mór83, AH84, Bab88, CC89, CM86a, CA89, DD89, Flu85, Geo87, Hay82b, Jos82, Kag88b, Kal85, Kha81, Kha82, KP87a, LP86, Lo89, Nag86a, Rao81, Sch87a, Sch89a, SM82, Sri80, Tam81, Tay89]. notice [Ano83e]. November [Ano89n]. nuisance [Bak89, Bel84, Bha89, CK84b, KK88]. null [Wie88a]. number [Bjø85, DN88, Fra80, Gup89, Mor84, SMK88]. numbers [AGLS84, KM86]. Numerical [Böh85].

ø-mixing [Har88]. observation [Whi89]. observations [Bar82, CCM89, GG86a, GR82, GCG88, KM89, MK82, PC81]. obtained [Ven81]. obtaining [EH83, Gas88]. occurrence [Gar83]. occurrence-exposure [Gar83]. October [Ano82f, Ano83g, Ano84k, Ano89k]. official [Ano86f]. Olkin [CW89]. One [CM88, CD83, GK83a, Gov84, HX87, Jur80, SS82b, Wie88b]. One-sample [CM88, Gov84]. one-way [CD83, SS82b, Wie88b]. ones [RJDT83]. opinion [DeG88]. opinions [Zid88]. Optimal [Bel89a, BI86, CL89, Cha88a, Cha89a, Fed89, Gil85, GG86c, JM89, Kha88, LaR86, Maj86, Muk82, Not85, Not89, PP87, Roe87, SM87, SS83d, Spr87, Tro82b, Weg84, BI84, CCM89, CKM87, CLS87, Dur89, Fel89, Gaf82, Gho80, Hed84, HM88, Hor88, Hor89, HPT84, Kal87, KY88, KC83, KF84, Kra85, Kuw82, Kuw88, Lau88, LJ87, Lee88, Man89, MR89, Mas88, MK82, MR85, Nab86, NS86, PR88, Rob85, Roy82, Sen84a, SO85, Spr85, Tam81, Tos80, Wie88b, WK87, Yeh88, YMO87, SS86a]. Optimality [GR82, MB83, SS88b, SW85, BS89, Cha88b, CA83, CMS82, EK81, GR85, Hay82b, JS80, JWM83, Jac84b, Jac85, Jac86, Jac89, KM89, MB89, PR85, Pes84, Puk83, RS84, SM82, Tah89, Jac88]. optimization [Fed89]. Optimum [CG83, Wie85, GK85, Man89, Mik89, WR87]. Order [Bel89b, AS85, AK87, AU89, Ara89, ABGZ84, BK86b, BM87, Bro84, CP84, Chi86, CS88, CMS82, Gas88, Gov80, Gov84, GG81, Hud83, HJ81, Jos82, KPY83, Kos89, LP88a, Lin88, Lin89, MLF81, MK82, Nag82, Nag86a, Nag86b, NS87a, Nag88, Not82, Rüs85, SAU83, Sch89b, SFM84, SK85, Tah89, ZdK81]. ordered [GG86b, Sac82]. Orderings [Tor84, GPW87, ZdK81]. Organization [Sri85c]. orientation [Pre82]. orthogonal [FHM84, GN85, Kag88a, KW85, Kuw86, KK88, MLF81, Muk81, SFM84, yS89a, dL87]. other [Goo87b]. outliers [GK85, KK89, PL89, dAV80]. Overall [MR89, RM88]. ovoids [Vec84].

P [GY85, Ylv81]. Pages [Ano80e, Ano80f, Ano80g, Ano80h, Ano81c, Ano81d, Ano81e, Ano81f, Ano82h, Ano82f, Ano82e, Ano82g, Ano82i, Ano82j, Ano83i, Ano83j, Ano83h, Ano83g, Ano83f, Ano84i, Ano84h, Ano84g, Ano84j, Ano84l, Ano84k, Ano85h, Ano85j, Ano85i, Ano85g, Ano86g, Ano86h, Ano86i, Ano87h, Ano88m, Ano88j, Ano88q, Ano88p, Ano88k, Ano88n, Ano88i, Ano88l, Ano88o, Ano89m, Ano89j, Ano89l, Ano89o, Ano89p, Ano89i, Ano89n, Ano89k, Ano89h]. paired [Bea83, CB88, CS84a, EH84, PC81]. paired-samples [PC81]. pairing [GR87]. pairs [Gho89b, Moo88, PC81]. Pairwise [FHM84, CH87b]. paper [Ano89d]. Papers [Ano85k, Ano87e]. paradoxes [Sob86]. parallel [Flu85, Gup89, NB85, PG88, SAM84]. parallelism [CP84, Chi86, RG81]. parameter [JS81a, Lia89, MB86, Pil86, Ruk83, San84]. parameters [Bak89, BlB85, Bel84, Bha89, Car82, CK84b, Chm81, Ebr86, FS89, Kub89, KK88, Muk84, Sac82, Sin87, Tay89]. parametric [BK86a, EL86, GL89, MPS84]. part [HK80, KH83]. Partial [LM88, MM87, Sin81, SS83c, SS85c]. partially [Agr87, EW89, Kuw87, Kuw88]. particular [Pfa88]. partition [DN86b]. partitioned [BP89, Mau85]. partitions [Gil85]. path [Ges86, Men86, Sha86]. paths [AA86, HM86, Tak86]. Paul [Ano82k]. PBIB [Kag86, CH87a, Kag82, Moh81, SD82]. PBIBD [GRS83]. PBIBDs [Str84]. PD [SD86]. PD-cycles [SD86]. Pearson [HG86]. percentile [GW88]. permutation [PRJ83, Ruk83, Sta88]. permuted [DN86b]. perturbed [LP88b]. phase [HM83]. Pitman [KP87a]. Plann [Sch89a]. Planning [CH83b, Oma85a]. plans [Gho80, HRS88, Mil88, SN83, yS89a]. plot [Roe87]. point [LP88b, Man89, Not82, Tah89]. points [DL84b, Gho89b, Moo88]. pointwise [Geo87, KP87b]. Poisson [BJ88, BEH84, DKPS88, HG86, RJDT83]. polynomial [CLS87, KN88]. polynomials [Hay82a, Kuw81, Nie86, Wat86]. polytomous [TW86, Tse86]. population [Bar86, Bau89, Bel89a, CG83, CW83, GG86c, GY85, GL89, Hay86, Hay89, Hay82b, Muk84, Muk87, Ris85, Rob85, SSP87, SK89, Tho84, VM83]. populations [Amo86, Bar82, Bjø85, Gil85, GH83, GM85, Nai82, Sch84, SS83c]. possible [Not89]. posterior [AWW87b]. power [CP80, Hab81, Tay89, WK87]. powerful [Gut87, Hod88]. pp [Mie82]. pp. [Wat82]. PPS [ACR89, BS85]. Practical [Wat82]. precise [NS86]. precision [SG86]. Prediction [Pei88, BK81, BR86, Jos88, LP86, Lia87, RMW85]. predictors [PL89]. Preface [Bas87, Moh86, Sri84]. preliminary [BA82, SD85b]. Prentice [Flu89a]. presence [Bel81, Bha89, CK84b, GR82, GK85, GV89, Hec87, Nac89, Not85]. preserving [GH89]. prevalence [GH89]. Priestley [Geo87]. principal [KW85]. principles [Bab89]. prior [VM83]. probabilistic [Ges86, Zab88]. probabilities [CA83, Csá86, Vís86, Wu87]. Probability [Flu89b, GG81, AM89, Böh85, Che83, Fin88, Gov84, Gup84, HG85, Hay82b, HG86, Jos89, KGN85, MPS84, ZB81]. probing [Men86]. problem [BdBP82, BJV89, CH87c, Gup84, Oma85a, Oma85b, RG81, Wat86, WS84]. problems [Bel84, BM86b, Pes84, SS88a, Seb85b, Sin82a]. procedure [Bau89, Goo89a, Tan85, Tos80, YC80]. Procedures [Sac82, BZ87, Bjø86, CH87b, GW88, GH80, GY85, JS81b, Kub89, LB89, Muk82, Muk84, Muk87, RM88, WS84, ZB81]. process [Hor89, PW85, ZdK81]. processes [Bha85, Fin88, Gar83, Mau85, Pou87, Sta88, Sti82]. product [GV88, Gup83, KPY83, Lin89]. product-limit [GV88]. Professors [Ano89d]. profile [Bha84a]. Projective [EFJ82]. proof [FH88]. proper [Jac84a]. Properties [Gar83, SC85, Weh86, Aic83b, BD88, BH89, CMS82, Gas82, GR87, Gov80, GK80, Gup87, LP88a, LP88b, LPY86, MMR84, Pou87, Puk83, SG84, Sta89]. property [ABGZ84, GRS83, GG81, HM86, Jac86, KM86, Loh85, MR85, Nag86a, SP81, SS83d]. proportional [DN84, KGN85, PR89]. proportions [Cam84]. Proschan [Bha87, May87]. PS [Ram82]. Pseudo [Chi80, Mon86]. pseudo-complements [Mon86]. publication [Wat82]. Publications [Ano85k]. published [Ano85k]. Publishers [Ano80c]. Publishing [Mie82, Ylv81].

Quadratic [GK83b, Gno87, YMO87, GKZ85, Man89, PP87]. quadrics [Cas89]. quadrilaterals [Mon86]. Qualitative [Cue88]. quantal [JJ83]. quantile [LPY86, SAU83]. quantiles [GSvR88, GV88, GV89, HSB86]. quantitative [EH83]. quasi [GT89, SS87]. quasi-likelihood [GT89]. quasi-residual [SS87]. Quasicyclic [NS87b]. quest [DL84b]. question [Seb82]. question-sets [Seb82].

R [Bha85, Flu89b]. random [Ara87, Ara89, BJ88, Cha86, GSV85, GR87, GR85, KW85, KS82, LP88b, Mus87, Nie86, PR89, Rao83, Roe87, Rou88, Seb85b, SS83a, Sen84a, SK85, SN89, SW85, Sta88]. Randomization [Jos88, WO82, GS87, Kai89, PRJ83, PR88, Weh86]. randomized [Bou81, CA89, EH83, Kai89, MB89, SK89]. Randomly [Sho89]. randomness [Weg88]. range [SS82a, SC83]. Rank [CP84, Alb85, Bak89, BN86, Chi86, DR85, Gov84, Har88, Hod88, Nab86, PS84a, PS84b, PW85, RF84, Sen84a, SK85, SM86, SM82, Ton81, Wu87, Wu89]. Rank-order [CP84, Chi86]. Ranked [Che83, Shi82, YC80]. ranking [ZdK81]. ranks [aS89]. Rao [Flu89b, Kag88a, LS84, Seb84, Seb85a, Kha80, Mór83, Sim80]. rare [GH89]. Rate [Jur80, Mac83, AWW87a, Gar83, KP87b, MH84, PS87, Rou89, Wan87]. Rates [Wu87, Bos86, EW89, Gou83, Lia89]. ratio [Fun82, GD88, Gov84, Hay82b, Iac85, KC81, LGB89, Rao81, Tay89]. ratios [BI86]. re [GR87]. re-pairing [GR87]. recapture [Zel88]. record [ZdK81]. recovering [GG86a]. rectangle [FHM84]. rectangular [Man89, yS89b]. Recurrence [KPY83]. Recursive [CC84, Ecc80, KW88, MPS84, Pra85]. redundant [De 87]. referee [Ano85a, Ano87a, Ano88a, Pur86, Pur89]. Reference [Flu89a]. regarding [BP87]. region [Man89]. regions [Chi80, Sko88, Sko89]. regressing [BP87, Bor82]. Regression [EL86, AL84, BP87, Bor82, CR82, CL89, CP84, Chi86, Cli89, CLS87, CM86c, DK89, DORG83, Dyk81, Flu85, Geo87, Goo89a, Gra89, GK80, GR85, GH88, Hod88, Iac85, Kap88, Kra85, KP87b, LN82, MB87, Mie87, Not81, Not89, OC84, OZ84, Pes84, Pil86, PP87, Puk80, PS84a, RG81, Ruk88, Sär82, Sch89b, Sen84a, Sen84b, SG84, SC85, Tay89, Ven81, Whi87, Wie85, Wie88b, Wie89, Woo84]. regressions [Nag88]. regressive [Cha89b]. Regular [Seb84, BK86c, Bus81, Jac89, SN83]. Regularity [MM87, iY84]. rejective [CH87b]. related [SK85, SS85b, Vec84, ZdK81]. relating [Jos89]. relations [HJ81, KPY83]. relationship [Chi80]. Relationships [EK81, Lin88, SG86]. relative [EH84, Hab81, SB86]. Reliability [De 87, AWW87a, AWW87b, Gup82, Gup87, May87, RM87, Sin87]. remarks [CRS87, McC86, Rüs85]. repair [BK83, Bha87]. repeated [Bil85, CR82, CH82, FG85, KM88, RFS86, Rob85, Roy88, SM87, SS82b]. replications [BT85]. representation [YMO87]. representations [CS88, GV88]. required [Gou83]. Resampling [BJV89]. research [Ovc88]. Residual [Mus87, Ahm83, BK89, CH82, Gup87, MMR84, SS87]. residuals [DORG83]. resolution [BI84, CCM89, Gho80, GK85, GN85, Kuw81, Kuw82, Kuw86, Kuw87, KK88, KN88, Kuw88, NS86]. resolution-IV [GN85]. Resolvable [Rok84, AT80, Sas88, SD82, dL87]. resources [SSP87]. respect [BK86a, Vís86]. response [AP83, Bou81, CA89, EH83, Gou83, Kos89, Man89, SK89, TW86, Tse86, Yan88]. responses [Hud85]. restricted [BM86a, Bel89b, GR80, Nie86, Sen84b]. restrictions [BP89, Pil86]. resulting [Gho89b, Moo88]. Results [HLT89, Amo86, Bha87, CA85, Jac88, SS85b, Vec84, YMO87]. Review [Ano80c, Bha85, Flu89a, Flu89b, Mie82, Roh88, Sid82a, Wat82, Ylv81]. rewards [BM89b]. Richard [Flu89a]. ridge [Woo84]. right [MP88]. right-censorship [MP88]. risk [Oma83, RM88]. risks [RMW85]. Robust [Car82, Hec87, Iac85, Kap88, LN82, Pes84, Zel88, HM88, HS88, KSY82, SM86, Tan85, Tik80, Whi89]. Robustness [Gho82, Tik80, BT87, Cue88, Gho80, GRS83, GK85, MB89, Spr85]. Ronald [Ylv81]. rotatable [Hud83]. rotation [Bel89a, Jai81]. row [AE85, BS89, EM85, IF88, SNS85]. row-column [AE85, BS89, EM85, IF88, SNS85]. RSRR [RMW85]. ruin [Csá86, Gup84]. rule [BM86b, KW88, PL88]. rules [GH83, HPT84, Sti82]. run [BM89b, CM88]. runs [Mor84].

S [Mie82, Roy82, Bau89, JS80]. Sabatier [Ano82k]. Sample [Gou83, AM89, BN86, BJV89, BH89, Che80, CM88, Day85, FS89, GR87, Gov84, HJ81, Jai81, JS81a, Jur80, Mar87, RG81, RF84, SS82a, SC83, SN89, Sko88, Sko89, Wat83]. samples [Bha84a, FS89, JSV87, LPY86, PC81, Ruk83, SN89, Tik80, iY84]. Sampling [AC89, HRS88, SS83c, ACR89, BS85, BZ87, Bel81, BR86, Bel89a, BEH84, CA83, CA85, Che83, CHT84, Jos88, KW89, KGN85, Mil88, PK83, Pou87, Rao83, Rob85, Sah80, SSP87, Shi82, Tho84, VM83, YC80]. satisfying [Bjø86, iY84]. Saturated [Not81, MK82]. Scalar [BPS87]. scale [Alb85, Chm81, Muk84, RB85, SAU83, Shi89]. scheme [CW83, Gup88, Jai81, KGN85]. schemes [CHT84, LM88, VD89]. Science [Ano80c, Flu89a]. Scientific [Mie82]. score [AK88, BH89, JS81b, OV85]. scores [HR88]. Scrambled [EH83]. screening [GA85]. Search [OS85, SO85, AT80, CM86b, Gho81, GA85, Mor84, Seb82, Seb85b, Tos80]. searches [Fel89]. searching [Tuz82]. Second [Ara89, CMS82, HJ81, Tah89, Not82]. Section [Ano86f]. selected [DKPS88, SAU83]. Selecting [GH88, GL89, Ris85, AM89, BB87, GH83, Muk84, Muk87, San81]. Selection [Pay81, Ber80, Bjø85, Bjø86, GY85, Hjo89, HPT84, PL89]. selective [Che83]. Semi [Bai88, BM89a, BM89b]. Semi-Latin [Bai88]. semi-Markov [BM89a, BM89b]. semigroup [DKPS88]. Seminar [Bha85]. sensitive [EH83]. Sensitivity [Vís86, Cox81]. separatory [McC86]. September [Ano83f, Ano89h]. sequences [DN86a, DN86b, Rou88, San84]. Sequential [Ara87, GV89, Gra89, JS81b, Muk84, Seb82, Sim80, Ara89, Bau89, CS88, DORG83, Gov84, Hay82b, Mar87, Muk82, Sko88, Sko89, SW85, SM85, Tah89, Ven81]. Sequentially [CH87b]. Serial [Aic83a]. Series [Flu89a, Flu89b, Wat82, BI84, Cha89b, CH87a, Cha86, KK88, ZB81, Wat82]. services [Ano85a, Ano87a, Ano88a, Pur86, Pur89, Sri85a]. set [Che83, DORG83, MLF81, SFM84, Shi82, YC80]. sets [DN88, Maj86, Ovc88, San84, Seb82, SNS85, dL87]. setups [Kra85]. several [Bha84a, CP84, Chi86, KG89, RMW85, RM88]. Shapiro [AU89]. Shapiro-Wilk [AU89]. Sheffer [Nie86]. shift [RB85, ZB81]. short [KP87a]. Shrunken [SR82]. Sidak [RMW85]. Sidak-type [RMW85]. sided [CP80]. signal [Bel84]. signed [PS84a, PS84b, PW85]. significance [BA82, CR82, SS82b]. significative [CM86c]. similarities [Gil86b]. Simple [BK81, BK83, Puk83]. simplex [Mik89]. simplicity [Sin82b, SS85b]. simulation [Sko88, Sko89]. Simultaneous [PL89, BD83, MH84, RMW85, RM88]. sine [Goo87c]. single [Whi89]. Singular [Gho89b, Moo88]. six [Gaf82]. size [BGH84, CF82, Don87, GLP84, KGN85, Mar87, Roy82, Seb84, Seb85a]. sizes [AM89, BT85, Gou83, LJ87, SN89]. skewness [LaR86]. slippage [JS81a]. slope [Bar82]. small [Lon86]. smoothing [Car88]. sobel [Kha81]. sobel-wald [Kha81]. solution [Chi80]. solutions [SD85a]. Some [Aic83b, BDK80, BP80, Cas89, CK84a, CM86b, CA85, CP80, Csá86, GA85, GR87, HPT84, JS80, Jac85, Jac88, Jac89, KT80, KT81, KC83, McC86, Mie86, NS87a, Nag88, PR85, Pou87, SSP87, Sin82b, Sta88, Str84, yS89b, Tak86, Vec84, WR87, YMO87, Zah85, BB82, Bou81, CRS87, Gho81, Gil86b, Hay82a, Kuw82, Kuw86, Kuw87, KK88, SS88a, SS85b, VD89, Weh86, Goo87b]. sortings [FG85]. Soviet [Ovc88]. space [KS83]. spaces [DL84b, Vec84]. spacings [DHK87]. Spatial [Bel81, Ylv81]. special [Sri86]. specific [LKM80]. Speed [Sin81]. spheres [FM88, Ruy89]. spherical [Chm81]. spherically [Wie88a]. splines [Car88, MW87]. split [Roe87]. split-plot [Roe87]. spread [Kla89]. spring [Jac86]. square [BLT89, MB85, Pra85]. Squared [PL88, EL86, HG86, Moo84, Oma85a, Oma85b]. squares [Bai88, BK81, BL84, Cli89, LGB88, LGB89, MLF81, Pes84, SFM84, Whi87, dL87]. SRGD [Kag89]. Srivastava [GK85, NS86]. stability [KC81]. stage [BR86, Jai81, Kub89, Mar87, Muk84, Muk87]. stand [De 87]. stand-by [De 87]. standard [FG85, Jac88]. Standardization [Sid82a]. Statist [Sch89a]. statistic [Fei85, MB85, Wie88a]. Statistical [And88, Ano86j, Ano89q, CH83b, Gas82, Gho88, Mau85, Oma85a, Tan87, CF82, CK84b, Cha88b, Cox81, Goo87b, Mie86, aS89, WS84, Ylv81, Mie82, Roh88]. statisticians [Ano86c]. Statistics [Flu89b, AK87, AU89, Ano86f, ABGZ84, AGLS84, Bab89, BK86b, BM87, BK89, Bha89, CS88, DR85, EL86, FW87, Gov80, GG81, Har88, HR88, JSV87, Jos82, KKL88, KPY83, LP88a, Lin88, Lin89, Nag82, Nag86a, Nag86b, NS87a, Nag88, Nie86, Now89, Pre82, PS84a, PS84b, PW85, Rüs85, RF84, SAU83, SK85, SP89, Sho89, Tik80, Wu89]. Stein [BM86b, Oma83, PL88]. Stein-rule [PL88]. step [MP88, San84, SS83b]. step-stress [MP88, SS83b]. Stochastic [ZdK81, Bha85, Pou87, Sta88]. Stopping [Gov84, KW88]. strategies [CA83]. strategy [GR87, Rao83]. stratification [CW83]. stratified [Jai81, Rao83]. strength [Kur88, Lon84, SMK88, SO85]. stress [MP88, SS83b]. string [Jac86, MR85, SS83d, Tuz82]. string-searching [Tuz82]. Strong [Ruy89, Mac83, Pra85]. strongly [Roy88]. structural [SG86]. Structure [Nag86b, SD86, Hwa86, Jac84a, Joh81, Khu89, Nag82, Ven81]. studentization [RF84]. studies [Gra89, Gup82, RM88, SM85]. study [Fel89, Goi88, HP85, RMW85, Sko88, Sko89]. Subhypotheses [Sen84b]. subject [SD85b]. subjects [GH89]. submatrices [KW85]. Subset [Bjø85, Ber80, Bjø86, GY85, HPT84]. success [AM89, ZB81]. sufficiency [BK86a, FP82]. sufficient [JS80, Jac85, Jac89, KKL88]. sum [DHK87, Hay82a]. sums [Cha86, EL86, Lo89]. super [Kal85]. superior [Bjø85]. superiority [SB86]. superpopulations [KW89]. support [HP85, HLT89]. surface [AP83, Man89]. surfaces [CP84, Chi86, Kos89]. Survey [Ovc88, Jos88, Lon81, Sär82]. surveys [SK89, Sin85]. Survival [PR89, Ara87, Ara89, Che89, GSV85, GSvR88, Gil86b, Zah85]. Suzuki [Mie82]. swapping [DR82]. Switzerland [Ano82d]. symbol [Kur88]. symbols [Lon84]. Symmetric [But85, CM86b, GG86c, Kap88, KP87a, Pil86, Pre82, SS82a, SP81, Wie88a]. symmetrical [Gup89, NB85, PR88]. symmetry [Chm81, CM88, NS85, NS87b]. system [AWW87b, Ahm83, Goi88]. Systematic [KW89, CW83]. systems [Bil85, De 87, Rok84].

table [Gil86b]. tables [Ash81, MB85, NS86]. tail [HG86]. tax [Gas82]. technique [DR82, LS81]. techniques [Ecc80, Gil86b, PL89]. Technische [Ano82d]. ternary [Don87]. Terry [CS84a]. test [Alb85, AU89, Bar82, BS82, BE87, BB82, BK89, BR85, BA82, Ebr86, Fei85, GW88, GG86b, Hab81, Hay82b, HM88, Jac88, JM89, JS81a, Kha81, Nab86, PSSZ87, PC81, SD85b, Sid82b, Stu88, Tik80, Vís86, Wie88a, WK87]. testimony [Zab88]. Testing [BA82, Cam84, CS84b, CW89, GR85, Khu89, KG89, Bau89, BFPR84, CR82, CH87b, GH89, Gil85, HSB86, Hay86, HX87, RB85, Sen84b, Sri87b, WR87]. Tests [Chi86, Gut87, AK88, Ang82, BN86, CG83, CP84, Chm81, CP80, CMS82, CM88, CH87c, DN86a, EL86, GCG88, GG86c, Gov84, GM85, HM83, Hod88, Hor88, Hor89, JT85, Jur80, KK89, KS82, LaR86, MH84, MP88, Moo84, PS87, RG81, Sen84a, SS83b, Sha89, Sko88, Sko89, SS82b, Woo84, Wu87]. th [Pra85, ZdK81]. their [BFPR84, Gup87]. them [SD86]. theorem [DK89, GW80, Wie89]. theorems [SP89]. Theory [SAM84, BJ88, Che83, DL80, Gar83, NS85, NS87b, Puk83, Shi82, Wat83, Weh86]. there [Car82, Mie87]. third [Hud83]. third-order [Hud83]. thirty [DL84b]. Thompson [Ano89d, Ram82, Rao83]. Three [Muk87, Gup85a, JT85, Roe87, Sch84, Seb84, Seb85a, Sko88, Sko89]. three-factorial [Roe87]. Three-stage [Muk87]. three-way [Gup85a]. ties [Bea83, GG81]. Tiku [Tan85]. time [Ara87, Ara89, Cha89b, CS80, GSV85, Sri87b, Wat82]. time-consuming [Sri87b]. times [Gov84]. total [Bar86]. Toulouse [Ano82k]. trades [Hwa86]. transformation [Alb85, SYT81, Tay89]. transformations [CP80]. transversal [dL87]. treatment [BT85, Hec87, Jac88, Kha80, Not85]. treatment-control [Not85]. treatments [HM88, Jac88, JM89, Maj86, SNS85, Stu88]. trend [Bel81]. trial [Aic83b]. trials [Gou83, ZB81]. triangular [Sko88, Sko89]. trimmed [JSV87, Sho89]. triple [Bil85, CHT84]. triples [Bil85]. true [Gou83, HG86]. truncated [BK86b, BM87, Jos82, Zel88]. truncation [BlB85, Hod88]. tumor [BD88]. tuples [Rok84]. Two [BN86, Bha84b, HM83, Hud83, Kub89, Mar87, Rüs85, Agr87, AC89, BGH84, BM84, BlB85, Bar82, BR86, BJV89, CW83, CP80, Das89, De 87, GK83a, GG86b, GK83b, GKZ85, Gup85a, Hay82a, Hay82b, Hec87, Hud85, HJ81, Kag82, Kag86, Kos89, Lin88, Lon84, Maj86, MH84, Moh81, Mor84, Muk84, Nai82, Not85, Pei88, RG81, RF84, San81, Seb85b, SNS85, SS85b, Sko88, Sko89, WR87]. two- [Gup85a]. two-associate [Kag82, Kag86, Moh81]. two-class [Agr87]. two-level [Mor84]. Two-phase [HM83]. two-sample [BJV89, HJ81, RG81, RF84, Sko88, Sko89]. two-sided [CP80]. Two-stage [Kub89, Mar87, BR86, Muk84]. two-unit [De 87]. two-way [CW83, Das89, GG86b, MH84, Not85]. type [AU89, Ang82, BN86, BH89, CS84a, CS80, EL86, Fed89, Gil85, Jac85, Jac89, LPY86, MB87, Moo84, NS87a, RMW85, SAM84, Wu87]. types [But85, VM83].

Uchida [Mie82]. UMVU [Kha84]. unavailability [BT87, Gho82]. unbalanced [CD83, Das89]. Unbiased [CK84b, BK81, BM84, BlB85, Cha88b, Sah80]. unbiasedness [Pra85, Rao81]. unbounded [HR88]. Uncertain [DK88]. uncertainty [DeG88, Doc88, Fin88, WS88]. unequal [AM89, LJ87, Muk84]. Unified [Kha80, DL80]. uniform [Bos86, CA85, DHK87, Mac83, Mór83, Pra85, RJDT83, Roy88, Ruy89]. uniformity [Gas82]. uniformly [Gut87]. Union [Ovc88]. Uniqueness [VD89, Dur89, Kal87, Tse86]. unit [De 87, Lau88]. units [GLP84, HRS88]. univariate [Wie88a]. universal [SM82]. universally [Hed84, Kra85, Yeh88]. Universite [Ano82k]. unknown [BK83, HM83, Muk84]. unstable [Cha89b]. upper [HG86]. upper-tail [HG86]. Use [MH84, SS85a]. used [BE88]. using [Alb85, AU89, CW83, CS80, Day85, FS89, GA85, Gov80, MW87, PL89, Seb82].

VA [Roh88]. vague [Goo87a]. valid [Fun82]. validity [BP88]. value [HSB86, Loh85]. values [Aic83a, Day85, Lo89]. variable [CP80, PL89]. variables [CC89, Cha86, GR85, GH88, Mus87, Rou88]. Variance [Hud85, BM84, BlB85, Cal86, CS89, CD83, Cli89, Das89, Flu89b, GD88, GLP84, GS87, GK83b, GKZ85, Gno87, Gup87, HM83, Kha82, Kos89, KC81, Muk82, Pei88, Sah80]. variances [BA82]. variates [RJDT83]. varieties [Gaf82]. various [Doc88, Sri86, VM83]. varying [CA83]. Vector [MW87, BK86a, Cha88b, Che83]. vectors [BT85]. versus [BK81, Jos88]. vertices [Fra80]. VI [But85, Kuw87]. via [Lin88, Mau85]. view [DeG88]. viewed [Pou87]. Vol [Flu89b]. Volume [Ano80b, Ano81a, Ano82a, Ano83c, Ano84b, Ano84a, Ano85c, Ano87d, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c]. volumes [Ano84f, Ano88h]. VPI&SU [Roh88]. vs [Goo89b, Rao83].

wald [Kha81]. walk [SS83a, SK85]. walks [Nie86]. way [CD83, CW83, Das89, GG86b, Gup85a, MH84, Mik89, Not85, SS82b, Wie88b]. Weak [CR82, GV88, Har88, Geo87, Wu87]. Weakly [AT80]. Wednesday [Ano82d]. Weibull [AWW87a, Sin87]. weighing [JWM83, KC83, KF84, Mas88, MR85, SS86a, Sch87b, SS83d]. weighted [Hay82a]. weighting [Gas88]. well [Mar87]. which [Puk80]. width [CS88, GSV85, GV89, Mes87]. Wilk [AU89]. winning [Gup84]. wise [Muk81]. within [Hed84, aS89]. within-block [aS89]. without [MH84, Mik89]. Wolfowitz [Wie89]. word [Rot86]. Worst [Tuz82]. Worst-case [Tuz82].

x [Mie82]. Xanthan [PK83].

York [Ano80c, Flu89b, Mie82]. Yoshino [Mie82]. Youden [Ash81]. young [Ano86c].

Zürich [Ano82d].

References

Agarwal:1986:HDL

[AA86] A. K. Agarwal and G. E. Andrews. Hook differences and lattice paths. Journal of Statistical Planning and Inference, 14(1): 5–14, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900042.

Arnold:1984:MPO

[ABGZ84] B. C. Arnold, A. Becker, U. Gather, and H. Zahedi. On the Markov property of order statistics. Journal of Statistical Planning and Inference, 9(2):147–154, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900156.

Alvo:1989:SDE

[AC89] M. Alvo and P. Cabilio. Sampling designs for the estimation of the difference between two means of a bivariate normal. Journal of Statistical Planning and Inference, 23(3):353– 369, November 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900785.

Amahia:1989:ENE

[ACR89] G. N. Amahia, Y. P. Chaubey, and T. J. Rao. Efficiency of a new estimator in PPS sampling for multiple characteristics. Journal of Statistical Planning and Inference, 21(1):75– 84, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900207.

Anderson:1985:CCE

[AE85] D. A. Anderson and J. A. Eccleston. On the construction of a class of efficient row-column designs. Journal of Statistical Planning and Inference, 11(1):131–134, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900333.

Athreya:1984:LLN

[AGLS84] Krishna B. Athreya, Malay Ghosh, Leone Y. Low, and Pranab K. Sen. Laws of large numbers for bootstrapped U-statistics. Journal of Statistical Planning and Inference, 9(2):185–194, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375884900193.

Agrawal:1987:CTC

[Agr87] H. C. Agrawal. Construction of two-class cyclic partially balanced incomplete block designs. Journal of Statistical Planning and Inference, 16(??):127–132, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900619.

Aubuchon:1984:NEI

[AH84] James C. Aubuchon and Thomas P. Hettmansperger. A note on the estimation of the integral of f^{2}(x). Journal of Statistical Planning and Inference, 9(3):321–331, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900077.

Ahmad:1983:NAE

[Ahm83] Ibrahim A. Ahmad. On the normal approximation of an estimate of the mean residual life of a multicomponent system. Journal of Statistical Planning and Inference, 7(3):195–207, March 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375883900149.

Aickin:1983:SV

[Aic83a] M. Aickin. Serial P -values. Journal of Statistical Planning and Inference, 7(3):243–256, March 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900186.

Aickin:1983:SLT

[Aic83b] Mikel Aickin. Some large trial properties of minimum likelihood allocation. Journal of Statistical Planning and Inference, 8(1):11– 20, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375883900575.

Ali:1987:OSL

[AK87] M. Masoom Ali and A. H. Khan. On order statistics from the loglogistic distribution. Journal of Statistical Planning and Inference, 17(??):103–108, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901042.

Akritas:1988:LBE

[AK88] Michael G. Akritas and Stavros Kourouklis. Local Bahadur efficiency of score tests. Journal of Statistical Planning and Inference, 19(2):187–199, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900729.

Ahmad:1984:FMR

[AL84] Ibrahim A. Ahmad and Pi-Erh Lin. Fitting a multiple regression function. Journal of Statistical Planning and Inference, 9(2): 163–176, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588490017X.

Albers:1985:ERT

[Alb85] Willem Albers. An exact rank test for scale under normality using Helmert’s transformation. Journal of Statistical Planning and Inference, 12(??):331–346, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900813.

Abughalous:1989:SLS

[AM89] Mansour M. Abughalous and Klaus J. Miescke. On selecting the largest success probability under unequal sample sizes. Journal of Statistical Planning and Inference, 21(1):53–68, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900189.

Amoh:1986:ARL

[Amo86] R. K. Amoh. Asymptotic results for a linear discriminant function estimated from a mixture of inverse Gaussian populations. Journal of Statistical Planning and Inference, 14(2–3):233– 243, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901606.

Anderson:1988:SDF

[And88] Oliver D. Anderson. Statistical discussion forum. Journal of Statistical Planning and Inference, 20(1):129, ???? 1988. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375888900894.

Angus:1982:GFT

[Ang82] John E. Angus. Goodness-of-fit tests for exponentiality based on a loss-of-memory type functional equation. Journal of Statistical Planning and Inference, 6(3):241–251, ???? 1982. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375882900295.

Anonymous:1980:A

[Ano80a] Anonymous. Announcement. Journal of Statistical Planning and Inference, 4(4):327, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588090018X.

Anonymous:1980:AIV

[Ano80b] Anonymous. Author index: Volume 4 (1980). Journal of Statistical Planning and Inference, 4(4):403–404, ???? 1980. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375880900257.

Anonymous:1980:BRA

[Ano80c] Anonymous. Book review: Applied Cybernetics: A. Ghosal (Gordon and Breach Science Publishers, New York, 1978). Journal of Statistical Planning and Inference, 4(3):325, ???? 1980. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375880900178.

Anonymous:1980:EB

[Ano80d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 4(1):ii, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375880900269.

Anonymous:1980:Pa

[Ano80e] Anonymous. Pages 1-114 (1980). Journal of Statistical Planning and Inference, 4(1):??, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1980:Pb

[Ano80f] Anonymous. Pages 115-212 (1980). Journal of Statistical Planning and Inference, 4(2):??, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1980:Pc

[Ano80g] Anonymous. Pages 213-326 (1980). Journal of Statistical Planning and Inference, 4(3):??, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1980:Pd

[Ano80h] Anonymous. Pages 327-404 (1980). Journal of Statistical Planning and Inference, 4(4):??, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1981:AIV

[Ano81a] Anonymous. Author index: Volume 5 (1981). Journal of Statistical Planning and Inference, 5(4):417–419, ???? 1981. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375881900458.

Anonymous:1981:EB

[Ano81b] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 5(1):ii, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588190046X.

Anonymous:1981:Pa

[Ano81c] Anonymous. Pages 1-106 (1981). Journal of Statistical Planning and Inference, 5(1):??, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1981:Pb

[Ano81d] Anonymous. Pages 107-212 (1981). Journal of Statistical Planning and Inference, 5(2):??, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1981:Pc

[Ano81e] Anonymous. Pages 213-308 (1981). Journal of Statistical Planning and Inference, 5(3):??, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1981:Pd

[Ano81f] Anonymous. Pages 309-420 (1981). Journal of Statistical Planning and Inference, 5(4):??, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1982:AIV

[Ano82a] Anonymous. Author index: Volume 6 (1982). Journal of Statistical Planning and Inference, 6(4):422–424, ???? 1982. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375882900088.

Anonymous:1982:EBa

[Ano82b] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 6(1):obc, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900611.

Anonymous:1982:EBb

[Ano82c] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 7(1):ifc, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375882900131.

Anonymous:1982:FIC

[Ano82d] Anonymous. Fourth international conference on mathematical modelling: Eidgenössische Technische Hochschule Zürich (Honggerberg Campus) Zürich, Switzerland, Monday, 15 August–Wednesday, 17 August, 1983. Journal of Statistical Planning and Inference, 7 (1):107–108, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900246.

Anonymous:1982:Pa

[Ano82e] Anonymous. Pages 1-104 (1982). Journal of Statistical Planning and Inference, 6(1):??, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1982:PO

[Ano82f] Anonymous. Pages 1-108 (October 1982). Journal of Statistical Planning and Inference, 7(1):??, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1982:Pb

[Ano82g] Anonymous. Pages 105-204 (1982). Journal of Statistical Planning and Inference, 6(2):??, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1982:PDa

[Ano82h] Anonymous. Pages 109-192 (December 1982). Journal of Statistical Planning and Inference, 7(2):??, December 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1982:Pc

[Ano82i] Anonymous. Pages 205-308 (1982). Journal of Statistical Planning and Inference, 6(3):??, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1982:Pd

[Ano82j] Anonymous. Pages 309-424 (1982). Journal of Statistical Planning and Inference, 6(4):??, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1982:TUP

[Ano82k] Anonymous. Toulouse 82 universite Paul Sabatier. Journal of Statistical Planning and Inference, 6(4):421, ???? 1982. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375882900076.

Anonymous:1983:A

[Ano83a] Anonymous. Announcement. Journal of Statistical Planning and Inference, 8(3):389–393, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588390054X.

Anonymous:1983:AI

[Ano83b] Anonymous. Author index. Journal of Statistical Planning and Inference, 7(4):419–421, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900125.

Anonymous:1983:AIV

[Ano83c] Anonymous. Author index: Volume 8 (1983). Journal of Statistical Planning and Inference, 8(3):394–396, December 1983. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375883900551.

Anonymous:1983:EB

[Ano83d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 8(1):ibc, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900678.

Anonymous:1983:EN

[Ano83e] Anonymous. Editorial notice. Journal of Statistical Planning and Inference, 7(3):193, March 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900137.

Anonymous:1983:PS

[Ano83f] Anonymous. Pages 1-128 (September 1983). Journal of Statistical Planning and Inference, 8(1):??, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1983:PO

[Ano83g] Anonymous. Pages 129-243 (October 1983). Journal of Statistical Planning and Inference, 8(2):??, October 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1983:PM

[Ano83h] Anonymous. Pages 193-294 (March 1983). Journal of Statistical Planning and Inference, 7(3):??, March 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1983:PD

[Ano83i] Anonymous. Pages 245-396 (December 1983). Journal of Statistical Planning and Inference, 8(3):??, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1983:PJ

[Ano83j] Anonymous. Pages 295-421 (June 1983). Journal of Statistical Planning and Inference, 7(4):??, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1984:AIVb

[Ano84a] Anonymous. Author index: Volume 10 (1984). Journal of Statistical Planning and Inference, 10(3):397–399, October 1984. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375884900636.

Anonymous:1984:AIVa

[Ano84b] Anonymous. Author index: Volume 9 (1984). Journal of Statistical Planning and Inference, 9(3):397–399, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900144.

Anonymous:1984:EBa

[Ano84c] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 9(1):iv, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900399.

Anonymous:1984:EBb

[Ano84d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 10(1):iv, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900260.

Anonymous:1984:F

[Ano84e] Anonymous. Forward. Journal of Statistical Planning and Inference, 9(3):271–272, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900028.

Anonymous:1984:MIV

[Ano84f] Anonymous. Master index to volumes 1-10. Journal of Statistical Planning and Inference, 10(3):400–415, October 1984. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375884900648.

Anonymous:1984:PJb

[Ano84g] Anonymous. Pages 1-130 (July 1984). Journal of Statistical Planning and Inference, 10(1):??, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1984:PJa

[Ano84h] Anonymous. Pages 1-146 (January 1984). Journal of Statistical Planning and Inference, 9(1):??, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1984:PA

[Ano84i] Anonymous. Pages 131-264 (August 1984). Journal of Statistical Planning and Inference, 10(2):??, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1984:PMa

[Ano84j] Anonymous. Pages 147-268 (March 1984). Journal of Statistical Planning and Inference, 9(2):??, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1984:PO

[Ano84k] Anonymous. Pages 265-415 (October 1984). Journal of Statistical Planning and Inference, 10(3):??, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1984:PMb

[Ano84l] Anonymous. Pages 269-399 (May 1984). Journal of Statistical Planning and Inference, 9(3):??, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1985:ARS

[Ano85a] Anonymous. Acknowledgement of referee’s services. Journal of Statistical Planning and Inference, 12(??):403, ???? 1985. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375885900904.

Anonymous:1985:AI

[Ano85b] Anonymous. Author index. Journal of Statistical Planning and Inference, 12(??):404–406, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900916.

Anonymous:1985:AIV

[Ano85c] Anonymous. Author index: Volume 11 (1985). Journal of Statistical Planning and Inference, 11(3):411–412, April 1985. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375885900461.

Anonymous:1985:D

[Ano85d] Anonymous. Dedication. Journal of Statistical Planning and Inference, 11(1):vii, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900151.

Anonymous:1985:EBa

[Ano85e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 11(1):ii, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588590014X.

Anonymous:1985:EBb

[Ano85f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 12(??):i, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900473.

Anonymous:1985:PJ

[Ano85g] Anonymous. Pages 1-134 (January 1985). Journal of Statistical Planning and Inference, 11(1):??, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1985:P

[Ano85h] Anonymous. Pages 1-406 (1985). Journal of Statistical Planning and Inference, 12(??):1–406, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1985:PF

[Ano85i] Anonymous. Pages 135-266 (February 1985). Journal of Statistical Planning and Inference, 11(2):??, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1985:PA

[Ano85j] Anonymous. Pages 267-412 (April 1985). Journal of Statistical Planning and Inference, 11(3):??, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1985:PJK

[Ano85k] Anonymous. Publications of Jack Kiefer: Papers (published). Journal of Statistical Planning and Inference, 11(1):viii–xiii, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900163.

Anonymous:1986:AIa

[Ano86a] Anonymous. Author index. Journal of Statistical Planning and Inference, 13(??):407–408, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901527.

Anonymous:1986:AIb

[Ano86b] Anonymous. Author index. Journal of Statistical Planning and Inference, 14(2–3):413–414, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901795.

Anonymous:1986:CYS

[Ano86c] Anonymous. Competitions for young statisticians from developing countries 1987/1989. Journal of Statistical Planning and Inference, 14(1):153, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886900182.

Anonymous:1986:EBa

[Ano86d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 13(??):ifc, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901138.

Anonymous:1986:EBb

[Ano86e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 14(1):ii, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886900017.

Anonymous:1986:ISO

[Ano86f] Anonymous. ISI section on official statistics. Journal of Statistical Planning and Inference, 14(1):154, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886900194.

Anonymous:1986:PM

[Ano86g] Anonymous. Pages 1-154 (May 1986). Journal of Statistical Planning and Inference, 14(1):??, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1986:Pa

[Ano86h] Anonymous. Pages 1-408 (1986). Journal of Statistical Planning and Inference, 13(??):1–408, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1986:Pb

[Ano86i] Anonymous. Pages 155-414 (1986). Journal of Statistical Planning and Inference, 14(2–3):??, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1986:SDF

[Ano86j] Anonymous. Statistical discussion forum. Journal of Statistical Planning and Inference, 14(2–3):411–412, ???? 1986. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375886901783.

Anonymous:1987:ARS

[Ano87a] Anonymous. Acknowledgment of referee’s services. Journal of Statistical Planning and Inference, 17(??):135–136, ???? 1987. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375887901066.

Anonymous:1987:A

[Ano87b] Anonymous. Announcement. Journal of Statistical Planning and Inference, 17(??):409, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887901327.

Anonymous:1987:AI

[Ano87c] Anonymous. Author index. Journal of Statistical Planning and Inference, 17(??):411–412, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887901339.

Anonymous:1987:AIV

[Ano87d] Anonymous. Author index volume 16 (1987). Journal of Statistical Planning and Inference, 16(??):431–432, ???? 1987. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375887900954.

Anonymous:1987:CPC

[Ano87e] Anonymous. Call for papers: Compstat’88. Journal of Statistical Planning and Inference, 17(??):277, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887901194.

Anonymous:1987:EBa

[Ano87f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 16(??):ifc, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900486.

Anonymous:1987:EBb

[Ano87g] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 17(??):ifc, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900966.

Anonymous:1987:P

[Ano87h] Anonymous. Pages 1-412 (1987). Journal of Statistical Planning and Inference, 17(??):1–412, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1988:ARS

[Ano88a] Anonymous. Acknowledgement of referees’ services. Journal of Statistical Planning and Inference, 20(2):237–238, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888901280.

Anonymous:1988:AIVa

[Ano88b] Anonymous. Author index: Volume 18 (1988). Journal of Statistical Planning and Inference, 18(3):405–406, March 1988. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375888901176.

Anonymous:1988:AIVb

[Ano88c] Anonymous. Author index: Volume 19 (1988). Journal of Statistical Planning and Inference, 19(3):401–402, July 1988. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837588890047X.

Anonymous:1988:AIVc

[Ano88d] Anonymous. Author index: Volume 20 (1988). Journal of Statistical Planning and Inference, 20(3):453–454, ???? 1988. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375888901024.

Anonymous:1988:EBa

[Ano88e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 18(1):ii, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900110.

Anonymous:1988:EBb

[Ano88f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 19(1):ii, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900481.

Anonymous:1988:EBc

[Ano88g] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 20(1):ii, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900791.

Anonymous:1988:MIV

[Ano88h] Anonymous. Master index to volumes 11-20. Journal of Statistical Planning and Inference, 20(3):455–471, ???? 1988. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375888901036.

Anonymous:1988:Pc

[Ano88i] Anonymous. Pages 1-130 (1988). Journal of Statistical Planning and Inference, 20(1):??, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1988:PJa

[Ano88j] Anonymous. Pages 1-134 (January 1988). Journal of Statistical Planning and Inference, 18(1):??, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1988:Pa

[Ano88k] Anonymous. Pages 1-136 (1988). Journal of Statistical Planning and Inference, 19(1):??, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1988:Pd

[Ano88l] Anonymous. Pages 131-238 (1988). Journal of Statistical Planning and Inference, 20(2):??, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1988:PF

[Ano88m] Anonymous. Pages 135-266 (February 1988). Journal of Statistical Planning and Inference, 18(2):??, February 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1988:Pb

[Ano88n] Anonymous. Pages 137-268 (1988). Journal of Statistical Planning and Inference, 19(2):??, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1988:Pe

[Ano88o] Anonymous. Pages 239-472 (1988). Journal of Statistical Planning and Inference, 20(3):??, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1988:PM

[Ano88p] Anonymous. Pages 267-406 (March 1988). Journal of Statistical Planning and Inference, 18(3):??, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1988:PJb

[Ano88q] Anonymous. Pages 269-402 (July 1988). Journal of Statistical Planning and Inference, 19(3):??, July 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1989:AIVa

[Ano89a] Anonymous. Author index: Volume 21 (1989). Journal of Statistical Planning and Inference, 21(3):411–412, March 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837588990058X.

Anonymous:1989:AIVb

[Ano89b] Anonymous. Author index: Volume 22 (1989). Journal of Statistical Planning and Inference, 22(3):405–406, July 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375889901055.

Anonymous:1989:AIVc

[Ano89c] Anonymous. Author index volume 23 (1989). Journal of Statistical Planning and Inference, 23(3):405–406, November 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375889900839.

Anonymous:1989:DPP

[Ano89d] Anonymous. Discussion of the paper by Professors Godambe and Thompson. Journal of Statistical Planning and Inference, 22(2): 153–172, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889901079. See [GT89].

Anonymous:1989:EBa

[Ano89e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 21(1):ii, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900128.

Anonymous:1989:EBb

[Ano89f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 22(1):ii, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900591.

Anonymous:1989:EBc

[Ano89g] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 23(1):ii, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900347.

Anonymous:1989:PS

[Ano89h] Anonymous. Pages 1-132 (September 1989). Journal of Statistical Planning and Inference, 23(1):??, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1989:PMb

[Ano89i] Anonymous. Pages 1-136 (May 1989). Journal of Statistical Planning and Inference, 22(1):??, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1989:PJa

[Ano89j] Anonymous. Pages 1-138 (January 1989). Journal of Statistical Planning and Inference, 21(1):??, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1989:PO

[Ano89k] Anonymous. Pages 133-266 (October 1989). Journal of Statistical Planning and Inference, 23(2):??, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1989:PJb

[Ano89l] Anonymous. Pages 137-270 (June 1989). Journal of Statistical Planning and Inference, 22(2):??, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1989:PF

[Ano89m] Anonymous. Pages 139-276 (February 1989). Journal of Statistical Planning and Inference, 21(2):??, February 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1989:PN

[Ano89n] Anonymous. Pages 267-406 (November 1989). Journal of Statistical Planning and Inference, 23(3):??, November 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1989:PJc

[Ano89o] Anonymous. Pages 271-406 (July 1989). Journal of Statistical Planning and Inference, 22(3):??, July 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1989:PMa

[Ano89p] Anonymous. Pages 277-412 (March 1989). Journal of Statistical Planning and Inference, 21(3):??, March 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1989:SDF

[Ano89q] Anonymous. Statistical discussion forum. Journal of Statistical Planning and Inference, 21(3):407–409, March 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375889900578.

Adhikary:1983:GDR

[AP83] Basudeb Adhikary and Rajendranath Panda. On group divisible response surface (GDRS) designs. Journal of Statistical Planning and Inference, 7(4):387–405, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900095.

Aras:1987:SEM

[Ara87] Girish Aras. Sequential estimation of the mean exponential survival time under random censoring. Journal of Statistical Planning and Inference, 16(??):147–158, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887900644.

Aras:1989:SOS

[Ara89] Girish Aras. Second order sequential estimation of the mean exponential survival time under random censoring. Journal of Statistical Planning and Inference, 21(1):3–17, January 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375889900141.

Aigner:1985:DDL

[AS85] M. Aigner and M. Schughart. Determining defectives in a linear order. Journal of Statistical Planning and Inference, 12(??): 359–368, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900849.

Shiraishi:1989:AES

[aS89] Taka aki Shiraishi. Asymptotic equivalence of statistical inference based on aligned ranks and on within-block ranks. Journal of Statistical Planning and Inference, 21(3):347–364, March 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900529.

Ash:1981:GYD

[Ash81] Arlene Ash. Generalized Youden designs: construction and tables. Journal of Statistical Planning and Inference, 5(1):1– 25, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900471.

Anderson:1980:WRI

[AT80] Donald A. Anderson and Ann M. Thomas. Weakly resolvable IV.3 search designs for the p^{n} factorial experiment. Journal of Statistical Planning and Inference, 4(3):299–312, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375880900142.

Ali:1989:SWT

[AU89] M. Masoom Ali and Dale Umbach. A Shapiro-wilk type goodnessof-fit test using a few order statistics. Journal of Statistical Planning and Inference, 22(2):251–261, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889901158.

Abdel-Wahid:1987:ABE

[AWW87a] A. A. Abdel-Wahid and A. Winterbottom. Approximate Bayesian estimates for the Weibull reliability function and hazard rate from censored data. Journal of Statistical Planning and Inference, 16 (??):277–283, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900802.

Abdel-Wahid:1987:ASR

[AWW87b] A. R. Abdel-Wahid and A. Winterbottom. The approximation of system reliability posterior distributions. Journal of Statistical Planning and Inference, 16(??):267–275, ???? 1987. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375887900796.

Bhattacharjee:1982:TEV

[BA82] G. P. Bhattacharjee and S. S. Alam. Testing equality of variances after a preliminary test of significance for correlation coefficient. Journal of Statistical Planning and Inference, 6(1):17– 28, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900520.

Babu:1988:NCC

[Bab88] Gutti Jogesh Babu. A note on comparison of conditional means. Journal of Statistical Planning and Inference, 19(2):253–259, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900778.

Babbel:1989:IPS

[Bab89] Beate Babbel. Invariance principles for U-statistics and von Mises functionals. Journal of Statistical Planning and Inference, 22(3): 337–354, July 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900992.

Bailey:1988:SLS

[Bai88] R. A. Bailey. Semi-Latin squares. Journal of Statistical Planning and Inference, 18(3):299–312, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888901073.

Baksalary:1989:RCL

[Bak89] Jerzy K. Baksalary. A rank characterization of linear models with nuisance parameters and its application to block designs. Journal of Statistical Planning and Inference, 22(2):173– 179, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889901080.

Barker:1982:AOB

[Bar82] L. E. Barker. Allocating observations between two normal populations to maximize the half slope of the F -test. Journal of Statistical Planning and Inference, 6(2):115–118, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900787.

Bartlett:1986:ETC

[Bar86] Roy F. Bartlett. Estimating the total of a continuous population. Journal of Statistical Planning and Inference, 13(??):51– 66, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901199.

Basu:1987:P

[Bas87] Asit P. Basu. Preface. Journal of Statistical Planning and Inference, 16(??):iv, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900504.

Bauer:1989:SEP

[Bau89] P. Bauer. A sequential elimination procedure for choosing the best population(s) based on multiple testing. Journal of Statistical Planning and Inference, 21(2):245–252, February 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900086.

Bemis:1982:EST

[BB82] Kerry G. Bemis and Vasant P. Bhapkar. On the equivalence of some test criteria based on BAN estimators for the multivariate exponential family. Journal of Statistical Planning and Inference, 6(3):277–286, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900313.

Bhandari:1987:SML

[BB87] Subir Kumar Bhandari and Arup Bose. On selecting the most likely event. Journal of Statistical Planning and Inference, 17(??): 227–240, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887901157.

Berger:1983:CCS

[BD83] James O. Berger and Dipak K. Dey. Combining coordinates in simultaneous estimation of normal means. Journal of Statistical Planning and Inference, 8(2):143–160, October 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900344.

Bartoszynski:1988:APT

[BD88] R. Bartoszyński and S. M. Dynin. Asymptotic properties of the tumor growth curve estimator. Journal of Statistical Planning and Inference, 18(1):45–56, January 1988. CODEN JSPIDN.

ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900158.

Basu:1982:BAC

[BdBP82] D. Basu and Carlos A. de B. Pereira. On the Bayesian analysis of categorical data: the problem of nonresponse. Journal of Statistical Planning and Inference, 6(4):345–362, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900040.

Baksalary:1980:SMC

[BDK80] J. K. Baksalary, A. Dobek, and R. Kala. Some methods for constructing efficiency-balanced block designs. Journal of Statistical Planning and Inference, 4(1):25–32, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375880900300.

Basu:1987:BAL

[BE87] A. P. Basu and N. Ebrahimi. On a bivariate accelerated life test. Journal of Statistical Planning and Inference, 16(??):297– 304, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887900826.

Basu:1988:MHN

[BE88] Asit P. Basu and Nader Ebrahimi. Multivariate harmonic new better than used in expectation distributions. Journal of Statistical Planning and Inference, 20(2):181–190, ???? 1988. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375888901218.

Beaver:1983:LLM

[Bea83] Robert J. Beaver. Log linear models for multivariate paired comparison experiments with ties. Journal of Statistical Planning and Inference, 7(3):209–218, March 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900150.

Brewer:1984:PMP

[BEH84] K. R. W. Brewer, L. J. Early, and Muhammad Hanif. Poisson, modified Poisson and collocated sampling. Journal of Statistical Planning and Inference, 10(1):15–30, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900296.

Bellhouse:1981:SSP

[Bel81] D. R. Bellhouse. Spatial sampling in the presence of a trend. Journal of Statistical Planning and Inference, 5(4):365–375, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375881900379.

Bell:1984:IGF

[Bel84] C. B. Bell. Inference for goodness-of-fit problems with nuisance parameters: (applications to signal detection). Journal of Statistical Planning and Inference, 9(3):273–284, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588490003X.

Bellhouse:1989:OEL

[Bel89a] D. R. Bellhouse. Optimal estimation of linear functions of finite population means in rotation sampling. Journal of Statistical Planning and Inference, 21(1):69–74, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900190.

Bellout:1989:ORE

[Bel89b] Djamel Bellout. Order restricted estimation of distributions with censored data. Journal of Statistical Planning and Inference, 21(1): 27–39, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900165.

Berger:1980:MSS

[Ber80] Roger L. Berger. Minimax subset selection for the multinomial distribution. Journal of Statistical Planning and Inference, 4(4): 391–402, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375880900245.

Bush:1984:NCD

[BFPR84] K. A. Bush, W. T. Federer, H. Pesotan, and D. Raghavarao. New combinatorial designs and their applications to group testing. Journal of Statistical Planning and Inference, 10(3):335– 343, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900582.

Bailey:1984:BDB

[BGH84] R. A. Bailey, D. C. Goldrei, and D. F. Holt. Block designs with block size two. Journal of Statistical Planning and Inference, 10 (2):257–263, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900776.

Burke:1989:LSP

[BH89] Murray D. Burke and Lajos Horvath. Large sample properties of kernel-type score function estimators. Journal of Statistical Planning and Inference, 22(3):307–321, July 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900979.

Bhapkar:1984:NPA

[Bha84a] Vasant P. Bhapkar. On nonparametric profile analysis of several multivariate samples. Journal of Statistical Planning and Inference, 9(3):285–296, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900041.

Bhave:1984:TCC

[Bha84b] S. V. Bhave. Two concepts of conditionality. Journal of Statistical Planning and Inference, 10(2):131–135, August 1984. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837588490065X.

Bhattacharya:1985:BRS

[Bha85] Rabi Bhattacharya. Book review: Seminar on stochastic processes, 1982: E. Çinlar, K. L. Chung, R. K. Getoor, Editors, Birkhäuser, Boston, 1983. Journal of Statistical Planning and Inference, 12 (??):133–135, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588590062X.

Bhattacharjee:1987:NRB

[Bha87] Manish C. Bhattacharjee. New results for the Brown–Proschan model of imperfect repair. Journal of Statistical Planning and Inference, 16(??):305–316, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900838.

Bhapkar:1989:CAS

[Bha89] Vasant P. Bhapkar. Conditioning on ancillary statistics and loss of information in the presence of nuisance parameters. Journal of Statistical Planning and Inference, 21(2):139–160, February 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900013.

Bose:1984:BOD

[BI84] R. C. Bose and H. K. Iyer. Balanced E-optimal designs of resolution III for the 2^{m} × 3^{n} series. Journal of Statistical Planning and Inference, 10(3):345–364, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900594.

Buonaccorsi:1986:ODR

[BI86] J. P. Buonaccorsi and H. K. Iyer. Optimal designs for ratios of linear combinations in the general linear model. Journal of Statistical Planning and Inference, 13(??):345–356, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886901448.

Billington:1985:MDT

[Bil85] Elizabeth J. Billington. On Mendelsohn and directed triple systems with repeated elements allowed in triples. Journal of Statistical Planning and Inference, 12(??):75–79, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900540.

Berg:1988:MBP

[BJ88] Sven Berg and Jerzy Jaworski. Modified binomial and Poisson distributions with applications in random mapping theory. Journal of Statistical Planning and Inference, 18(3):313–322, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888901085.

Bjornstad:1985:SSS

[Bjø85] Jan F. Bjørnstad. Subset selection of superior populations when the number of populations is large. Journal of Statistical Planning and Inference, 11(2):207–216, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900072.

Bjornstad:1986:ACS

[Bjø86] Jan F. Bjørnstad. Asymptotic consistency for subset selection procedures satisfying the P ^{∗}-condition. Journal of Statistical Planning and Inference, 13(??):319–335, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901424.

Boos:1989:RCD

[BJV89] D. Boos, P. Janssen, and N. Veraverbeke. Resampling from centered data in the two-sample problem. Journal of Statistical Planning and Inference, 21(3):327–345, March 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900517.

Baksalary:1981:SLS

[BK81] J. K. Baksalary and R. Kala. Simple least squares estimation versus best linear unbiased prediction. Journal of Statistical Planning and Inference, 5(2):147–151, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375881900240.

Bhattacharjee:1983:SBA

[BK83] Manish C. Bhattacharjee and Ranjit Kandar. Simple bounds on availability in a model with unknown life and repair distributions. Journal of Statistical Planning and Inference, 8(2):129– 142, October 1983. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375883900332.

Baksalary:1986:LSR

[BK86a] Jerzy K. Baksalary and Radoslaw Kala. Linear sufficiency with respect to a given vector of parametric functions. Journal of Statistical Planning and Inference, 14(2–3):331–338, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886901710.

Balakrishnan:1986:MOS

[BK86b] N. Balakrishnan and Subrahmaniam Kocherlakota. On the moments of order statistics from the doubly truncated logistic distribution. Journal of Statistical Planning and Inference, 13(??): 117–129, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901266.

Banerjee:1986:MCR

[BK86c] Shakti Banerjee and Sanpei Kageyama. A method of constructing regular block designs. Journal of Statistical Planning and Inference, 13(??):399–401, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901503.

Bergman:1989:FTS

[BK89] Bo Bergman and Bengt Klefsjö. A family of test statistics for detecting monotone mean residual life. Journal of Statistical Planning and Inference, 21(2):161–178, February 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900025.

Berger:1984:LLS

[BL84] Roger L. Berger and Naftali A. Langberg. Linear least squares estimates and nonlinear means. Journal of Statistical Planning and Inference, 10(3):277–288, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900545.

Bar-lev:1985:MVU

[BlB85] Shaul K. Bar-lev and Benzion Boukai. Minimum variance unbiased estimation for families of distributions involving two truncation parameters. Journal of Statistical Planning and Inference, 12(??):379–384, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900862.

Baksalary:1989:MSE

[BLT89] Jerzy K. Baksalary, Erkki P. Liski, and Götz Trenkler. Mean square error matrix improvements and admissibility of linear estimators. Journal of Statistical Planning and Inference, 23(3):313– 325, November 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588990075X.

Baksalary:1984:NUE

[BM84] Jerzy K. Baksalary and Anna Molińska. Nonnegative unbiased estimability of linear combinations of two variance components. Journal of Statistical Planning and Inference, 10(1): 1–8, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900272.

Baksalary:1986:CAL

[BM86a] Jerzy K. Baksalary and Augustyn Markiewicz. Characterizations of admissible linear estimators in restricted linear models. Journal of Statistical Planning and Inference, 13(??):395– 398, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901497.

Bohrer:1986:CSE

[BM86b] Robert Bohrer and Adam Martinsek. On combining Stein estimation problems: an adaptive rule under classical criteria. Journal of Statistical Planning and Inference, 14(2–3):301–309, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886901680.

Balakrishnan:1987:MOS

[BM87] N. Balakrishnan and H. J. Malik. Moments of order statistics from truncated log-logistic distribution. Journal of Statistical Planning and Inference, 17(??):251–267, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887901170.

Baksalary:1988:ALE

[BM88] Jerzy K. Baksalary and Augustyn Markiewicz. Admissible linear estimators in the general Gauss–Markov model. Journal of Statistical Planning and Inference, 19(3):349–359, July 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900420.

Bhattacharya:1989:CSMa

[BM89a] Rabi N. Bhattacharya and Mukul Majumdar. Controlled semi-Markov models — the discounted case. Journal of Statistical Planning and Inference, 21(3):365–381, March 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900530.

Bhattacharya:1989:CSMb

[BM89b] Rabi N. Bhattacharya and Mukul Majumdar. Controlled semi-Markov models under long-run average rewards. Journal of Statistical Planning and Inference, 22(2):223–242, June 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375889901134.

Behnen:1986:TSA

[BN86] Konrad Behnen and Georg Neuhaus. Two sample adaptive rank tests under type II censoring. Journal of Statistical Planning and Inference, 13(??):203–213, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901333.

Bohning:1985:NEP

[Böh85] Dankmar Böhning. Numerical estimation of a probability measure. Journal of Statistical Planning and Inference, 11(1):57– 69, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900242.

Borjas:1982:RRC

[Bor82] George J. Borjas. On regressing regression coefficients. Journal of Statistical Planning and Inference, 7(2):131–137, December 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900714. See comment [BP87].

Bose:1986:CNU

[Bos86] Arup Bose. Certain non-uniform rates of convergence to normality for martingale differences. Journal of Statistical Planning and Inference, 14(2–3):155–167, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901539.

Bourke:1981:ASM

[Bou81] P. D. Bourke. On the analysis of some multivariate randomized response designs for categorical data. Journal of Statistical Planning and Inference, 5(2):165–170, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375881900264.

Bhagwandas:1980:SNG

[BP80] Bhagwandas and J. S. Parihar. Some new group divisible designs. Journal of Statistical Planning and Inference, 4(3):321– 323, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375880900166.

Baksalary:1987:FCR

[BP87] Jerzy K. Baksalary and P. D. Puri. F3. A comment regarding “On regressing regression coefficients”. Journal of Statistical Planning and Inference, 16(??):133–135, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887900620.

See [Bor82].

Baksalary:1988:CVF

[BP88] Jerzy K. Baksalary and P. D. Puri. Criteria for the validity of Fisher’s condition for balanced block designs. Journal of Statistical Planning and Inference, 18(1):119–123, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900225.

Baksalary:1989:IPM

[BP89] Jerzy K. Baksalary and Pawel R. Pordzik. Inverse-partitioned-matrix method for the general Gauss–Markov model with linear restrictions. Journal of Statistical Planning and Inference, 23(2): 133–143, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900840.

Birkes:1987:SMA

[BPS87] David Birkes, Cliff Pereira, and Justus Seely. Scalar multiples of admissible linear estimators. Journal of Statistical Planning and Inference, 17(??):337–344, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901248.

Best:1985:LTN

[BR85] D. J. Best and J. C. W. Rayner. Lancaster’s test of normality. Journal of Statistical Planning and Inference, 12(??):395– 400, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900886.

Bellhouse:1986:EPE

[BR86] D. R. Bellhouse and J. N. K. Rao. On the efficiency of prediction estimators in two-stage sampling. Journal of Statistical Planning and Inference, 13(??):269–281, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901394.

Brouwer:1984:FMO

[Bro84] A. E. Brouwer. Four mols of order 10 with a hole of order 2. Journal of Statistical Planning and Inference, 10(2):203– 205, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900727.

Barker:1982:BET

[BS82] Lawrence Barker and J. Sethuraman. Bahadur efficiency of the F -test. Journal of Statistical Planning and Inference, 6(3):301– 304, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900349.

Bansal:1985:AEM

[BS85] M. L. Bansal and Ravindra Singh. An alternative estimator for multiple characteristics in PPS sampling. Journal of Statistical Planning and Inference, 11(3):313–320, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900369.

Bagchi:1989:OCR

[BS89] S. Bagchi and K. R. Shah. On the optimality of a class of rowcolumn designs. Journal of Statistical Planning and Inference, 23 (3):397–402, November 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900815.

Baksalary:1985:ECC

[BT85] Jerzy K. Baksalary and Zenon Tabis. Existence and constructions of connected block designs with given vectors of treatment replications and block sizes. Journal of Statistical Planning and Inference, 12(??):285–293, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900771.

Baksalary:1987:CRB

[BT87] Jerzy K. Baksalary and Zenon Tabis. Conditions for the robustness of block designs against the unavailability of data. Journal of Statistical Planning and Inference, 16(??):49–54, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375887900541.

Bush:1981:FRD

[Bus81] K. A. Bush. Families of regular designs based on geometries. Journal of Statistical Planning and Inference, 5(4):391– 398, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900409.

Butler:1985:SDT

[But85] N. T. Butler. Symmetric designs of types V and VI. Journal of Statistical Planning and Inference, 11(3):355–361, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900400.

Barlow:1987:BAI

[BZ87] Richard E. Barlow and Xiang Zhang. Bayesian analysis of inspection sampling procedures discussed by deming. Journal of Statistical Planning and Inference, 16(??):285–296, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375887900814.

Chaudhuri:1983:ODS

[CA83] Arijit Chaudhuri and Arun Kumar Adhikary. On optimality of double sampling strategies with varying probabilities. Journal of Statistical Planning and Inference, 8(3):257–265, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375883900447.

Chaudhuri:1985:SRA

[CA85] Arijit Chaudhuri and Arun Kumar Adhikary. Some results on admissibility and uniform admissibility in double sampling. Journal of Statistical Planning and Inference, 12(??):199–202, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900692.

Chaudhuri:1989:NHL

[CA89] Arijit Chaudhuri and Arun Kumar Adhikary. A note on handling linear randomized response. Journal of Statistical Planning and Inference, 22(2):263–264, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588990116X.

Calvin:1986:NCV

[Cal86] James A. Calvin. A new class of variance balanced designs. Journal of Statistical Planning and Inference, 14(2–3):251–254, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837588690162X.

Campbell:1984:TEP

[Cam84] Gregory Campbell. Testing equality of proportions with incomplete correlated data. Journal of Statistical Planning and Inference, 10 (3):311–321, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900569.

Carroll:1982:REC

[Car82] R. J. Carroll. Robust estimation in certain heteroscedastic linear models when there are many parameters. Journal of Statistical Planning and Inference, 7(1):1–12, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375882900143.

Carmody:1988:DMS

[Car88] T. J. Carmody. Diagnostics for multivariate smoothing splines. Journal of Statistical Planning and Inference, 19(2):171–186, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900717.

Casaglia:1989:SBD

[Cas89] Ivan Casaglia. Some block designs associated with quadrics of PG(4, q) and PG(5, q). Journal of Statistical Planning and Inference, 21(2):265–272, February 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900104.

Catchpole:1984:GIB

[Cat84] E. A. Catchpole. Generalized inverses in block designs. Journal of Statistical Planning and Inference, 10(2):195–202, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375884900715.

Cressie:1986:MGF

[CB86] Noel Cressie and Marinus Borkent. The moment generating function has its moments. Journal of Statistical Planning and Inference, 13(??):337–344, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901436.

Charnet:1988:DAP

[CB88] Reinaldo Charnet and Robert J. Beaver. Design and analysis of paired comparison experiments involving mixtures. Journal of Statistical Planning and Inference, 20(1):91–103, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900857.

Colbourn:1984:RCC

[CC84] Marlene J. Colbourn and Charles J. Colbourn. Recursive constructions for cyclic block designs. Journal of Statistical Planning and Inference, 10(1):97–103, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900351.

Calinski:1989:NAC

[CC89] Tadeusz Caliński and Henri Caussinus. A note on the analysis of covariance: efficiency of concomitant variables. Journal of Statistical Planning and Inference, 21(3):315–326, March 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900505.

Chadjiconstantinidis:1989:COF

[CCM89] Stathis Chadjiconstantinidis, Ching-Shui Cheng, and Chronis Moyssiadis. Construction of optimal fractional factorial resolution V designs with N ≡ 2 mod 16 observations. Journal of Statistical Planning and Inference, 23(2):153–161, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900864.

Chatterjee:1983:EVC

[CD83] Shoutir Kishore Chatterjee and Kalyan Das. Estimation of variance components in an unbalanced one-way classification. Journal of Statistical Planning and Inference, 8(1):27–41, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375883900599.

Capobianco:1982:CSG

[CF82] Michael Capobianco and Ove Frank. Comparison of statistical graph-size estimators. Journal of Statistical Planning and Inference, 6(1):87–97, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900593.

Chakravorti:1983:OIT

[CG83] S. R. Chakravorti and N. C. Giri. Optimum invariant tests on discriminant coefficients or means of multinormal population with additional information. Journal of Statistical Planning and Inference, 8(1):1–10, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900563.

Constantine:1982:CRM

[CH82] G. Constantine and A. Hedayat. A construction of repeated measurements designs with balance for residual effects. Journal of Statistical Planning and Inference, 6(2):153–164, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900842.

Constantine:1983:CDBa

[CH83a] G. M. Constantine and A. S. Hedayat. Complete designs with blocks of maximal multiplicity. Journal of Statistical Planning and Inference, 7(3):289–294, March 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900204.

Constantine:1983:CDBb

[CH83b] G. M. Constantine and A. S. Hedayat. Complete designs with blocks of maximal multiplicity: Journal of statistical planning and inference 7 (1983) 289–294. Journal of Statistical Planning and Inference, 7(4):417, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900113.

Chang:1987:NSE

[CH87a] Cheng-Tao Chang and Klaus Hinkelmann. A new series of EGD-PBIB designs. Journal of Statistical Planning and Inference, 16 (??):1–13, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887900516.

Conforti:1987:SRP

[CH87b] Michele Conforti and Yosef Hochberg. Sequentially rejective pairwise testing procedures. Journal of Statistical Planning and Inference, 17(??):193–208, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901121.

Csorgo:1987:NTC

[CH87c] Miklós Csörgő and Lajos Horváth. Nonparametric tests for the changepoint problem. Journal of Statistical Planning and Inference, 17(??):1–9, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900978.

Charalambides:1986:GSD

[Cha86] Ch. A. Charalambides. Gould series distributions with applications to fluctuations of sums of random variables. Journal of Statistical Planning and Inference, 14(1):15–28, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886900054.

Chan:1988:ODL

[Cha88a] Ling-Yau Chan. Optimal design for a linear log contrast model for experiments with mixtures. Journal of Statistical Planning and Inference, 20(1):105–113, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900869.

Chandrasekar:1988:OCV

[Cha88b] B. Chandrasekar. An optimality criterion for vector unbiased statistical estimation functions. Journal of Statistical Planning and Inference, 18(1):115–117, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900213.

Chan:1989:ODL

[Cha89a] Ling-Yau Chan. Optimal design for a linear log contrast model for experiments with mixtures. Journal of Statistical Planning and Inference, 21(1):129, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900268.

Chan:1989:AIU

[Cha89b] Ngai Hang Chan. Asymptotic inference for unstable auto- regressive time series with drifts. Journal of Statistical Planning and Inference, 23(3):301–312, November 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900748.

Cheng:1980:ANM

[Che80] Smiley W. Cheng. On the ABLUE of the normal mean from a censored sample. Journal of Statistical Planning and Inference, 4(3):259–265, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375880900117.

Chen:1983:RSS

[Che83] Shan-Huo Chen. Ranked set sampling theory with selective probability vector. Journal of Statistical Planning and Inference, 8 (2):161–174, October 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900356.

Cheng:1989:NES

[Che89] Philip E. Cheng. Nonparametric estimation of survival curve under dependent censorship. Journal of Statistical Planning and Inference, 23(2):181–191, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900888.

Chikuse:1980:PCR

[Chi80] Yasuko Chikuse. Pseudo confidence regions for the solution of a multivariate linear functional relationship. Journal of Statistical Planning and Inference, 4(4):329–337, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375880900191.

Chiang:1986:TPS

[Chi86] Ching-Yuan Chiang. Tests of the parallelism of several regression surfaces based on rank-order estimates. Journal of Statistical Planning and Inference, 14(2–3):373–387, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901746.

Chmielewski:1981:ITE

[Chm81] M. A. Chmielewski. Invariant tests for the equality of k scale parameters under spherical symmetry. Journal of Statistical Planning and Inference, 5(4):341–346, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375881900343.

Chen:1984:AMC

[CHT84] T. Timothy Chen, Yosef Hochberg, and Aaron Tenenbein. Analysis of multivariate categorical data with misclassification errors by triple sampling schemes. Journal of Statistical Planning and Inference, 9(2):177–184, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900181.

Ceranka:1984:SMC

[CK84a] B. Ceranka and M. Kozlowska. Some methods for constructing Cdesigns. Journal of Statistical Planning and Inference, 9(2):253– 258, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900247.

Chandrasekar:1984:USE

[CK84b] B. Chandrasekar and B. K. Kale. Unbiased statistical estimation functions for parameters in presence of nuisance parameters. Journal of Statistical Planning and Inference, 9(1):45– 54, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900430.

Cohen:1986:GCB

[CK86] Daniel I. A. Cohen and Talbot M. Katz. A geometric connection between the cubic lattice and the honeycomb lattice. Journal of Statistical Planning and Inference, 14(1):29–34, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900066.

Chadjipantelis:1987:MDM

[CKM87] Theo Chadjipantelis, Stratis Kounias, and Chronis Moyssiadis. The maximum determinant of 21 × 21 (+1, −1)-matrices and d-optimal designs. Journal of Statistical Planning and Inference, 16(??): 167–178, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887900668.

Chaloner:1989:OBD

[CL89] Kathryn Chaloner and Kinley Larntz. Optimal Bayesian design applied to logistic regression experiments. Journal of Statistical Planning and Inference, 21(2):191–208, February 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375889900049.

Cline:1989:CLS

[Cli89] Daren B. H. Cline. Consistency for least squares regression estimators with infinite variance data. Journal of Statistical Planning and Inference, 23(2):163–179, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900876.

Constantine:1987:AOE

[CLS87] Kenneth B. Constantine, Yong B. Lim, and W. J. Studden. Admissible and optimal exact designs for polynomial regression. Journal of Statistical Planning and Inference, 16(??):15– 32, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887900528.

Ceranka:1986:NIB

[CM86a] Bronislaw Ceranka and Stanislaw Mejza. A note on the inequality b ≥ t for block designs. Journal of Statistical Planning and Inference, 13(??):97–101, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901230.

Chatterjee:1986:SSD

[CM86b] Kashinath Chatterjee and Rahul Mukerjee. Some search designs for symmetric and asymmetric factorials. Journal of Statistical Planning and Inference, 13(??):357–363, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588690145X.

Cuesta:1986:LRC

[CM86c] J. A. Cuesta and C. Matrán. L_{p}-linear regression, consistency and significative regression in median. Journal of Statistical Planning and Inference, 13(??):15–30, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901151.

Cohen:1988:OSR

[CM88] Jonathan P. Cohen and Shailendra S. Menjoge. One-sample run tests of symmetry. Journal of Statistical Planning and Inference, 18(1):93–100, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900195.

Cohen:1982:SOA

[CMS82] Arthur Cohen, John I. Marden, and Kesar Singh. Second order asymptotic and non-asymptotic optimality properties of combined tests. Journal of Statistical Planning and Inference, 6(3): 253–276, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900301.

Cox:1981:LSM

[Cox81] Lawrence H. Cox. Linear sensitivity measures in statistical disclosure control. Journal of Statistical Planning and Inference, 5(2): 153–164, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900252.

Chu:1980:SCV

[CP80] S. Sylvia Chu and K. C. S. Pillai. Some complex variable transformations and exact power comparisons of two-sided tests of equality of two Hermitian covariance matrices. Journal of Statistical Planning and Inference, 4(3):267–290, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375880900129.

Chiang:1984:ROT

[CP84] Ching-Yuan Chiang and Madan L. Puri. Rank-order tests for the parallelism of several regression surfaces. Journal of Statistical Planning and Inference, 10(1):43–57, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900314.

Carroll:1982:WCB

[CR82] Raymond J. Carroll and David Ruppert. Weak convergence of bounded influence regression estimates with applications to repeated significance testing. Journal of Statistical Planning and Inference, 7(2):117–129, December 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375882900702.

Cohen:1987:CMN

[CRS87] Arthur Cohen, Andrew Rukhin, and William E. Strawderman. A characterization of the multivariate normal distribution and some remarks on linear estimators. Journal of Statistical Planning and Inference, 17(??):361–365, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887901261.

Crow:1980:EUT

[CS80] L. H. Crow and I. N. Shimi. Estimation using type I censoring for a class of life-time distributions. Journal of Statistical Planning and Inference, 4(1):91–97, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375880900373.

Chen:1984:BTE

[CS84a] C. Chen and Theodore M. Smith. A Bayes-type estimator for the Bradley-Terry model for paired comparison. Journal of Statistical Planning and Inference, 10(1):9–14, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900284.

Cohen:1984:THA

[CS84b] Arthur Cohen and H. B. Sackrowitz. Testing hypotheses about the common mean of normal distributions. Journal of Statistical Planning and Inference, 9(2):207–227, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900223.

Choudhury:1988:GOS

[CS88] J. Choudhury and R. J. Serfling. Generalized order statistics, Bahadur representations, and sequential nonparametric fixed-width confidence intervals. Journal of Statistical Planning and Inference, 19(3):269–282, July 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900365.

Calvin:1989:MCV

[CS89] James A. Calvin and Kishore Sinha. A method for constructing variance balanced designs. Journal of Statistical Planning and Inference, 23(1):127–131, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900451.

Csaki:1986:SAC

[Csá86] E. Csáki. Some applications of the classical formula on ruin probabilities. Journal of Statistical Planning and Inference, 14(1): 35–42, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900078.

Cuevas:1988:QRA

[Cue88] Antonio Cuevas. Qualitative robustness in abstract inference. Journal of Statistical Planning and Inference, 18(3):277–289, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837588890105X.

Chernick:1983:EPM

[CW83] Michael R. Chernick and Tommy Wright. Estimation of a population mean with two-way stratification using a systematic allocation scheme. Journal of Statistical Planning and Inference, 7(3):219–231, March 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900162.

Csorgo:1989:TEM

[CW89] Sándor Csörgő and A. H. Welsh. Testing for exponential and Marshall–Olkin distributions. Journal of Statistical Planning and Inference, 23(3):287–300, November 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900736.

Das:1989:EVC

[Das89] Kalyan Das. Estimation of variance components in an unbalanced two-way mixed model under heteroskedasticity. Journal of Statistical Planning and Inference, 21(3):285–291, March 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900475.

deAlba:1980:EBA

[dAV80] Enrique de Alba and John Van Ryzin. An empirical Bayes approach to outliers. Journal of Statistical Planning and Inference, 4(3):217–236, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375880900087.

Dawson:1985:CGH

[Daw85] Jeremy E. Dawson. A construction for generalized Hadamard matrices GH(4q, EA(q)). Journal of Statistical Planning and Inference, 11(1):103–110, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900308.

Dayal:1985:ASU

[Day85] Shambhu Dayal. Allocation of sample using values of auxiliary characteristic. Journal of Statistical Planning and Inference, 11(3): 321–328, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900370.

Das:1989:NBB

[DD89] Ashish Das and A. Dey. A note on balanced block designs. Journal of Statistical Planning and Inference, 22(2):265–268, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889901171.

DeDominicis:1987:RAT

[De 87] R. De Dominicis. Reliability analysis of two-unit stand-by redundant systems allowing for aging. Journal of Statistical Planning and Inference, 16(??):317–328, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588790084X.

DeGroot:1988:BVA

[DeG88] Morris H. DeGroot. A Bayesian view of assessing uncertainty and comparing expert opinion. Journal of Statistical Planning and Inference, 20(3):295–306, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900948.

Does:1987:EES

[DHK87] R. J. M. M. Does, R. Helmers, and C. A. J. Klaassen. On the Edgeworth expansion for the sum of a function of uniform spacings. Journal of Statistical Planning and Inference, 17(??):149– 157, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588790108X.

Dempster:1988:UEA

[DK88] A. P. Dempster and Augustine Kong. Uncertain evidence and artificial analysis. Journal of Statistical Planning and Inference, 20(3):355–368, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900973.

Devroye:1989:ETC

[DK89] Luc Devroye and Adam Krzy[Zdot]ak. An equivalence theorem for L _{1} convergence of the kernel regression estimate. Journal of Statistical Planning and Inference, 23(1):71–82, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900402.

Deheuvels:1988:PAS

[DKPS88] Paul Deheuvels, Alan Karr, Dietmar Pfeifer, and Robert Serfling. Poisson approximations in selected metrics by coupling and semigroup methods with applications. Journal of Statistical Planning and Inference, 20(1):1–22, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900808.

Dean:1980:UTG

[DL80] A. M. Dean and S. M. Lewis. A unified theory for generalized cyclic designs. Journal of Statistical Planning and Inference, 4(1): 13–23, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375880900294.

deLauney:1984:NEG

[dL84a] Warwick de Launey. On the non-existence of generalised Hadamard matrices. Journal of Statistical Planning and Inference, 10(3):385– 396, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900624.

Drake:1984:QCL

[DL84b] David A. Drake and Jean A. Larson. A quest for certain linear spaces on thirty points. Journal of Statistical Planning and Inference, 10(2):241–255, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900764.

deLauney:1987:DMT

[dL87] Warwick de Launey. On difference matrices, transversal designs, resolvable transversal designs and large sets of mutually orthogonal F -squares. Journal of Statistical Planning and Inference, 16 (??):107–125, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900607.

Doksum:1984:EPH

[DN84] Kjell A. Doksum and Seiji Nabeya. Estimation in proportional hazard and log-linear models. Journal of Statistical Planning and Inference, 9(3):297–303, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900053.

Dale:1986:NTC

[DN86a] M. R. T. Dale and T. V. Narayana. Nonparametric tests for Catalan sequences. Journal of Statistical Planning and Inference, 14 (1):43–47, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588690008X.

Dale:1986:PCP

[DN86b] M. R. T. Dale and T. V. Narayana. A partition of Catalan permuted sequences with applications. Journal of Statistical Planning and Inference, 14(2–3):245–249, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901618.

Dale:1988:NDS

[DN88] M. R. T. Dale and T. V. Narayana. The number of decision sets involving n means. Journal of Statistical Planning and Inference, 18(3):371–376, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888901139.

Dockery:1988:CMD

[Doc88] John T. Dockery. C3 managerial decision aids for dealing with uncertainty — various approaches. Journal of Statistical Planning and Inference, 20(3):429–452, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888901012.

Donovan:1987:MBT

[Don87] Diane Donovan. More balanced ternary designs with block size four. Journal of Statistical Planning and Inference, 17(??):109– 133, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887901054.

Diaz:1983:SIS

[DORG83] Joaquín Diaz, Federico J. O’Reilly, and Santiago Rincon-Gallardo. A set of independent sequential residuals for the multivariate regression model. Journal of Statistical Planning and Inference, 8 (1):21–25, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900587.

Dalenius:1982:DST

[DR82] Tore Dalenius and Steven P. Reiss. Data-swapping: a technique for disclosure control. Journal of Statistical Planning and Inference, 6(1):73–85, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900581.

Denker:1985:MER

[DR85] Manfred Denker and Uwe Rösler. A moment estimate for rank statistics. Journal of Statistical Planning and Inference, 12(??): 269–284, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588590076X.

Durairajan:1989:CUO

[Dur89] T. M. Durairajan. Characterisation and uniqueness of optimal estimating function. Journal of Statistical Planning and Inference, 22(3):391–395, July 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588990102X.

Dupreez:1989:GAE

[DV89] J. P. Dupreez and J. H. Venter. A group adaptive empirical Bayes estimator of normal means. Journal of Statistical Planning and Inference, 22(1):29–41, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900621.

Dykstra:1981:IRA

[Dyk81] Richard L. Dykstra. An isotonic regression algorithm. Journal of Statistical Planning and Inference, 5(4):355–363, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375881900367.

Ebrahimi:1986:EPE

[Ebr86] Nader Ebrahimi. Estimating the parameters of an exponential distribution from a hybrid life test. Journal of Statistical Planning and Inference, 14(2–3):255–261, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901631.

Eccleston:1980:RTC

[Ecc80] J. A. Eccleston. Recursive techniques for the construction of experimental designs. Journal of Statistical Planning and Inference, 4(3):291–297, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375880900130.

Erdos:1982:PD

[EFJ82] P. Erdős, V. Faber, and F. Jones. Projective (2n, n, λ, 1)- designs. Journal of Statistical Planning and Inference, 7(2):181– 191, December 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900763.

Eichhorn:1983:SRR

[EH83] Benjamin H. Eichhorn and Lakhbir S. Hayre. Scrambled randomized response methods for obtaining sensitive quantitative data. Journal of Statistical Planning and Inference, 7(4):307– 316, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375883900022.

El-Helbawy:1984:ARE

[EH84] Abdalla T. El-Helbawy. Asymptotic relative efficiency of designs for factorial paired comparison experiments. Journal of Statistical Planning and Inference, 10(1):105–113, July 1984. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375884900363.

Eccleston:1981:ROI

[EK81] J. Eccleston and J. Kiefer. Relationships of optimality for individual factors of a design. Journal of Statistical Planning and Inference, 5(3):213–219, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588190001X.

Eubank:1986:RTT

[EL86] R. L. Eubank and V. N. LaRiccia. Regression type tests for parametric hypotheses based on sums of squared L-statistics. Journal of Statistical Planning and Inference, 14(2–3):401–407, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837588690176X.

Eccleston:1985:ARC

[EM85] J. A. Eccleston and C. A. McGilchrist. Algebra of a row-column design. Journal of Statistical Planning and Inference, 12(??):305– 310, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900795.

Eubank:1989:CRE

[EW89] R. L. Eubank and P. Whitney. Convergence rates for estimation in certain partially linear models. Journal of Statistical Planning and Inference, 23(1):33–43, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900372.

Fedorov:1989:ODB

[Fed89] V. V. Fedorov. Optimal design with bounded density: optimization algorithms of the exchange type. Journal of Statistical Planning and Inference, 22(1):1–13, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900608.

Feingold:1985:TSC

[Fei85] Marcia Feingold. A test statistic for combined intra- and interblock estimates. Journal of Statistical Planning and Inference, 12(??):103–114, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900588.

Fellman:1989:ESC

[Fel89] Johan Fellman. An empirical study of a class of iterative searches for optimal designs. Journal of Statistical Planning and Inference, 21(1):85–92, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900219.

Friedman:1985:ESM

[FG85] Lea Friedman and Ilya Gertsbakh. Estimation of a standard measuring error by repeated measurements and sortings. Journal of Statistical Planning and Inference, 11(1):1–14, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900205.

Foody:1988:GPN

[FH88] W. M. Foody and A. S. Hedayat. A graphical proof of the nonexistence of BIB(7, b, r, 3, λ|16) designs. Journal of Statistical Planning and Inference, 20(1):77–90, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900845.

Federer:1984:POR

[FHM84] W. T. Federer, A. S. Hedayat, and J. P. Mandeli. Pairwise orthogonal F -rectangle designs. Journal of Statistical Planning and Inference, 10(3):365–374, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900600.

Fine:1988:LPM

[Fin88] Terrence L. Fine. Lower probability models for uncertainty and nondeterministic processes. Journal of Statistical Planning and Inference, 20(3):389–411, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900997.

Frank:1980:ICI

[FK80] Ove Frank and Gösta Karlsson. Inferences concerning interpersonal choices given anonymously. Journal of Statistical Planning and Inference, 4(1):51–65, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375880900336.

Flury:1985:NMP

[Flu85] Bernhard N. Flury. A note on multivariate parallel regression. Journal of Statistical Planning and Inference, 12(??):369– 377, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900850.

Flury:1989:BRA

[Flu89a] Bernhard Flury. Book review: Algorithms for clustering data: Anil K. Jain and Richard C. Dubes Prentice Hall Advanced Reference Series in Computer Science Prentice Hall, Englewood Cliffs, NJ, 1988. Journal of Statistical Planning and Inference, 21(1):137– 138, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900335.

Flury:1989:BRE

[Flu89b] Bernhard Flury. Book review: Estimation of variance components and applications: C. R. Rao and J. Kleffe, North-Holland Series in Statistics and Probability, Vol. 3 North-Holland, New York, 1988. Journal of Statistical Planning and Inference, 21(1):136– 137, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900323.

Fraser:1988:LIS

[FM88] D. A. S. Fraser and H. Massam. Location inference on spheres and cylinders. Journal of Statistical Planning and Inference, 18 (2):195–201, February 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900055.

Ferentinos:1982:IES

[FP82] K. Ferentinos and T. Papaioannou. Information in experiments and sufficiency. Journal of Statistical Planning and Inference, 6(4): 309–317, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900015.

Frank:1980:ENV

[Fra80] Ove Frank. Estimation of the number of vertices of different degrees in a graph. Journal of Statistical Planning and Inference, 4 (1):45–50, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375880900324.

From:1989:EPM

[FS89] Steven G. From and K. M. Lal Saxena. Estimating parameters from mixed samples using sample fractional moments. Journal of Statistical Planning and Inference, 21(2):231–244, February 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900074.

Funatsu:1982:MDV

[Fun82] Y. Funatsu. A method of deriving valid approximate expressions for bias in ratio estimation. Journal of Statistical Planning and Inference, 6(3):215–225, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588290026X.

Frank:1987:CII

[FW87] Ove Frank and Per Weidenman. Controlling individual information in statistics by coding. Journal of Statistical Planning and Inference, 17(??):321–336, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901236.

Ghosh:1985:SNF

[GA85] Subir Ghosh and Diana Avila. Some new factor screening designs using the search linear model. Journal of Statistical Planning and Inference, 11(2):259–266, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900138.

Gaffke:1982:OBD

[Gaf82] N. Gaffke. D-optimal block designs with at most six varieties. Journal of Statistical Planning and Inference, 6(2):183– 200, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588290088X.

Gardiner:1983:POE

[Gar83] Joseph C. Gardiner. Properties of occurrence-exposure rate processes in multiple decrement theory. Journal of Statistical Planning and Inference, 8(3):301–314, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900472.

Gastwirth:1982:SPM

[Gas82] Joseph L. Gastwirth. Statistical properties of a measure of tax assessment uniformity. Journal of Statistical Planning and Inference, 6(1):1–12, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900507.

Gasim:1988:FOA

[Gas88] Ali Abul Gasim. First-order autoregressive models: a method for obtaining eigenvalues for weighting matrices. Journal of Statistical Planning and Inference, 18(3):391–398, March 1988. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375888901152.

Gil:1988:MIE

[GCG88] M. A. Gil, N. Corral, and P. Gil. The minimum inaccuracy estimates in χ^{2} tests for goodness of fit with fuzzy observations. Journal of Statistical Planning and Inference, 19(1):95– 115, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900559.

Gelfand:1988:EVR

[GD88] Alan E. Gelfand and Dipak K. Dey. On the estimation of a variance ratio. Journal of Statistical Planning and Inference, 19(1): 121–131, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900572.

Georgiev:1987:FNW

[Geo87] Alexander A. Georgiev. F3. A note on the weak pointwise convergence of the Priestley–Chao multiple regression estimate. Journal of Statistical Planning and Inference, 16(??):261–262, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375887900760.

Gessel:1986:PML

[Ges86] Ira M. Gessel. A probabilistic method for lattice path enumeration. Journal of Statistical Planning and Inference, 14(1): 49–58, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900091.

Gupta:1981:PTM

[GG81] Pushpa Lata Gupta and Ramesh C. Gupta. Probability of ties and Markov property in discrete order statistics. Journal of Statistical Planning and Inference, 5(3):273–279, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375881900070.

Georgiev:1986:NFR

[GG86a] Alexander A. Georgiev and Wlodzimierz Greblicki. Nonparametric function recovering from noisy observations. Journal of Statistical Planning and Inference, 13(??):1–14, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588690114X.

Ghiassi:1986:ADF

[GG86b] S. H. Mansouri Ghiassi and Z. Govindarajulu. An asymptotically distribution-free test for ordered alternatives in two-way layouts. Journal of Statistical Planning and Inference, 13(??):239– 249, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901369.

Gokhale:1986:OTC

[GG86c] D. V. Gokhale and Ashis Sen Gupta. Optimal tests for the correlation coefficient in a symmetric multivariate normal population. Journal of Statistical Planning and Inference, 14(2–3):263– 268, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901643.

Gupta:1980:ECC

[GH80] Shanti S. Gupta and Deng-Yuang Huang. An essentially complete class of multiple decision procedures. Journal of Statistical Planning and Inference, 4(2):115–121, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375880900014.

Gupta:1983:EBR

[GH83] Shanti S. Gupta and Ping Hsiao. Empirical Bayes rules for selecting good populations. Journal of Statistical Planning and Inference, 8 (1):87–101, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900642.

Gupta:1988:SII

[GH88] Shanti S. Gupta and Deng-Yuan Huang. Selecting important independent variables in linear regression models. Journal of Statistical Planning and Inference, 20(2):155–167, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837588890119X.

Gastwirth:1989:EPR

[GH89] Joseph L. Gastwirth and Patricia A. Hammick. Estimation of the prevalence of a rare disease, preserving the anonymity of the subjects by group testing: application to estimating the prevalence of aids antibodies in blood donors. Journal of Statistical Planning and Inference, 22(1):15–27, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588990061X.

Ghosh:1980:ROB

[Gho80] Subir Ghosh. On robustness of optimal balanced resolution v plans. Journal of Statistical Planning and Inference, 4(3):313– 319, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375880900154.

Ghosh:1981:SNS

[Gho81] Subir Ghosh. On some new search designs for 2^{m} factorial experiments. Journal of Statistical Planning and Inference, 5(4): 381–389, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900392.

Ghosh:1982:RBA

[Gho82] Subir Ghosh. Robustness of BIBD against the unavailability of data. Journal of Statistical Planning and Inference, 6(1):29– 32, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900532.

Ghosh:1988:SDF

[Gho88] Subir Ghosh. Statistical discussion forum. Journal of Statistical Planning and Inference, 18(1):133–134, January 1988. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375888900249.

Ghosh:1989:CCD

[Gho89a] D. K. Ghosh. Construction of confounded designs for mixed factorial experiments. Journal of Statistical Planning and Inference, 23(2):253–261, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900931.

Ghosh:1989:FCS

[Gho89b] Subir Ghosh. F19. A comment on “Singular factorial designs resulting from missing pairs of design points”. Journal of Statistical Planning and Inference, 21(1):133–134, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900293. See [Moo88].

Gilstein:1985:OPF

[Gil85] C. Zachary Gilstein. Optimal partitions of finite populations for Dorfman-type group testing. Journal of Statistical Planning and Inference, 12(??):385–394, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900874.

Gill:1986:BIA

[Gil86a] P. S. Gill. Balanced incomplete arrays. Journal of Statistical Planning and Inference, 14(2–3):179–185, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901552.

Gilula:1986:SSB

[Gil86b] Zvi Gilula. On some similarities between contingency table methods and grouped data survival analytic techniques. Journal of Statistical Planning and Inference, 14(2–3):203–211, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886901576.

Greblicki:1980:APK

[GK80] Wlodzimierz Greblicki and Adam Krzyzak. Asymptotic properties of kernel estimates of a regression function. Journal of Statistical Planning and Inference, 4(1):81–90, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375880900361.

Galil:1983:CDE

[GK83a] Z. Galil and J. Kiefer. Comparison of designs equivalent under one or two criteria. Journal of Statistical Planning and Inference, 8 (1):103–116, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900654.

Gnot:1983:QEM

[GK83b] Stanislaw Gnot and Jürgen Kleffe. Quadratic estimation in mixed linear models with two variance components. Journal of Statistical Planning and Inference, 8(3):267–279, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375883900459.

Ghosh:1985:ROB

[GK85] Subir Ghosh and Wilson A. K. Kipngeno. On the robustness of the optimum balanced 2^{m} factorial designs of resolution v (given by Srivastava and Chopra) in the presence of outliers. Journal of Statistical Planning and Inference, 11(1):119– 129, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900321.

Gnot:1985:NAI

[GKZ85] S. Gnot, J. Kleffe, and R. Zmyślony. Nonnegativity of admissible invariant quadratic estimates in mixed linear models with two variance components. Journal of Statistical Planning and Inference, 12(??):249–258, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900734.

Gupta:1989:SBB

[GL89] Shanti S. Gupta and TaChen Liang. Selecting the best binomial population: parametric empirical Bayes approach. Journal of Statistical Planning and Inference, 23(1):21–31, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900360.

Genizi:1984:EVB

[GLP84] A. Genizi, E. Lahav, and J. Putter. Estimation of the variance between units of given size from experimental data. Journal of Statistical Planning and Inference, 9(1):83–101, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375884900478.

Gupta:1985:MMT

[GM85] Shanti S. Gupta and Klaus J. Miescke. Minimax multiple ttests for comparing k normal populations with a control. Journal of Statistical Planning and Inference, 12(??):161–169, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900655.

Gupta:1985:CAO

[GN85] V. K. Gupta and A. K. Nigam. A class of asymmetrical orthogonal resolution-IV designs. Journal of Statistical Planning and Inference, 11(3):381–383, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900436.

Gnot:1987:QEV

[Gno87] Stanislw Gnot. Quadratic estimation of variance components in mixed block designs. Journal of Statistical Planning and Inference, 16(??):231–236, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900723.

Goicoechea:1988:ESM

[Goi88] Ambrose Goicoechea. Expert system models for inference with imperfect knowledge: a comparative study. Journal of Statistical Planning and Inference, 20(3):245–277, ???? 1988. CODEN JSPIDN.

ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900924.

Good:1987:FGV

[Goo87a] I. J. Good. F6. Graig the vague. Journal of Statistical Planning and Inference, 16(??):262–263, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887900772.

Good:1987:FSI

[Goo87b] I. J. Good. F7. Some improvements in statistical and other ‘literature’. Journal of Statistical Planning and Inference, 16(??): 263–265, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887900784.

Good:1987:FAS

[Goo87c] I. J. Good. F8. Is the arc-sine law that counter-intuitive? Journal of Statistical Planning and Inference, 16(??):429–430, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375887900930.

Good:1989:FER

[Goo89a] I. J. Good. F20. An enveloping regression procedure. Journal of Statistical Planning and Inference, 21(1):134, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837588990030X.

Good:1989:FCV

[Goo89b] I. J. Good. F21. Chaotics vs. chaology. Journal of Statistical Planning and Inference, 21(1):134–135, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900311.

Gould:1983:SSR

[Gou83] A. Lawrence Gould. Sample sizes required for binomial trials when the true response rates are estimated. Journal of Statistical Planning and Inference, 8(1):51–58, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900617.

Govindarajulu:1980:CGD

[Gov80] Z. Govindarajulu. Characterization of the geometric distribution using properties of order statistics. Journal of Statistical Planning and Inference, 4(3):237–247, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375880900099.

Govindarajulu:1984:STO

[Gov84] Z. Govindarajulu. Stopping times of one-sample rank order sequential probability ratio tests. Journal of Statistical Planning and Inference, 9(3):305–320, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900065.

Giovagnoli:1987:GIO

[GPW87] Alessandra Giovagnoli, Friedrich Pukelsheim, and Henry P. Wynn. Group invariant orderings and experimental designs. Journal of Statistical Planning and Inference, 17(??):159–171, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375887901091.

Gupta:1980:ERM

[GR80] A. K. Gupta and V. K. Rohatgi. On the estimation of restricted mean. Journal of Statistical Planning and Inference, 4(4):369– 379, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375880900221.

Ghosh:1982:OPD

[GR82] S. Ghosh and J. Roy. Optimality in presence of damaged observations. Journal of Statistical Planning and Inference, 6(2): 123–126, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900805.

Gupta:1985:TOE

[GR85] Rameshwar Gupta and Donald Richards. Testing optimality of experimental designs for a regression model with random variables. Journal of Statistical Planning and Inference, 11(1):75– 80, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900266.

Goel:1987:SPM

[GR87] Prem K. Goel and T. Ramalingam. Some properties of the maximum likelihood strategy for re-pairing a broken random sample. Journal of Statistical Planning and Inference, 16(??):237– 248, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887900735.

Grambsch:1989:SML

[Gra89] Patricia Grambsch. Sequential maximum likelihood estimation with applications to logistic regression in case-control studies. Journal of Statistical Planning and Inference, 22(3):355–369, July 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889901006.

Ghosh:1983:RPP

[GRS83] S. Ghosh, S. B. Rao, and N. M. Singhi. On a robustness property of PBIBD. Journal of Statistical Planning and Inference, 8 (3):355–363, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900514.

Gill:1987:RMA

[GS87] D. S. Gill and Minoru Siotani. On randomization in multivariate analysis of variance. Journal of Statistical Planning and Inference, 17(??):217–226, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901145.

Gardiner:1985:FWC

[GSV85] Joseph C. Gardiner, V. Susarla, and John Van Ryzin. Fixed width confidence intervals for the exponential mean survival time under random censorship. Journal of Statistical Planning and Inference, 12(??):153–159, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900643.

Gardiner:1988:BLC

[GSvR88] Joseph C. Gardiner, V. Susarla, and John van Ryzin. On bounded length confidence intervals for quantiles of the survival distribution. Journal of Statistical Planning and Inference, 18(2):245– 253, February 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900092.

Godambe:1989:EQL

[GT89] V. P. Godambe and M. E. Thompson. An extension of quasilikelihood estimation. Journal of Statistical Planning and Inference, 22(2):137–152, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889901067. See discussion

[Ano89d].

Gupta:1982:DMD

[Gup82] Ramesh C. Gupta. On the determination of mixing density in reliability studies. Journal of Statistical Planning and Inference, 7 (1):101–105, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900234.

Gupta:1983:BLA

[Gup83] S. C. Gupta. A basic lemma and the analysis of block and Kronecker product designs. Journal of Statistical Planning and Inference, 7(4):407–416, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900101.

Gupta:1984:EPW

[Gup84] Ramesh C. Gupta. Estimating the probability of winning (losing) in a gambler’s ruin problem with applications. Journal of Statistical Planning and Inference, 9(1):55–62, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375884900442.

Gupta:1985:IAT

[Gup85a] S. C. Gupta. Iterative analysis of two- and three-way designs. Journal of Statistical Planning and Inference, 11(1):95– 102, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900291.

Gupta:1985:KBD

[Gup85b] S. C. Gupta. On Kronecker block designs for factorial experiments. Journal of Statistical Planning and Inference, 11(2):227– 236, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900096.

Gupta:1986:IEK

[Gup86] S. C. Gupta. Interaction efficiencies in Kronecker block designs. Journal of Statistical Planning and Inference, 14(2–3):275– 279, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901667.

Gupta:1987:MPR

[Gup87] Ramesh C. Gupta. On the monotonic properties of the residual variance and their applications in reliability. Journal of Statistical Planning and Inference, 16(??):329–335, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887900851.

Gupta:1988:AME

[Gup88] Sudhir Gupta. The association matrices of extended group divisible scheme. Journal of Statistical Planning and Inference, 20(1): 115–119, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900870.

Gupta:1989:BDS

[Gup89] Sudhir Gupta. On block designs for symmetrical parallel line assays with even number of doses. Journal of Statistical Planning and Inference, 21(3):383–389, March 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900542.

Gutmann:1987:TUM

[Gut87] Sam Gutmann. Tests uniformly more powerful than uniformly most powerful monotone tests. Journal of Statistical Planning and Inference, 17(??):279–292, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901200.

Gijbels:1988:WAR

[GV88] I. Gijbels and N. Veraverbeke. Weak asymptotic representations for quantiles of the product-limit estimator. Journal of Statistical Planning and Inference, 18(2):151–160, February 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837588890002X.

Gijbels:1989:SFW

[GV89] I. Gijbels and N. Veraverbeke. Sequential fixed-width confidence intervals for quantiles in the presence of censoring. Journal of Statistical Planning and Inference, 22(2):213–222, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889901122.

Giovagnoli:1980:MTM

[GW80] Alessandra Giovagnoli and H. P. Wynn. A majorization theorem for the C-matrices of binary designs. Journal of Statistical Planning and Inference, 4(2):145–154, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375880900038.

Gastwirth:1988:CPT

[GW88] Joseph L. Gastwirth and Jane-Ling Wang. Control percentile test procedures for censored data. Journal of Statistical Planning and Inference, 18(3):267–276, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888901048.

Gupta:1985:BPS

[GY85] Shanti S. Gupta and Hwa-Ming Yang. Bayes-p ^{∗} subset selection procedures for the best population. Journal of Statistical Planning and Inference, 12(??):213–233, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900710.

Haber:1981:APR

[Hab81] Michael Haber. On the asymptotic power and relative efficiency of the frequency X ^{2} test. Journal of Statistical Planning and Inference, 5(3):299–308, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375881900100.

Halteman:1986:CCM

[Hal86] Edward J. Halteman. The Chebyshev center: a multidimensional estimate of location. Journal of Statistical Planning and Inference, 13(??):389–394, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901485.

Harel:1988:WCM

[Har88] Michel Harel. Weak convergence of multidimensional rank statistics under ø-mixing conditions. Journal of Statistical Planning and Inference, 20(1):41–63, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900821.

Hayakawa:1982:SFW

[Hay82a] Takesi Hayakawa. On some formulas for weighted sum of invariant polynomials of two matrix arguments. Journal of Statistical Planning and Inference, 6(2):105–114, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375882900775.

Hayre:1982:NAO

[Hay82b] Lakhbir Singh Hayre. A note on the asymptotic optimality of a two population sequential probability ratio test. Journal of Statistical Planning and Inference, 6(2):127–130, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900817.

Hayakawa:1986:THC

[Hay86] T. Hayakawa. On testing hypothesis of covariance matrices under elliptical population. Journal of Statistical Planning and Inference, 13(??):193–202, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901321.

Hayakawa:1989:DFM

[Hay89] T. Hayakawa. On the distributions of the functions of the F - matrix under an elliptical population. Journal of Statistical Planning and Inference, 21(1):41–52, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900177.

Hamada:1989:CMH

[HD89] Noboru Hamada and Michel Deza. A characterization of ν_{μ} + 1 + ϵ, ν_{μ}; t, q-min.hypers and its applications to error-correcting codes and factorial designs. Journal of Statistical Planning and Inference, 22(3):323–336, July 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900980.

Heckman:1987:RDT

[Hec87] Nancy E. Heckman. Robust design in a two treatment comparison in the presence of a covariate. Journal of Statistical Planning and Inference, 16(??):75–81, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887900577.

Hedayat:1984:CUO

[Hed84] A. Hedayat. A characterization of a universally optimal design within a class of block designs. Journal of Statistical Planning and Inference, 9(1):143–145, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588490051X.

Hart:1985:AMA

[HG85] Jeffrey D. Hart and H. L. Gray. The ARMA method of approximating probability density functions. Journal of Statistical Planning and Inference, 12(??):137–152, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900631.

Holtzman:1986:PCS

[HG86] Golde I. Holtzman and I. J. Good. The Poisson and chisquared approximations as compared with the true upper-tail probability of Pearson’s X ^{2} for equiprobable multinomials. Journal of Statistical Planning and Inference, 13(??):283–295, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886901400.

Hogg:1984:AE

[HHL84] Robert V. Hogg, Paul S. Horn, and Russell V. Lenth. On adaptive estimation. Journal of Statistical Planning and Inference, 9(3): 333–343, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900089.

Huskova:1981:SOA

[HJ81] Marie Husková and Jana Jurecková. Second order asymptotic relations of M-estimators and R-estimators in two-sample location model. Journal of Statistical Planning and Inference, 5(4): 309–328, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900318.

Hjorth:1989:MSC

[Hjo89] Urban Hjorth. On model selection in the computer age. Journal of Statistical Planning and Inference, 23(1):101–115, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900438.

Hedayat:1980:FDP

[HK80] A. Hedayat and S. Kageyama. The family of t-designs-part I. Journal of Statistical Planning and Inference, 4(2):173–212, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375880900063.

Hedayat:1989:RSB

[HLT89] A. S. Hedayat, I. N. Landgev, and V. D. Tonchev. Results on the support of BIB designs. Journal of Statistical Planning and Inference, 22(3):295–306, July 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900967.

Handa:1982:FUG

[HM82] B. R. Handa and S. G. Mohanty. A formula for P (b_{i} ≤ R_{i} ≤ a_{i}, 1 ≤ i ≤ m) under a general class of alternatives. Journal of Statistical Planning and Inference, 6(4):335– 343, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900039.

Hochberg:1983:TPT

[HM83] Yosef Hochberg and Ruth Marcus. Two-phase tests on a normal mean when variance is unknown. Journal of Statistical Planning and Inference, 7(3):233–242, March 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900174.

Handa:1986:PLP

[HM86] B. R. Handa and S. G. Mohanty. On a property of lattice paths. Journal of Statistical Planning and Inference, 14(1):59– 62, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900108.

Hedayat:1988:MRO

[HM88] A. S. Hedayat and Dibyen Majumdar. Model robust optimal designs for comparing test treatments with a control. Journal of Statistical Planning and Inference, 18(1):25–33, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900134.

Hodoshima:1988:LMP

[Hod88] Jiro Hodoshima. Locally most powerful rank tests for regression under truncation. Journal of Statistical Planning and Inference, 18 (2):239–244, February 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900080.

Horn:1985:BBM

[Hor85] Paul S. Horn. Back-to-back mixtures of discrete distributions. Journal of Statistical Planning and Inference, 11(3):267–276, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900345.

Hornik:1988:AOT

[Hor88] K. Hornik. On asymptotically optimal tests. Journal of Statistical Planning and Inference, 20(2):131–154, ???? 1988. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375888901188.

Hornik:1989:AOT

[Hor89] K. Hornik. Asymptotically optimal tests of hypotheses on the generator of a finite Markov process. Journal of Statistical Planning and Inference, 23(3):345–352, November 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900773.

Hedayat:1985:SBD

[HP85] A. Hedayat and H. Pesotan. A study of BIB designs through support matrices. Journal of Statistical Planning and Inference, 11(3):363–372, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900412.

Huang:1984:SLO

[HPT84] Deng-Yuan Huang, S. Panchapakesan, and Sheng-Tsaing Tseng. Some locally optimal subset selection rules for comparison with a control. Journal of Statistical Planning and Inference, 9(1):63– 72, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900454.

Helmers:1988:ANG

[HR88] R. Helmers and F. H. Ruymgaart. Asymptotic normality of generalized L-statistics with unbounded scores. Journal of Statistical Planning and Inference, 19(1):43–53, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900511.

Hedayat:1988:SPE

[HRS88] A. S. Hedayat, C. R. Rao, and J. Stufken. Sampling plans excluding contiguous units. Journal of Statistical Planning and Inference, 19(2):159–170, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900705.

Huang:1988:MRE

[HS88] Mong-Na Lo Huang and William J. Studden. Model robust extrapolation designs. Journal of Statistical Planning and Inference, 18(1):1–24, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900122.

Hassanein:1986:ETQ

[HSB86] Khatab M. Hassanein, A. K. Md. Ehsanes Saleh, and Edward F. Brown. Estimation and testing of quantiles of the extreme-value distribution. Journal of Statistical Planning and Inference, 14(2– 3):389–400, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901758.

Huda:1983:TTO

[Hud83] S. Huda. Two third-order rotatable designs in four dimensions. Journal of Statistical Planning and Inference, 8(2):241– 243, October 1983. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375883900423.

Huda:1985:VDB

[Hud85] S. Huda. Variance of the difference between two estimated responses. Journal of Statistical Planning and Inference, 11(1):89– 93, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588590028X.

Hwang:1986:ST

[Hwa86] H. L. Hwang. On the structure of (v, k, t) trades. Journal of Statistical Planning and Inference, 13(??):179–191, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837588690131X.

Hwang:1987:GTI

[HX87] F. K. Hwang and Y. H. Xu. Group testing to identify one defective and one mediocre item. Journal of Statistical Planning and Inference, 17(??):367–373, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901273.

Iachan:1985:RDR

[Iac85] Ronaldo Iachan. Robust designs for ratio and regression estimation. Journal of Statistical Planning and Inference, 11(2):149– 161, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900023.

Ipinyomi:1988:EBN

[IF88] R. A. Ipinyomi and G. H. Freeman. Equineighboured balanced nested row-column designs. Journal of Statistical Planning and Inference, 20(2):191–199, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588890122X.

Yoshihara:1984:DES

[iY84] Ken ichi Yoshihara. Density estimation for samples satisfying a certain absolute regularity condition. Journal of Statistical Planning and Inference, 9(1):19–32, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900417.

Jacroux:1984:BSP

[Jac84a] Mike Jacroux. On the block structure of proper block designs. Journal of Statistical Planning and Inference, 9(2):259– 267, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900259.

Jacroux:1984:OGD

[Jac84b] Mike Jacroux. On the D-optimality of group divisible designs. Journal of Statistical Planning and Inference, 9(1):119– 129, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900491.

Jacroux:1985:SSC

[Jac85] Mike Jacroux. Some sufficient conditions for the type 1 optimality of block designs. Journal of Statistical Planning and Inference, 11(3):385–398, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900448.

Jacroux:1986:OSB

[Jac86] Mike Jacroux. On the E- and MV -optimality of spring balance designs with string property. Journal of Statistical Planning and Inference, 13(??):89–96, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901229.

Jacroux:1988:SFR

[Jac88] Mike Jacroux. Some further results on the MV-optimality of block designs for comparing test treatments to a standard treatment. Journal of Statistical Planning and Inference, 20(2):201– 214, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888901231.

Jacroux:1989:SSC

[Jac89] Mike Jacroux. Some sufficient conditions for type 1 optimality with applications to regular graph designs. Journal of Statistical Planning and Inference, 23(2):193–215, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588990089X.

Jain:1981:DRS

[Jai81] Aridaman K. Jain. Design of a rotation scheme for a stratified multi-stage sample. Journal of Statistical Planning and Inference, 5(1):57–69, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375881900513.

Johnson:1985:EIM

[JG85] Wesley Johnson and Seymour Geisser. Estimative influence measures for the multivariate general linear model. Journal of Statistical Planning and Inference, 11(1):33–56, January 1985. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375885900230.

James:1983:ICQ

[JJ83] Barry R. James and Kang Ling James. On the influence curve for quantal bioassay. Journal of Statistical Planning and Inference, 8 (3):331–345, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900496.

Jacroux:1989:OBD

[JM89] Mike Jacroux and Dibyen Majumdar. Optimal block designs for comparing test treatments with a control when k > υ. Journal of Statistical Planning and Inference, 23(3):381–396, November 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900803.

John:1981:FBA

[Joh81] J. A. John. Factorial-balance and the analysis of designs with factorial structure. Journal of Statistical Planning and Inference, 5(1): 99–105, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900550.

Joshi:1982:NMM

[Jos82] P. C. Joshi. A note on the mixed moments of order statistics from exponential and truncated exponential distributions. Journal of Statistical Planning and Inference, 6(1):13–16, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900519.

Joshi:1988:RVP

[Jos88] V. M. Joshi. Randomization versus the prediction approach in survey sampling. Journal of Statistical Planning and Inference, 20(2):215–219, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888901243.

Joshi:1989:FFA

[Jos89] V. M. Joshi. F17. Fallacies in an argument of the Finetti relating to the extension of probability. Journal of Statistical Planning and Inference, 21(1):131–132, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588990027X.

Jacroux:1980:SSC

[JS80] Michael Jacroux and Justus Seely. Some sufficient conditions for establishing ( M, S )-optimality. Journal of Statistical Planning and Inference, 4(1):3–11, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375880900282.

Joshi:1981:SST

[JS81a] Suneeta Joshi and Y. S. Sathe. A k-sample slippage test for location parameter. Journal of Statistical Planning and Inference, 5 (1):93–98, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900549.

Jureckova:1981:SPB

[JS81b] Jana Jurecková and Pranab Kumar Sen. Sequential procedures based on M-estimators with discontinuous score functions. Journal of Statistical Planning and Inference, 5(3):253– 266, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900057.

Janssen:1987:ANS

[JSV87] Paul Janssen, Robert Serfling, and Noël Veraverbeke. Asymptotic normality of U-statistics based on trimmed samples. Journal of Statistical Planning and Inference, 16(??):63–74, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375887900565.

Jammalamadaka:1985:ACT

[JT85] S. Rao Jammalamadaka and Ram C. Tiwari. Asymptotic comparison of three tests for goodness of fit. Journal of Statistical Planning and Inference, 12(??):295–304, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900783.

Jureckova:1980:RCO

[Jur80] Jana Jurecková. Rate of consistency of one sample tests of location. Journal of Statistical Planning and Inference, 4(3):249– 257, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375880900105.

Jacroux:1983:OCB

[JWM83] Mike Jacroux, C. S. Wong, and J. C. Masaro. On the optimality of chemical balance weighing designs. Journal of Statistical Planning and Inference, 8(2):231–240, October 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900411.

Kageyama:1982:CTA

[Kag82] Sanpei Kageyama. Connectedness of two-associate PBIB designs. Journal of Statistical Planning and Inference, 7(1):77– 82, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900209.

Kageyama:1985:CGD

[Kag85] Sanpei Kageyama. A construction of group divisible designs. Journal of Statistical Planning and Inference, 12(??):123–125, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900606.

Kageyama:1986:ACT

[Kag86] Sanpei Kageyama. Addendum to “connectedness of two-associate PBIB designs”. Journal of Statistical Planning and Inference, 13 (??):131–132, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901278.

Kageyama:1988:OAA

[Kag88a] S. Kageyama. On orthogonal arrays attaining Rao’s bounds. Journal of Statistical Planning and Inference, 19(3):395–400, July 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900468.

Kageyama:1988:NEC

[Kag88b] Sanpei Kageyama. A note on A-efficiency of connected designs. Journal of Statistical Planning and Inference, 20(2):221– 224, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888901255.

Kageyama:1989:FCS

[Kag89] Sanpei Kageyama. F18. Construction of SRGD designs. Journal of Statistical Planning and Inference, 21(1):132–133, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900281.

Kaiser:1989:NBU

[Kai89] Lee D. Kaiser. Nonexistence of blue’s under a randomization model for the randomized block design. Journal of Statistical Planning and Inference, 22(1):63–69, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900657.

Kale:1985:NSE

[Kal85] B. K. Kale. A note on the super efficient estimator. Journal of Statistical Planning and Inference, 12(??):259–263, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900746.

Kale:1987:EUO

[Kal87] B. K. Kale. Essential uniqueness of optimal estimating functions. Journal of Statistical Planning and Inference, 17(??):405– 407, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887901315.

Kappenman:1988:RSD

[Kap88] Russell F. Kappenman. Robust symmetric distribution location estimation and regression. Journal of Statistical Planning and Inference, 19(1):55–72, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900523.

Krewski:1981:SJV

[KC81] D. Krewski and R. P. Chakrabarty. On the stability of the jackknife variance estimator in ratio estimation. Journal of Statistical Planning and Inference, 5(1):71–78, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375881900525.

Kounias:1983:SOW

[KC83] Stratis Kounias and Th. Chadjipantelis. Some d-optimal weighing designs for n ≡ (mod 4). Journal of Statistical Planning and Inference, 8(1):117–127, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900666.

Kestenband:1986:BIB

[Kes86] Barbu C. Kestenband. Balanced incomplete block designs on q + 1 elements. Journal of Statistical Planning and Inference, 13(??): 45–50, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901187.

Kounias:1984:COW

[KF84] Stratis Kounias and Nikos Farmakis. A construction of d-optimal weighing designs when n□3 mod 4. Journal of Statistical Planning and Inference, 10(2):177–187, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900697.

Konishi:1989:TES

[KG89] Sadanori Konishi and Arjun K. Gupta. Testing the equality of several intraclass correlation coefficients. Journal of Statistical Planning and Inference, 21(1):93–105, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900220.

Kumar:1985:IPP

[KGN85] Pranesh Kumar, V. K. Gupta, and A. K. Nigam. On inclusion probability proportional to size sampling scheme. Journal of Statistical Planning and Inference, 12(??):127–131, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900618.

Kageyama:1983:FDP

[KH83] S. Kageyama and A. Hedayat. The family of t-designs — part II. Journal of Statistical Planning and Inference, 7(3):257– 287, March 1983. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375883900198.

Khatri:1980:UTC

[Kha80] C. G. Khatri. Unified treatment of Cramér-Rao bound for the nonregular density functions. Journal of Statistical Planning and Inference, 4(1):75–79, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588090035X.

Khan:1981:NAE

[Kha81] Rasul A. Khan. A note on the asymptotic efficiency of the sobelwald test. Journal of Statistical Planning and Inference, 5(3): 243–251, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900045.

Khatri:1982:NVB

[Kha82] C. G. Khatri. A note on variance balanced designs. Journal of Statistical Planning and Inference, 6(2):173–177, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900866.

Khan:1984:UEB

[Kha84] Rasul A. Khan. On UMVU estimators and Bhattacharya bounds in exponential distributions. Journal of Statistical Planning and Inference, 9(2):199–206, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900211.

Khan:1988:OBE

[Kha88] M. K. Khan. Optimal Bayesian estimation of the median effective dose. Journal of Statistical Planning and Inference, 18(1): 69–81, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900171.

Khuri:1989:TCM

[Khu89] A. I. Khuri. Testing a covariance matrix structure in a mixed model with no empty cells. Journal of Statistical Planning and Inference, 22(1):117–125, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900700.

Kuwada:1988:BFF

[KK88] Masahide Kuwada and Shinji Kuriki. Balanced fractional factorial designs of resolution 2l+1 for interesting effects orthogonal to some nuisance parameters: 2_{1}^{m} + m_{2} series. Journal of Statistical Planning and Inference, 19(3):373–382, July 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900444.

Kimura:1989:MID

[KK89] Miyoshi Kimura and Itsuro Kakiuchi. A majorization inequality for distributions on hyperplanes and its applications to tests for outliers. Journal of Statistical Planning and Inference, 21(1):19– 26, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900153.

Kapadia:1988:MSS

[KKL88] C. H. Kapadia, Alan H. Kvanli, and Kwan R. Lee. Minimal sufficient statistics for incomplete block designs with interaction under an Eisenhart model III. Journal of Statistical Planning and Inference, 19(3):317–324, July 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900390.

Klaassen:1989:ASE

[Kla89] Chris A. J. Klaassen. The asymptotic spread of estimators. Journal of Statistical Planning and Inference, 23(3):267–285, November 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900724.

Kreweras:1986:NEP

[KM86] G. Kreweras and P. Moszkowski. A new enumerative property of the Narayana numbers. Journal of Statistical Planning and Inference, 14(1):63–67, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588690011X.

Khosrovshahi:1988:DRB

[KM88] G. B. Khosrovshahi and E. S. Mahmoodian. BIB(9, 18t, 8t, 4, 3t) designs with repeated blocks. Journal of Statistical Planning and Inference, 18(1):125–131, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900237.

Kolyva-Machera:1989:DOD

[KM89] F. Kolyva-Machera. D-optimality in 3^{k}-designs for N ≡ 1 mod 9 observations. Journal of Statistical Planning and Inference, 22(1): 95–103, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900682.

Kuwada:1988:CPI

[KN88] Masahide Kuwada and Ryuei Nishii. On the characteristic polynomial of the information matrix of balanced fractional s^{m} factorial designs of resolution V_{p,q}. Journal of Statistical Planning and Inference, 18(1):101–114, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900201.

Koske:1989:VFD

[Kos89] J. K. Arap Koske. The variance function of the difference between two estimated fourth order response surfaces. Journal of Statistical Planning and Inference, 23(2):263–266, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900943.

Khattree:1987:SNP

[KP87a] Ravindra Khattree and Shyamal Das Peddada. A short note on Pitman nearness for elliptically symmetric estimators. Journal of Statistical Planning and Inference, 16(??):257–260, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375887900759.

Krzyzak:1987:PRC

[KP87b] Adam Krzyzak and Miroslaw Pawlak. The pointwise rate of convergence of the kernel regression estimate. Journal of Statistical Planning and Inference, 16(??):159–166, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887900656.

Khan:1983:RRB

[KPY83] A. H. Khan, Saud Parvez, and Mohd. Yaqub. Recurrence relations between product moments of order statistics. Journal of Statistical Planning and Inference, 8(2):175–183, October 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375883900368.

Krafft:1985:UOD

[Kra85] Olaf Krafft. On universally optimal designs for regression setups. Journal of Statistical Planning and Inference, 12(??):347– 352, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900825.

Kulperger:1982:RGG

[KS82] R. J. Kulperger and A. C. Singh. On random grouping in goodness of fit tests of discrete distributions. Journal of Statistical Planning and Inference, 7(2):109–115, December 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375882900696.

Katona:1983:MCF

[KS83] Gyula Katona and Jaya Srivastava. Minimal 2-coverings of a finite affine space based on GF(2). Journal of Statistical Planning and Inference, 8(3):375–388, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900538.

Knafl:1982:MRC

[KSY82] G. Knafl, J. Sacks, and D. Ylvisaker. Model robust confidence intervals. Journal of Statistical Planning and Inference, 6(4):319– 334, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900027.

Kageyama:1980:SBB

[KT80] Sanpei Kageyama and Takumi Tsuji. Some bounds on balanced block designs. Journal of Statistical Planning and Inference, 4(2): 155–167, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588090004X.

Kageyama:1981:SFG

[KT81] Sanpei Kageyama and Tatsuzo Tanaka. Some families of group divisible designs. Journal of Statistical Planning and Inference, 5(3):231–241, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375881900033.

Kubokawa:1989:TSP

[Kub89] Tatsuya Kubokawa. Two-stage procedures for parameters in a growth curve model. Journal of Statistical Planning and Inference, 22(1):105–115, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900694.

Kuriki:1988:ESB

[Kur88] Shinji Kuriki. Existence of 2-symbol balanced arrays of strength t and t + 2 constraints. Journal of Statistical Planning and Inference, 20(2):225–228, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888901267.

Kuwada:1981:CPI

[Kuw81] Masahide Kuwada. Characteristic polynomials of the information matrices of balanced fractional 3^{m} factorial designs of resolution v. Journal of Statistical Planning and Inference, 5(2): 189–209, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588190029X.

Kuwada:1982:SOF

[Kuw82] Masahide Kuwada. On some optimal fractional 2^{m} factorial designs of resolution V. Journal of Statistical Planning and Inference, 7(1):39–48, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900180.

Kuwada:1986:BFF

[Kuw86] Masahide Kuwada. Balanced fractional 2_{1}^{m} factorial designs of resolution V for interesting effects orthogonal to some effects concerning m_{2} factors. Journal of Statistical Planning and Inference, 13(??):365–376, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901461.

Kuwada:1987:SPB

[Kuw87] Masahide Kuwada. On some partially balanced fractional 2^{m_{1}+m_{2}} factorial designs of resolution VI^{−}. Journal of Statistical Planning and Inference, 17(??):93–102, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887901030.

Kuwada:1988:OPB

[Kuw88] Masahide Kuwada. A-optimal partially balanced fractional 2_{1}^{m} + m_{2} factorial designs of resolution V, with 4 ≤ m_{1>} + m_{2} ≤ 6. Journal of Statistical Planning and Inference, 18(2):177–193, February 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900043.

Kariya:1985:NPS

[KW85] Takeaki Kariya and C. F. J. Wu. On the nonsingularity of principal submatrices of a random orthogonal matrix. Journal of Statistical Planning and Inference, 12(??):353–357, ????

1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900837.

Koronacki:1988:GSR

[KW88] Jacek Koronacki and Wolfgang Wertz. A global stopping rule for recursive density estimators. Journal of Statistical Planning and Inference, 20(1):23–39, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588890081X.

Karakostas:1989:SSA

[KW89] K. X. Karakostas and H. P. Wynn. Systematic sampling for autocorrelated superpopulations. Journal of Statistical Planning and Inference, 22(2):181–195, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889901092.

Khan:1988:ODB

[KY88] M. K. Khan and A. A. Yazdi. On d-optimal designs for binary data. Journal of Statistical Planning and Inference, 18(1):83– 91, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900183.

LaRiccia:1986:OGF

[LaR86] Vincent N. LaRiccia. Optimal goodness-of-fit tests for normality against skewness and kurtosis alternatives. Journal of Statistical Planning and Inference, 13(??):67–79, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901205.

Lau:1988:ODU

[Lau88] Tai-Shing Lau. D-optimal designs on the unit q-ball. Journal of Statistical Planning and Inference, 19(3):299–315, July 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900389.

Lawrence:1988:AAC

[Law88] John D. Lawrence. Automating argument construction. Journal of Statistical Planning and Inference, 20(3):369–387, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900985.

Lu:1989:ECP

[LB89] K. L. Lu and James O. Berger. Estimated confidence procedures for multivariate normal means. Journal of Statistical Planning and Inference, 23(1):1–19, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900359.

Lee:1988:COD

[Lee88] Carl M.-S. Lee. Constrained optimal designs. Journal of Statistical Planning and Inference, 18(3):377–389, March 1988. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375888901140.

Lewis:1982:GAF

[Lew82] S. M. Lewis. Generators for asymmetrical factorial experiments. Journal of Statistical Planning and Inference, 6(1):59– 64, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900568.

Lu:1988:CID

[LGB88] Tai-Fang C. Lu, Franklin A. Graybill, and Richard K. Burdick. Confidence intervals on a difference of expected mean squares. Journal of Statistical Planning and Inference, 18(1):35– 43, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900146.

Lu:1989:CIR

[LGB89] Tai-Fang C. Lu, Franklin A. Graybill, and Richard K. Burdick. Confidence intervals on the ratio of expected mean squares (θ_{1} − dθ_{2})θ_{3}. Journal of Statistical Planning and Inference, 21(2):179– 190, February 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900037.

Liang:1987:CMF

[Lia87] W. Liang. Consistency of modified Fisher prediction function. Journal of Statistical Planning and Inference, 17(??):401– 403, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887901303.

Liang:1989:CRE

[Lia89] TaChen Liang. Convergence rates for empirical Bayes estimation of binomial parameter. Journal of Statistical Planning and Inference, 21(1):117–124, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900244.

Lin:1988:CDR

[Lin88] Gwo Dong Lin. Characterizations of distributions via relationships between two moments of order statistics. Journal of Statistical Planning and Inference, 19(1):73–80, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900535.

Lin:1989:PMO

[Lin89] Gwo Dong Lin. The product moments of order statistics with applications to characterizations of distributions. Journal of Statistical Planning and Inference, 21(3):395–406, March 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900566.

Lee:1987:COG

[LJ87] K. Y. Lee and M. Jacroux. On the construction of E- and MV - optimal group divisible designs with unequal block sizes. Journal of Statistical Planning and Inference, 16(??):193–201, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375887900681.

Lairson:1980:DIM

[LKM80] David R. Lairson, Joseph Krislov, and James R. Marsden. The demand to initiate medical care: a morbidity specific analysis. Journal of Statistical Planning and Inference, 4(1):99– 113, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375880900385.

Liebler:1988:PGL

[LM88] R. A. Liebler and A. D. Meyerowitz. Partial geometric lattices, II: association schemes. Journal of Statistical Planning and Inference, 18(2):161–176, February 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900031.

Li:1982:RDN

[LN82] K. C. Li and W. Notz. Robust designs for nearly linear regression. Journal of Statistical Planning and Inference, 6(2):135– 151, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900830.

Lo:1989:NAN

[Lo89] Gane Samb Lo. A note on the asymptotic normality of sums of extreme values. Journal of Statistical Planning and Inference, 22(1): 127–136, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900712.

Loh:1985:CMV

[Loh85] Wei-Yin Loh. The Cauchy mean value property for M-estimates. Journal of Statistical Planning and Inference, 12(??):265–267, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900758.

Longyear:1981:SND

[Lon81] J. Q. Longyear. A survey of nested designs. Journal of Statistical Planning and Inference, 5(2):181–187, ???? 1981. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375881900288.

Longyear:1984:AST

[Lon84] J. Q. Longyear. Arrays of strength s on two symbols. Journal of Statistical Planning and Inference, 10(2):227–239, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375884900752.

Longyear:1986:NGD

[Lon86] Judith Q. Longyear. Nested group divisible designs and small nested designs. Journal of Statistical Planning and Inference, 13 (??):81–87, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901217.

Levy:1986:NCM

[LP86] M. S. Levy and S. K. Perng. A note on the consistency of the modified Fisher prediction function. Journal of Statistical Planning and Inference, 13(??):263–267, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901382.

Lea:1988:APL

[LP88a] Cheun-Der Lea and Madan L. Puri. Asymptotic properties of linear functions of order statistics. Journal of Statistical Planning and Inference, 18(2):203–223, February 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900067.

Lea:1988:APP

[LP88b] Cheun-Der Lea and Madan L. Puri. Asymptotic properties of perturbed empirical distribution functions evaluated at a random point. Journal of Statistical Planning and Inference, 19(2):201– 215, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900730.

Lio:1986:APK

[LPY86] Y. L. Lio, W. J. Padgett, and K. F. Yu. On the asymptotic properties of a kernel type quantile estimator from censored samples. Journal of Statistical Planning and Inference, 14(2–3):169– 177, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901540.

Ludman:1981:TGG

[LS81] J. E. Ludman and J. L. Sampson. A technique for generating Gray codes. Journal of Statistical Planning and Inference, 5(2): 171–180, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900276.

Lam:1984:GBR

[LS84] Clement Lam and Jennifer Seberry. Generalized Bhaskar Rao designs. Journal of Statistical Planning and Inference, 10(1): 83–95, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588490034X.

Mack:1983:RSU

[Mac83] Y. P. Mack. Rate of strong uniform convergence of k-NN density estimates. Journal of Statistical Planning and Inference, 8 (2):185–192, October 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588390037X.

Majumdar:1986:ODC

[Maj86] Dibyen Majumdar. Optimal designs for comparisons between two sets of treatments. Journal of Statistical Planning and Inference, 14(2–3):359–372, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901734.

Mandal:1989:ODE

[Man89] Nripes Kumar Mandal. D-optimal designs for estimating the optimum point in a quadratic response surface — rectangular region. Journal of Statistical Planning and Inference, 23(2):243– 252, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588990092X.

Martinsek:1987:TSF

[Mar87] Adam T. Martinsek. Two-stage and fully sequential determination of estimator as well as sample size. Journal of Statistical Planning and Inference, 17(??):11–20, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588790098X.

Masaro:1988:OBM

[Mas88] Joseph C. Masaro. On A-optimal block matrices and weighing designs when N ≡ 3 mod 4. Journal of Statistical Planning and Inference, 18(3):363–370, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888901127.

Mau:1985:SMP

[Mau85] J. Mau. Statistical modelling via partitioned counting processes. Journal of Statistical Planning and Inference, 12(??):171– 176, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900667.

Maymin:1987:CBP

[May87] Zakhar Maymin. On a conjecture of Barlow and Proschan concerning reliability bounds. Journal of Statistical Planning and Inference, 16(??):337–344, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900863.

Mathew:1983:OBG

[MB83] Thomas Mathew and P. Bhimasankaram. Optimality of blue’s in a general linear model with incorrect design matrix. Journal of Statistical Planning and Inference, 8(3):315–329, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375883900484.

Mielke:1985:NAI

[MB85] Paul W. Mielke and Kenneth J. Berry. Non-asymptotic inferences based on the chi-square statistic for r by c contingency tables. Journal of Statistical Planning and Inference, 12(??):41– 45, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900515.

Medhi:1986:GFP

[MB86] J. Medhi and M. Borah. On generalized four-parameter Charlier distribution. Journal of Statistical Planning and Inference, 14(1): 69–77, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900121.

Mehrotra:1987:EGL

[MB87] Kishan G. Mehrotra and Gouri K. Bhattacharyya. Estimation in a Gaussian linear regression model under type II censoring. Journal of Statistical Planning and Inference, 17(??):241– 250, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887901169.

Mukerjee:1988:NEK

[MB88] Rahul Mukerjee and Mausumi Bose. Non-equireplicate Kronecker factorial designs. Journal of Statistical Planning and Inference, 19(2):261–267, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588890078X.

Mathew:1989:MRO

[MB89] Thomas Mathew and Dulal Kumar Bhaumik. The model-robustness and optimality of randomized designs. Journal of Statistical Planning and Inference, 23(3):371–379, November 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375889900797.

McCulloch:1986:SRA

[McC86] Robert E. McCulloch. Some remarks on allocatory and separatory linear discrimination. Journal of Statistical Planning and Inference, 14(2–3):323–330, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901709.

Mendelson:1986:LPC

[Men86] Haim Mendelson. Lattice path combinatorics and linear probing. Journal of Statistical Planning and Inference, 14(1):79– 93, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900133.

Meslem:1987:AEF

[Mes87] Abd-El-Hakim Meslem. Asymptotic expansions for fixed width confidence interval. Journal of Statistical Planning and Inference, 17 (??):51–65, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887901005.

McAllister:1984:UEE

[MH84] Paul R. McAllister and Burt S. Holland. Use of the experimentwise error rate in simultaneous tests of main effects in two-way disproportionate ANOVA without interaction. Journal of Statistical Planning and Inference, 9(2):155–162, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900168.

Mielke:1982:BRS

[Mie82] Paul W. Mielke. Book review: Statistical climatology: S. Ikeda, E. Suzuki, E. Uchida and M. M. Yoshino, editors (Elsevier Scientific Publishing Company, New York, 1980) x + 388 pp. $68.25. Journal of Statistical Planning and Inference, 6(2):201– 202, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900891.

Mielke:1986:NMS

[Mie86] Paul W. Mielke. Non-metric statistical analyses: Some metric alternatives. Journal of Statistical Planning and Inference, 13(??): 377–387, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901473.

Mielke:1987:RMT

[Mie87] Paul W. Mielke. F_{9}. L_{1}, L_{2} and L_{∞} regression models: Is there a difference? Journal of Statistical Planning and Inference, 16 (??):430, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887900942.

Mikaeili:1989:DOD

[Mik89] Fattah Mikaeili. D-optimum design for cubic without 3-way effect on the simplex. Journal of Statistical Planning and Inference, 21 (1):107–115, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900232.

Milbrodt:1988:LEA

[Mil88] Hartmut Milbrodt. Limits of experiments associated with sampling plans. Journal of Statistical Planning and Inference, 19(1): 1–29, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900493.

Moyssiadis:1982:EOF

[MK82] Chronis Moyssiadis and Stratis Kounias. The exact D-optimal first order saturated design with 17 observations. Journal of Statistical Planning and Inference, 7(1):13–27, October 1982. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375882900155.

Mandeli:1981:COS

[MLF81] John P. Mandeli, F.-C. Helen Lee, and Walter T. Federer. On the construction of orthogonal F -squares of order n from an orthogonal array (n, k, s, 2) and an OL(s, t) set. Journal of Statistical Planning and Inference, 5(3):267–272, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375881900069.

Meyerowitz:1987:PGL

[MM87] A. Meyerowitz and R. Miskimins. Partial geometric lattices, i: Regularity conditions. Journal of Statistical Planning and Inference, 17(??):21–50, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900991.

Mavron:1984:FPG

[MMR84] V. C. Mavron, R. C. Mullin, and A. Rosa. Further properties of generalized residual designs. Journal of Statistical Planning and Inference, 10(1):59–68, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900326.

Mohan:1981:CCT

[Moh81] N. R. Mohan. A criterion for connectedness in two-associate class PBIB designs. Journal of Statistical Planning and Inference, 5(2): 211–212, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900306.

Mohanty:1986:P

[Moh86] S. G. Mohanty. Preface. Journal of Statistical Planning and Inference, 14(1):3–4, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886900030.

Montakhab:1986:EFP

[Mon86] M. S. Montakhab. Embedding of finite pseudo-complements of quadrilaterals. Journal of Statistical Planning and Inference, 13 (??):103–110, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901242.

Moore:1984:MLF

[Moo84] David S. Moore. Measures of lack of fit from tests of chi-squared type. Journal of Statistical Planning and Inference, 10(2):151– 166, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900673.

Moore:1988:SFD

[Moo88] L. M. Moore. Singular factorial designs resulting from missing pairs of design points. Journal of Statistical Planning and Inference, 19(3):325–340, July 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900407. See comment

[Gho89b].

Mori:1983:NCR

[Mór83] T. F. Móri. Note on the Cramér-Rao inequality in the nonregular case: the family of uniform distributions. Journal of Statistical Planning and Inference, 7(4):353–358, June 1983. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837588390006X.

Morris:1984:MNR

[Mor84] Max D. Morris. Minimum number of runs for two-level factorial search designs. Journal of Statistical Planning and Inference, 10 (1):115–117, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900375.

McNichols:1988:ISS

[MP88] D. T. McNichols and W. J. Padgett. Inference for step-stress accelerated life tests under arbitrary right-censorship. Journal of Statistical Planning and Inference, 20(2):169–179, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888901206.

Menon:1984:NPR

[MPS84] V. V. Menon, B. Prasad, and R. S. Singh. Non-parametric recursive estimates of a probability density function and its derivatives. Journal of Statistical Planning and Inference, 9(1):73– 82, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900466.

Mukerjee:1985:AOW

[MR85] Rahul Mukerjee and Rita Saha Ray. Asymptotically optimal weighing designs with string property. Journal of Statistical Planning and Inference, 12(??):87–91, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900564.

Marrero:1989:OOB

[MR89] Osvaldo Marrero and Damaraju Raghavarao. Overall A-optimal balanced incomplete block designs. Journal of Statistical Planning and Inference, 21(1):125–127, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900256.

Mielniczuk:1989:LDD

[MSV89] Jan Mielniczuk, Pascal Sarda, and Philippe Vieu. Local datadriven bandwidth choice for density estimation. Journal of Statistical Planning and Inference, 23(1):53–69, September 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375889900396.

Mukerjee:1981:CEW

[Muk81] Rahul Mukerjee. Construction of effect-wise orthogonal factorial designs. Journal of Statistical Planning and Inference, 5(3):221– 229, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900021.

Mukhopadhyay:1982:OSE

[Muk82] N. Mukhopadhyay. Optimal sequential estimation procedures for the normal mean when the variance is known. Journal of Statistical Planning and Inference, 6(2):131–134, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900829.

Mukhopadhyay:1984:STS

[Muk84] N. Mukhopadhyay. Sequential and two-stage procedures for selecting the better exponential population covering the case of unknown and unequal scale parameters. Journal of Statistical Planning and Inference, 9(1):33–43, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900429.

Mukhopadhyay:1987:TSP

[Muk87] Nitis Mukhopadhyay. Three-stage procedures for selecting the best exponential population. Journal of Statistical Planning and Inference, 16(??):345–352, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900875.

Mustafi:1987:RMN

[Mus87] C. K. Mustafi. Residual median for non-identically distributed random variables. Journal of Statistical Planning and Inference, 17(??):389–400, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901297.

Miller:1987:VFE

[MW87] John J. Miller and Edward J. Wegman. Vector function estimation using splines. Journal of Statistical Planning and Inference, 17(??):173–180, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901108.

Nabeya:1986:CLR

[Nab86] Seiji Nabeya. On a condition for the log-rank test to be a locally optimal rank test. Journal of Statistical Planning and Inference, 13(??):403–405, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901515.

Nachtsheim:1989:DEP

[Nac89] Christopher J. Nachtsheim. On the design of experiments in the presence of fixed covariates. Journal of Statistical Planning and Inference, 22(2):203–212, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889901110.

Nagaraja:1982:NMS

[Nag82] H. N. Nagaraja. On the non-Markovian structure of discrete order statistics. Journal of Statistical Planning and Inference, 7(1): 29–33, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900167.

Nagaraja:1986:NCM

[Nag86a] H. N. Nagaraja. A note on the conditional Markov property of discrete order statistics. Journal of Statistical Planning and Inference, 13(??):37–43, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901175.

Nagaraja:1986:SDO

[Nag86b] H. N. Nagaraja. Structure of discrete order statistics. Journal of Statistical Planning and Inference, 13(??):165–177, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886901308.

Nagaraja:1988:SCD

[Nag88] H. N. Nagaraja. Some characterizations of discrete distributions based on linear regressions of adjacent order statistics. Journal of Statistical Planning and Inference, 20(1):65–75, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900833.

Nair:1982:ECM

[Nai82] K. Aiyappan Nair. An estimator of the common mean of two normal populations. Journal of Statistical Planning and Inference, 6(2):119–122, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900799.

Nigam:1985:IBD

[NB85] A. K. Nigam and G. M. Boopathy. Incomplete block designs for symmetrical parallel line assays. Journal of Statistical Planning and Inference, 11(1):111–117, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588590031X.

Niederhausen:1986:ERR

[Nie86] Heinrich Niederhausen. The enumeration of restricted random walks by Sheffer polynomials with applications to statistics. Journal of Statistical Planning and Inference, 14(1):95– 114, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900145.

Notz:1981:SDM

[Not81] W. Notz. Saturated designs for multivariate cubic regression. Journal of Statistical Planning and Inference, 5(1):37–43, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375881900495.

Notz:1982:MPS

[Not82] W. Notz. Minimal point second order designs. Journal of Statistical Planning and Inference, 6(1):47–58, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375882900556.

Notz:1985:ODT

[Not85] William I. Notz. Optimal designs for treatment-control comparisons in the presence of two-way heterogeneity. Journal of Statistical Planning and Inference, 12(??):61–73, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900539.

Notz:1989:ODR

[Not89] William I. Notz. Optimal designs for regression models with possible bias. Journal of Statistical Planning and Inference, 22(1): 43–54, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900633.

Nowicki:1989:ANG

[Now89] Krzysztof Nowicki. Asymptotic normality of graph statistics. Journal of Statistical Planning and Inference, 21(2):209–222, February 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900050.

Nath:1985:SDD

[NS85] Prem Nath and Ranjit Singh. On symmetry and the directed divergence in information theory. Journal of Statistical Planning and Inference, 12(??):235–248, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900722.

Nishii:1986:MPT

[NS86] Ryuei Nishii and Teruhiro Shirakura. More precise tables of Srivastava–Chopra balanced optimal 2^{m} fractional factorial designs of resolution V, m ≤ 6. Journal of Statistical Planning and Inference, 13(??):111–116, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901254.

Nagaraja:1987:SCG

[NS87a] H. N. Nagaraja and R. C. Srivastava. Some characterizations of geometric type distributions based on order statistics. Journal of Statistical Planning and Inference, 17(??):181–191, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837588790111X.

Nath:1987:QSD

[NS87b] Prem Nath and Ranjit Singh. Quasicyclic symmetry and the directed divergence in information theory. Journal of Statistical Planning and Inference, 16(??):33–47, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588790053X.

OReilly:1984:ELN

[OC84] Federico J. O’Reilly and Alberto Castillo. Estimability and linearity in normal regression. Journal of Statistical Planning and Inference, 9(2):195–198, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588490020X.

Oman:1983:CMS

[Oma83] Samuel D. Oman. A class of modified Stein estimators with easily computable risk functions. Journal of Statistical Planning and Inference, 7(4):359–369, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900071.

Oman:1985:CEF

[Oma85a] S. D. Oman. Corrigendum to “An exact formula for the mean squared error of the inverse estimator in the linear calibration problem”: (journal of statistical planning and inference 11 (1985) 189–196). Journal of Statistical Planning and Inference, 12(??): 401, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900898. See [Oma85b].

Oman:1985:EFM

[Oma85b] Samuel D. Oman. An exact formula for the mean squared error of the inverse estimator in the linear calibration problem. Journal of Statistical Planning and Inference, 11(2):189–196, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900059. See corrigendum [Oma85a].

Ohnishi:1985:SDF

[OS85] Toshio Ohnishi and Teruhiro Shirakura. Search designs for 2^{m} factorial experiments. Journal of Statistical Planning and Inference, 11(2):241–245, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900114.

Ouyang:1985:ADE

[OV85] Soo Peter Ouyang and John Van Ryzin. Asymptotic distributions of estimators of the derivative of a density and the score function. Journal of Statistical Planning and Inference, 12(??): 11–25, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900497.

Ovchinnikov:1988:SRF

[Ovc88] Sergei V. Ovchinnikov. Survey of research on fuzzy sets in the Soviet union. Journal of Statistical Planning and Inference, 20(3):

413–428, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888901000.

OReilly:1984:MDM

[OZ84] Federico J. O’Reilly and Guillermo P. Zárate. On the mvue for the distribution in the multivariate normal regression. Journal of Statistical Planning and Inference, 10(3):273–276, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375884900533.

Payne:1981:SCC

[Pay81] R. W. Payne. Selection criteria for the construction of efficient diagnostic keys. Journal of Statistical Planning and Inference, 5 (1):27–36, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900483.

Pazman:1989:IMN

[Páz89] Andrej Pázman. On information matrices in nonlinear experimental design. Journal of Statistical Planning and Inference, 21 (2):253–263, February 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900098.

Pollak:1981:CIS

[PC81] Moshe Pollak and Jiska Cohen. A comparison of the independentsamples t-test and the paired-samples t-test when the observations are nonnegatively correlated pairs. Journal of Statistical Planning and Inference, 5(2):133–146, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375881900239.

Peixoto:1988:PCM

[Pei88] Julio L. Peixoto. Prediction in a class of mixed models with two variance components. Journal of Statistical Planning and Inference, 19(1):81–94, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900547.

Pesotchinsky:1984:RDO

[Pes84] L. Pesotchinsky. Robust designs and optimality of least squares for regression problems. Journal of Statistical Planning and Inference, 9(1):103–117, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588490048X.

Pfanzagl:1988:CML

[Pfa88] J. Pfanzagl. Consistency of maximum likelihood estimators for certain nonparametric families, in particular: mixtures. Journal of Statistical Planning and Inference, 19(2):137–158, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900699.

Puri:1988:IBD

[PG88] P. D. Puri and L. R. Gupta. Incomplete block designs for parallel line assays. Journal of Statistical Planning and Inference, 20(1):121–128, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900882.

Pilz:1986:MLR

[Pil86] Jürgen Pilz. Minimax linear regression estimation with symmetric parameter restrictions. Journal of Statistical Planning and Inference, 13(??):297–318, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901412.

Park:1983:MCS

[PK83] C. J. Park and K. S. Kang. A modified cluster sampling method for estimating the bacterial density in Xanthan gum. Journal of Statistical Planning and Inference, 7(4):317–327, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375883900034.

Peddada:1988:EMS

[PL88] Shyamal Das Peddada and Parthasarathi Lahiri. The exact mean squared error of Stein-rule estimator in linear models. Journal of Statistical Planning and Inference, 18(3):345–353, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888901103.

Peixoto:1989:SIO

[PL89] Julio L. Peixoto and Lynn R. LaMotte. Simultaneous identification of outliers and predictors using variable selection techniques. Journal of Statistical Planning and Inference, 23(3):327– 343, November 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900761.

Pourahmadi:1987:SSP

[Pou87] Moshen Pourahmadi. Some sampling properties of empirical characteristic functions viewed as harmonizable stochastic processes. Journal of Statistical Planning and Inference, 17(??):345– 359, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588790125X.

Preitschopf:1987:ODQ

[PP87] Franz Preitschopf and Friedrich Pukelsheim. Optimal designs for quadratic regression. Journal of Statistical Planning and Inference, 16(??):213–218, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588790070X.

Pesotan:1985:SDO

[PR85] H. Pesotan and B. L. Raktoe. Some determinant optimality aspects of main effect fold-over designs. Journal of Statistical Planning and Inference, 11(3):399–410, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588590045X.

Pesotan:1988:IRF

[PR88] H. Pesotan and B. L. Raktoe. On invariance and randomization in factorial designs with applications to D-optimal main effect designs of the symmetrical factorial. Journal of Statistical Planning and Inference, 19(3):283–298, July 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900377.

Pena:1989:SFE

[PR89] Edsel A. Peña and Vijay K. Rohatgi. Survival function estimation for a generalized proportional hazards model of random censorship. Journal of Statistical Planning and Inference, 22(3): 371–389, July 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889901018.

Prasad:1985:EUA

[Pra85] B. Prasad. Equivalence of uniform asymptotic unbiasedness, mean square and strong consistencies of recursive estimates of a density and its p-th derivative. Journal of Statistical Planning and Inference, 12(??):81–86, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900552.

Prentice:1982:ASD

[Pre82] Michael J. Prentice. Antipodally symmetric distributions for orientation statistics. Journal of Statistical Planning and Inference, 6(3):205–214, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900258.

Press:1987:MCI

[Pre87] S. James Press. The MISER criterion for imbalance in the analysis of covariance. Journal of Statistical Planning and Inference, 17(??):375–388, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901285.

Pesotan:1983:IRU

[PRJ83] H. Pesotan, B. L. Raktoe, and J. Joiner. On invariance and randomization under factor permutation in fractional factorial designs. Journal of Statistical Planning and Inference, 8(2):209– 221, October 1983. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375883900393.

Puri:1984:EESa

[PS84a] Madan L. Puri and Munsup Seoh. Edgeworth expansion for signed linear rank statistics with regression constants. Journal of Statistical Planning and Inference, 10(2):137–149, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375884900661.

Puri:1984:EESb

[PS84b] Madan L. Puri and Munsup Seoh. Edgeworth expansions for signed linear rank statistics under near location alternatives. Journal of Statistical Planning and Inference, 10(3):289–309, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375884900557.

Pace:1987:CTE

[PS87] Luigi Pace and Alessandra Salvan. A class of tests for exponentiality against monotone failure rate alternatives with incomplete data. Journal of Statistical Planning and Inference, 16(??): 249–256, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887900747.

Parihar:1988:MCG

[PS88] J. S. Parihar and R. Shrivastava. Methods of construction of group divisible designs. Journal of Statistical Planning and Inference, 18(3):399–404, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888901164.

Pathak:1987:EBE

[PSSZ87] Pramod K. Pathak, Anita Singh, Ashok K. Singh, and William J. Zimmer. Empirical Bayesian estimation of mean life from an accelerated life test. Journal of Statistical Planning and Inference, 16(??):353–363, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900887.

Pukelsheim:1980:LRD

[Puk80] Friedrich Pukelsheim. On linear regression designs which maximize information. Journal of Statistical Planning and Inference, 4(4):339–364, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375880900208.

Pukelsheim:1983:OPS

[Puk83] Friedrich Pukelsheim. On optimality properties of simple block designs in the approximate design theory. Journal of Statistical Planning and Inference, 8(2):193–208, October 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900381.

Puri:1986:ARS

[Pur86] Madan L. Puri. Acknowledgement of referee’s services. Journal of Statistical Planning and Inference, 14(2–3):409, ???? 1986. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375886901771.

Puri:1989:ARS

[Pur89] Madan L. Puri. Acknowledgement of referees’ services. Journal of Statistical Planning and Inference, 23(3):403–404, November 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900827.

Puri:1985:GAS

[PW85] Madan L. Puri and Tiee-Jian Wu. Gaussian approximation of signed rank statistics process. Journal of Statistical Planning and Inference, 11(3):277–312, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900357.

Ramachandran:1982:HTE

[Ram82] Geetha Ramachandran. Horvitz–Thompson estimator and generalized πPS designs. Journal of Statistical Planning and Inference, 7 (2):151–153, December 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900738.

Rao:1981:NUR

[Rao81] T. J. Rao. A note on unbiasedness in ratio estimation. Journal of Statistical Planning and Inference, 5(4):335–340, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375881900331.

Rao:1983:HTS

[Rao83] T. J. Rao. Horvitz–Thompson strategy vs. stratified random sampling strategy. Journal of Statistical Planning and Inference, 8 (1):43–50, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900605.

Rayner:1985:ECT

[RB85] J. C. W. Rayner and D. J. Best. The effect of categorisation on testing for an exponential scale shift. Journal of Statistical Planning and Inference, 11(2):197–206, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900060.

Rust:1984:STS

[RF84] Steven W. Rust and Michael A. Fligner. On the studentization of two-sample linear rank statistics. Journal of Statistical Planning and Inference, 9(3):345–353, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900090.

Raghavarao:1986:CDA

[RFS86] D. Raghavarao, W. T. Federer, and S. J. Schwager. Characteristics for distinguishing among balanced incomplete block designs with repeated blocks. Journal of Statistical Planning and Inference, 13(??):151–163, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375886901291.

Rao:1981:DFT

[RG81] K. S. Madhava Rao and A. P. Gore. Distribution-free tests for parallelism and concurrence in two-sample regression problem. Journal of Statistical Planning and Inference, 5(3):281– 286, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900082.

Risko:1985:SBB

[Ris85] Kenneth J. Risko. Selecting the better binomial population under a budget constraint. Journal of Statistical Planning and Inference, 11(3):341–353, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900394.

Rosenblatt:1983:UAP

[RJDT83] Judah Rosenblatt, D. L. Jackson, W. P. Dole, and W. L. Thompson. On uniform approximation of Poisson variates by normal ones. Journal of Statistical Planning and Inference, 8(3):281– 300, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375883900460.

Richards:1987:GMD

[RM87] Dale O. Richards and James B. McDonald. A general methodology for determining distributional forms with applications in reliability. Journal of Statistical Planning and Inference, 16(??): 365–376, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887900899.

Rao:1988:ASI

[RM88] B. Raja Rao and Gary M. Marsh. Approximate simultaneous inferential procedures for overall risk assessment of several competing causes in biomedical and epidemiologic studies. Journal of Statistical Planning and Inference, 18(3):323–343, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888901097.

Rao:1985:STS

[RMW85] B. Raja Rao, Gary M. Marsh, and Joan Winwood. Sidak-type simultaneous prediction intervals for the mortality measures RSRR_{i} about the corresponding S_{e}PMR_{i} for several competing risks of death in an epidemiologic study. Journal of Statistical Planning and Inference, 12(??):311–329, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900801.

Robinson:1985:AOA

[Rob85] P. M. Robinson. Approximate optimal allocation in repeated sampling from a finite population. Journal of Statistical Planning and Inference, 11(2):135–148, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900011.

Roebruck:1987:ODT

[Roe87] P. Roebruck. Optimal design for three-factorial split-plot models with random block effects. Journal of Statistical Planning and Inference, 17(??):293–305, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901212.

Rohatgi:1988:BRS

[Roh88] Vijay K. Rohatgi. Book review: Statistical discussion forum: Edited by I. J. Good VPI&SU, Blacksburg, VA 24061. Journal of Statistical Planning and Inference, 19(1):133–136, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900584.

Rokowska:1984:RST

[Rok84] Barbara Rokowska. Resolvable systems of 8-tuples. Journal of Statistical Planning and Inference, 9(1):131–141, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375884900508.

Rothschild:1986:DED

[Rot86] Lord Rothschild. The distribution of English dictionary word lengths. Journal of Statistical Planning and Inference, 14(2–3): 311–322, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901692.

Roussas:1988:NEM

[Rou88] George G. Roussas. Nonparametric estimation in mixing sequences of random variables. Journal of Statistical Planning and Inference, 18(2):135–149, February 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900018.

Roussas:1989:HRE

[Rou89] George G. Roussas. Hazard rate estimation under dependence conditions. Journal of Statistical Planning and Inference, 22(1): 81–93, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900670.

Roy:1982:CMO

[Roy82] Bimal K. Roy. Construction of (m,s)-optimal design for block size 3. Journal of Statistical Planning and Inference, 7(1):35– 37, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900179.

Roy:1988:CSB

[Roy88] B. K. Roy. Construction of strongly balanced uniform repeated measurements designs. Journal of Statistical Planning and Inference, 19(3):341–348, July 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900419.

Roy:1984:OCM

[RS84] B. K. Roy and K. R. Shah. On the optimality of a class of minimal covering designs. Journal of Statistical Planning and Inference, 10(2):189–194, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900703.

Rukhin:1983:IAP

[Ruk83] Andrew L. Rukhin. Inference about permutation parameter in large samples. Journal of Statistical Planning and Inference, 8 (1):59–74, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900629.

Rukhin:1988:IEL

[Ruk88] Andrew L. Rukhin. Improved estimation in lognormal regression models. Journal of Statistical Planning and Inference, 18(3):291– 297, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888901061.

Ruschendorf:1985:TRO

[Rüs85] Ludger Rüschendorf. Two remarks on order statistics. Journal of Statistical Planning and Inference, 11(1):71–74, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900254.

Ruymgaart:1989:SUC

[Ruy89] F. H. Ruymgaart. Strong uniform convergence of density estimators on spheres. Journal of Statistical Planning and Inference, 23 (1):45–52, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900384.

Sackrowitz:1982:PIM

[Sac82] Harold Sackrowitz. Procedures for improving the mle for ordered binomial parameters. Journal of Statistical Planning and Inference, 6(3):287–296, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900325.

Sahai:1980:IVB

[Sah80] Ashok Sahai. Improved variance bounds for unbiased estimation in inverse sampling. Journal of Statistical Planning and Inference, 4(3):213–216, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375880900075.

Srivastava:1984:TFD

[SAM84] Jaya Srivastava, Donald Anderson, and Jacy Mardekian. Theory of factorial designs of the parallel flats type. I: the coefficient matrix. Journal of Statistical Planning and Inference, 9(2):229– 252, March 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900235.

Santner:1981:DTF

[San81] Thomas J. Santner. Designing two factor experiments for selecting interactions. Journal of Statistical Planning and Inference, 5 (1):45–55, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900501.

Sane:1984:ASM

[San84] Sharad S. Sane. Auxiliary sets of matrices with non-increasing step parameter sequences. Journal of Statistical Planning and Inference, 10(3):375–384, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900612.

Sarndal:1982:ISD

[Sär82] Carl Erik Särndal. Implications of survey design for generalized regression estimation of linear functions. Journal of Statistical Planning and Inference, 7(2):155–170, December 1982. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837588290074X.

Sastry:1988:CRI

[Sas88] D. V. S. Sastry. A class of resolvable incomplete block designs. Journal of Statistical Planning and Inference, 19(1):117– 120, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900560.

Saleh:1983:EQF

[SAU83] A. K. Md. Ehsanes Saleh, M. Masoom Ali, and Dale Umbach. Estimating the quantile function of a location-scale family of distributions based on few selected order statistics. Journal of Statistical Planning and Inference, 8(1):75–86, September 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375883900630.

Srivastava:1981:EIM

[SB81] V. K. Srivastava and S. Bhatnagar. Estimation of the inverse of mean. Journal of Statistical Planning and Inference, 5(4):329– 334, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588190032X.

Srivastava:1986:SNM

[SB86] J. N. Srivastava and R. J. Beaver. On the superiority of the nested multi-dimensional block designs, relative to the classical incomplete block designs. Journal of Statistical Planning and Inference, 13(??):133–150, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588690128X.

Samuel-Cahn:1983:MES

[SC83] Ester Samuel-Cahn. Minimizing the expected sample range. Journal of Statistical Planning and Inference, 8(3):347–354, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375883900502.

Srivastava:1985:PMR

[SC85] V. K. Srivastava and R. Chandra. Properties of the mixed regression estimator when disturbances are not necessarily normal. Journal of Statistical Planning and Inference, 11(1):15– 21, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900217.

Schervish:1984:LDT

[Sch84] Mark J. Schervish. Linear discrimination for three known normal populations. Journal of Statistical Planning and Inference, 10(2):167–175, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900685.

Schick:1987:NCA

[Sch87a] Anton Schick. A note on the construction of asymptotically linear estimators. Journal of Statistical Planning and Inference, 16 (??):89–105, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900590. See correction

[Sch89a].

Schwabe:1987:ACB

[Sch87b] Rainer Schwabe. On adaptive chemical balance weighing designs. Journal of Statistical Planning and Inference, 17(??):209– 216, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887901133.

Schick:1989:CNC

[Sch89a] Anton Schick. Correction to “A note on the construction of asymptotically linear estimators” [J. Statist. Plann. Inference 16 (1987) 89–105]. Journal of Statistical Planning and Inference, 22(2): 269–270, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889901183. See [Sch87a].

Schucany:1989:NRH

[Sch89b] William R. Schucany. On nonparametric regression with higherorder kernels. Journal of Statistical Planning and Inference, 23 (2):145–151, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900852.

Sinha:1982:RPD

[SD82] Kishore Sinha and Aloke Dey. On resolvable PBIB designs. Journal of Statistical Planning and Inference, 6(2):179–181, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900878.

Sharma:1985:DTA

[SD85a] H. C. Sharma and S. N. Dube. Designs through automorphisms and solutions of T _{2} (45, 5; 0, 1). Journal of Statistical Planning and Inference, 12(??):115–122, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588590059X.

Shukla:1985:CIE

[SD85b] G. K. Shukla and P. Datta. Comparison of the inverse estimator with the classical estimator subject to a preliminary test in linear calibration. Journal of Statistical Planning and Inference, 12(??):93–102, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900576.

Sharma:1986:SPC

[SD86] Hoshyar Chand Sharma and S. N. Dube. Structure of PD-cycles and MOLS through them. Journal of Statistical Planning and Inference, 14(2–3):221–232, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588690159X.

Seberry:1980:CGH

[Seb80] Jennifer Seberry. A construction for generalized Hadamard matrices. Journal of Statistical Planning and Inference, 4(4):365– 368, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588090021X.

Sebo:1982:SSU

[Seb82] András Sebő. Sequential search using question-sets with bounded intersections. Journal of Statistical Planning and Inference, 7(2): 139–150, December 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900726.

Seberry:1984:RGD

[Seb84] Jennifer Seberry. Regular group divisible designs and Bhaskar Rao designs with block size three. Journal of Statistical Planning and Inference, 10(1):69–82, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900338.

Seberry:1985:GBR

[Seb85a] Jennifer Seberry. Generalized Bhaskar Rao designs of block size three. Journal of Statistical Planning and Inference, 11(3):373– 379, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900424.

Sebo:1985:TRS

[Seb85b] András Sebő. On two random search problems. Journal of Statistical Planning and Inference, 11(1):23–31, January 1985. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375885900229.

Sen:1984:CRM

[Sen84a] Pranab Kumar Sen. The Cox regression model, random censoring and locally optimal rank tests. Journal of Statistical Planning and Inference, 9(3):355–366, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900107.

Sen:1984:STA

[Sen84b] Pranab Kumar Sen. Subhypotheses testing against restricted alternatives for the Cox regression model. Journal of Statistical Planning and Inference, 10(1):31–42, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900302.

Schwager:1984:ECL

[SFM84] S. J. Schwager, W. T. Federer, and J. P. Mandeli. Embedding cyclic latin squares of order 2^{n} in a complete set of orthogonal F -squares. Journal of Statistical Planning and Inference, 10(2): 207–218, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900739.

Sinha:1984:FCD

[SG84] Bikas Kumar Sinha and Thomas M. Gerig. On a further characterization of dispersion matrices based on the properties of regression. Journal of Statistical Planning and Inference, 10(3):265– 271, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900521.

Shyr:1986:IAC

[SG86] J. Y. Shyr and L. J. Gleser. Inference about comparative precision in linear structural relationships. Journal of Statistical Planning and Inference, 14(2–3):339–358, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901722.

Shapiro:1986:LPL

[Sha86] Louis W. Shapiro. A lattice path lemma and an application in enzyme kinetics. Journal of Statistical Planning and Inference, 14(1): 115–122, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900157.

Shanmugam:1989:AHT

[Sha89] Ramalingam Shanmugam. Asymptotic homogeneity tests for mean exponential family distributions. Journal of Statistical Planning and Inference, 23(2):227–241, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900918.

Shirahata:1982:ERS

[Shi82] Shingo Shirahata. An extension of the ranked set sampling theory. Journal of Statistical Planning and Inference, 6(1):65– 72, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588290057X.

Shimizu:1989:ESM

[Shi89] R. Shimizu. Expansion of scale mixtures of the gamma distribution. Journal of Statistical Planning and Inference, 21(3):305– 314, March 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900499.

Shorack:1989:RTS

[Sho89] G. R. Shorack. Randomly trimmed L-statistics. Journal of Statistical Planning and Inference, 21(3):293–304, March 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375889900487.

Siddiqui:1982:BRS

[Sid82a] M. M. Siddiqui. Book review: Standardization; mathematical methods in assortment determination: C. Bongers. Journal of Statistical Planning and Inference, 6(3):305–307, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900350.

Siddiqui:1982:CMT

[Sid82b] M. M. Siddiqui. The consistency of a matching test. Journal of Statistical Planning and Inference, 6(3):227–233, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900271.

Simons:1980:SEC

[Sim80] Gordon Simons. Sequential estimators and the Cramér-Rao lower bound. Journal of Statistical Planning and Inference, 4(1):67– 74, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375880900348.

Singh:1981:SCN

[Sin81] R. S. Singh. Speed of convergence in nonparametric estimation of a multivariate μ-density and its mixed partial derivatives. Journal of Statistical Planning and Inference, 5(3):287– 298, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900094.

Sinha:1982:CCE

[Sin82a] Bikas Kumar Sinha. On complete classes of experiments for certain invariant problems of linear inference. Journal of Statistical Planning and Inference, 7(2):171–180, December 1982. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375882900751.

Sinha:1982:SAS

[Sin82b] Bikas Kumar Sinha. Some aspects of simplicity in the analysis of block designs. Journal of Statistical Planning and Inference, 6(2):165–172, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900854.

Singh:1985:EID

[Sin85] Randhir Singh. Estimation from incomplete data in longitudinal surveys. Journal of Statistical Planning and Inference, 11(2):163– 170, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900035.

Sinha:1987:BEP

[Sin87] S. K. Sinha. Bayesian estimation of the parameters and reliability function of a mixture of Weibull life distributions. Journal of Statistical Planning and Inference, 16(??):377–387, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375887900905.

Sen:1985:DRO

[SK85] Kanwar Sen and C. L. Kaul. Distributions of rank order statistics related to a generalised random walk. Journal of Statistical Planning and Inference, 12(??):177–191, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900679.

Sengupta:1989:EFP

[SK89] Samindranath Sengupta and Debasree Kundu. Estimation of finite population mean in randomized response surveys. Journal of Statistical Planning and Inference, 23(1):117–125, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837588990044X.

Skovlund:1988:SST

[Sko88] Eva Skovlund. A simulation study of three sequential two-sample tests with triangular continuation regions. Journal of Statistical Planning and Inference, 19(3):383–394, July 1988. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375888900456.

Skovlund:1989:SST

[Sko89] Eva Skovlund. A simulation study of three sequential two-sample tests with triangular continuation regions. Journal of Statistical Planning and Inference, 22(3):403–404, July 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375889901043.

Sinha:1982:NUO

[SM82] Bikas Kumar Sinha and Rahul Mukerjee. A note on the universal optimality criterion for full rank models. Journal of Statistical Planning and Inference, 7(1):97–100, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375882900222.

Stone:1985:ESN

[SM85] Mervyn Stone and Julie A. Morris. Exploratory studies in nonsequential non-local non-linear design. Journal of Statistical Planning and Inference, 12(??):1–9, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900485.

Sievers:1986:RRA

[SM86] Gerald L. Sievers and Joseph W. McKean. On the robust rank analysis of linear models with nonsymmetric error distributions. Journal of Statistical Planning and Inference, 13(??):215– 230, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901345.

Sen:1987:ORM

[SM87] Mausumi Sen and Rahul Mukerjee. Optimal repeated measurements designs under interaction. Journal of Statistical Planning and Inference, 17(??):81–91, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887901029.

Saha:1988:BNC

[SMK88] G. M. Saha, Rahul Mukerjee, and Sanpei Kageyama. Bounds on the number of constraints for balanced arrays of strength t. Journal of Statistical Planning and Inference, 18(2):255–265, February 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900109.

Sinha:1983:BAM

[SN83] K. Sinha and A. K. Nigam. Balanced arrays and main-effect plans from regular group-divisible designs. Journal of Statistical Planning and Inference, 8(2):223–229, October 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588390040X.

Singh:1989:MES

[SN89] Randhir Singh and P. Narain. Method of estimation from samples with random sample sizes. Journal of Statistical Planning and Inference, 23(2):217–225, October 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889900906.

Singh:1985:ACR

[SNS85] N. P. Singh, A. K. Nigam, and M. Singh. Analysis and characterization of row-column experiments involving two non-interacting sets of treatments. Journal of Statistical Planning and Inference, 12(??):193–198, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900680.

Shirakura:1985:SDF

[SO85] Teruhiro Shirakura and Toshio Ohnishi. Search designs for 2^{m} factorials derived from balanced arrays of strength 2(l+1) and ad-optimal search designs. Journal of Statistical Planning and Inference, 11 (2):247–258, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900126.

Sobel:1986:FPA

[Sob86] Marc Sobel. Further paradoxes in admissibility. Journal of Statistical Planning and Inference, 14(2–3):213–220, ???? 1986. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375886901588.

Sathe:1981:EDP

[SP81] Y. S. Sathe and M. Pradhan. On existence of dual property in symmetric group divisible designs. Journal of Statistical Planning and Inference, 5(4):377–380, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375881900380.

Seoh:1989:CLT

[SP89] Munsup Seoh and Madan L. Puri. Central limit theorems under alternatives for a broad class of nonparametric statistics. Journal of Statistical Planning and Inference, 22(3):271– 294, July 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900955.

Spruill:1985:MRH

[Spr85] Carl Spruill. Model robustness of Hoel–Levine optimal designs. Journal of Statistical Planning and Inference, 11(2):217– 225, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900084.

Spruill:1987:ODI

[Spr87] M. C. Spruill. Optimal designs for interpolation. Journal of Statistical Planning and Inference, 16(??):219–229, ???? 1987. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375887900711.

Srivastava:1982:CSE

[SR82] V. K. Srivastava and Ramkaran. A class of shrunken estimators for normal mean. Journal of Statistical Planning and Inference, 6(3):297–299, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900337.

Srivastava:1980:EN

[Sri80] Jaya Srivastava. Editorial note. Journal of Statistical Planning and Inference, 4(1):1, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375880900270.

Srivastava:1981:A

[Sri81] Jaya N. Srivastava. Acknowledgement. Journal of Statistical Planning and Inference, 5(4):415–416, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375881900446.

Srivastava:1984:API

[Sri84] Jaya Srivastava. Announcement and preface to this issue. Journal of Statistical Planning and Inference, 9(3):269, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375884900016.

Srivastava:1985:AS

[Sri85a] Jaya Srivastava. Acknowledgement of services. Journal of Statistical Planning and Inference, 11(1):xvii–xix, January 1985. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375885900199.

Srivastava:1985:JK

[Sri85b] Jaya Srivastava. Jack Kiefer (1924–1981). Journal of Statistical Planning and Inference, 11(1):xiv, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900175.

Srivastava:1985:OJ

[Sri85c] Jaya Srivastava. Organization of JSPI. Journal of Statistical Planning and Inference, 11(1):xv–xvi, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375885900187.

Srivastava:1986:FIA

[Sri86] Jaya Srivastava. Foreword to this issue and announcement of future special issues in various fields of this journal. Journal of Statistical Planning and Inference, 14(1):1, May 1986. CODEN JSPIDN.

ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886900029.

Srivastava:1987:FA

[Sri87a] Jaya Srivastava. Foreword and announcement. Journal of Statistical Planning and Inference, 16(??):iii, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375887900498.

Srivastava:1987:MEL

[Sri87b] Jaya Srivastava. More efficient and less time-consuming censoring designs for life testing. Journal of Statistical Planning and Inference, 16(??):389–413, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900917.

Srivastava:1988:F

[Sri88] Jaya Srivastava. Foreword. Journal of Statistical Planning and Inference, 20(3):239, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900900.

Srivastava:1989:A

[Sri89] Jaya Srivastava. Announcement. Journal of Statistical Planning and Inference, 21(1):1, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837588990013X.

Salama:1982:ESR

[SS82a] Ibrahim A. Salama and Pranab K. Sen. On expected sample range from heterogeneous symmetric distributions. Journal of Statistical Planning and Inference, 6(3):235–239, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900283.

So:1982:NRS

[SS82b] Y. C. So and P. K. Sen. Nonparametric repeated significance tests for one-way anova with adaptation to multiple comparisons. Journal of Statistical Planning and Inference, 7(1):83– 96, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375882900210.

Sen:1983:DCG

[SS83a] Kanwar Sen and Jagdish Saran. On the distribution of crossings in a generalized random walk. Journal of Statistical Planning and Inference, 7(4):371–385, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900083.

Shaked:1983:ISS

[SS83b] Moshe Shaked and Nozer D. Singpurwalla. Inference for step-stress accelerated life tests. Journal of Statistical Planning and Inference, 7(4):295–306, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900010.

Singh:1983:SPE

[SS83c] Randhir Singh and D. Singh. Sampling with partial enumeration from bivariate populations. Journal of Statistical Planning and Inference, 7(4):343–351, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375883900058.

Sinha:1983:OWD

[SS83d] Bikas Kumar Sinha and Rita Saha. Optimal weighing designs with a string property. Journal of Statistical Planning and Inference, 8 (3):365–374, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375883900526.

Sahai:1985:EUA

[SS85a] Ashok Sahai and Ajit Sahai. On efficient use of auxiliary information. Journal of Statistical Planning and Inference, 12(??): 203–212, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900709.

Sinha:1985:ETA

[SS85b] Bikas Kumar Sinha and Gour Mohan Saha. On the equivalence of two approaches to simplicity in the analysis of block designs and some related results. Journal of Statistical Planning and Inference, 11(2):237–240, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900102.

Street:1985:DPN

[SS85c] Deborah J. Street and Anne Penfold Street. Designs with partial neighbour balance. Journal of Statistical Planning and Inference, 12(??):47–59, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900527.

Sathe:1986:MOW

[SS86a] Y. S. Sathe and R. G. Shenoy. MV-optimal weighing designs. Journal of Statistical Planning and Inference, 13(??):31– 36, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901163.

Singer:1986:MMG

[SS86b] Julio M. Singer and Pranab K. Sen. M-methods in growth curve analysis. Journal of Statistical Planning and Inference, 13(??): 251–261, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901370.

Sane:1987:GQR

[SS87] Sharad S. Sane and S. S. Shrikhande. On generalised quasi-residual designs. Journal of Statistical Planning and Inference, 17(??): 269–276, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887901182.

Schick:1988:EES

[SS88a] A. Schick and V. Susarla. Efficient estimation in some missing data problems. Journal of Statistical Planning and Inference, 19 (2):217–228, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900742.

Sinha:1988:OAC

[SS88b] B. K. Sinha and K. R. Shah. Optimality aspects of 3-concurrence most balanced designs. Journal of Statistical Planning and Inference, 20(2):229–236, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888901279.

Sengupta:1987:SIA

[SSP87] Samindranath Sengupta, Bikas K. Sinha, and Sastry G. Pantula. Some inferential aspects of finite population sampling with additional resources. Journal of Statistical Planning and Inference, 16(??):203–211, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900693.

Stam:1988:SSP

[Sta88] A. J. Stam. Some stochastic processes in a random permutation. Journal of Statistical Planning and Inference, 19(2):229– 244, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900754.

Staiger:1989:CPH

[Sta89] Ludwig Staiger. Combinatorial properties of the Hausdorff dimension. Journal of Statistical Planning and Inference, 23(1):95– 100, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900426.

Stijnen:1982:EBR

[Sti82] Theo Stijnen. Empirical Bayes rules and Gaussian processes. Journal of Statistical Planning and Inference, 6(4):363–372, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900052.

Street:1984:SCP

[Str84] Deborah J. Street. Some constructions for PBIBDs. Journal of Statistical Planning and Inference, 10(1):119–129, July 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375884900387.

Stufken:1988:BEB

[Stu88] John Stufken. On bounds for the efficiency of block designs for comparing test treatments with a control. Journal of Statistical Planning and Inference, 19(3):361–372, July 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900432.

Soares:1985:ORA

[SW85] JoséF. Soares and C. F. J. Wu. Optimality of random allocation design for the control of accidental bias in sequential experiments. Journal of Statistical Planning and Inference, 11(1):81– 87, January 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900278.

Shohoji:1981:GFT

[SYT81] Takao Shohoji, Yuji Yamashita, and Tomoyuki Tarumi. Generalization of Fisher’s z-transformation. Journal of Statistical Planning and Inference, 5(4):347–354, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375881900355.

Tahir:1989:SOA

[Tah89] Mohamed Tahir. Second order asymptotic optimality in sequential point estimation. Journal of Statistical Planning and Inference, 23 (1):83–93, September 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900414.

Takacs:1986:SAF

[Tak86] Lajos Takács. Some asymptotic formulas for lattice paths. Journal of Statistical Planning and Inference, 14(1):123–142, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900169.

Tamari:1981:NCO

[Tam81] Fumikazu Tamari. A note on the construction of optimal linear codes. Journal of Statistical Planning and Inference, 5(4): 405–411, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900422.

Tan:1985:TRP

[Tan85] W. Y. Tan. On Tiku’s robust procedure — a Bayesian insight. Journal of Statistical Planning and Inference, 11(3):329– 340, April 1985. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375885900382.

Taneja:1987:SAD

[Tan87] Inder Jeet Taneja. Statistical aspects of divergence measures. Journal of Statistical Planning and Inference, 16(??):137–145, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375887900632.

Taylor:1989:NCE

[Tay89] Jeremy M. G. Taylor. A note on the cost of estimating the ratio of regression parameters after fitting a power transformation. Journal of Statistical Planning and Inference, 21(2):223– 230, February 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889900062.

Thompson:1984:MDC

[Tho84] M. E. Thompson. Model and design correspondence in finite population sampling. Journal of Statistical Planning and Inference, 10(3):323–334, October 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900570.

Tiku:1980:RME

[Tik80] M. L. Tiku. Robustness of MML estimators based on censored samples and robust test statistics. Journal of Statistical Planning and Inference, 4(2):123–143, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375880900026.

Thompson:1987:IDC

[TL87] W. A. Thompson and William R. Lower. Indirect determination of correlation. Journal of Statistical Planning and Inference, 16 (??):83–87, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887900589.

Tonchev:1981:BDA

[Ton81] V. D. Tonchev. On block designs arising from rank 3 graphs. Journal of Statistical Planning and Inference, 5(4):399–403, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375881900410.

Torgersen:1984:OLM

[Tor84] Erik Torgersen. Orderings of linear models. Journal of Statistical Planning and Inference, 9(1):1–17, January 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900405.

Tosic:1980:OSP

[Tos80] Ratko Tosić. An optimal search procedure. Journal of Statistical Planning and Inference, 4(2):169–171, ???? 1980. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375880900051.

Tosic:1989:FCC

[Tos89] Ratko Tosić. Five counterfeit coins. Journal of Statistical Planning and Inference, 22(2):197–202, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889901109.

Troya:1982:CDC

[Tro82a] Josefa Lopes Troya. Cyclic designs for a covariate model. Journal of Statistical Planning and Inference, 7(1):49–75, October 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900192.

Troya:1982:ODC

[Tro82b] Josefa Lopes Troya. Optimal designs for covariates’ models. Journal of Statistical Planning and Inference, 6(4):373–419, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375882900064.

Tse:1986:EUM

[Tse86] Siu-Keung Tse. On the existence and uniqueness of maximum likelihood estimates in polytomous response models. Journal of Statistical Planning and Inference, 14(2–3):269–273, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886901655.

Tuza:1982:WCB

[Tuz82] Zsolt Tuza. Worst-case behavior of string-searching algorithms. Journal of Statistical Planning and Inference, 6(1):99–103, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837588290060X.

Tse:1986:CDP

[TW86] Siu-Keung Tse and C. F. Jeff Wu. Comparison of dichotomous and polytomous response models. Journal of Statistical Planning and Inference, 14(2–3):187–202, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375886901564.

Vartak:1989:USA

[VD89] M. N. Vartak and S. M. Diwanji. Uniqueness of some 4-associate schemes. Journal of Statistical Planning and Inference, 22(3): 397–401, July 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889901031.

Vecchi:1984:SRC

[Vec84] Italo Vecchi. Some results on coverings of Galois spaces with ovoids and related bib designs. Journal of Statistical Planning and Inference, 10(2):219–225, August 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900740.

Venezia:1981:CSS

[Ven81] Itzhak Venezia. The covariance structure of sequential forecasts obtained by regression analysis. Journal of Statistical Planning and Inference, 5(2):121–132, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375881900227.

Visek:1986:STE

[Vís86] Jan Ámos Vísek. Sensitivity of the test error probabilities with respect to the level of contamination in general model of contaminacy. Journal of Statistical Planning and Inference, 14(2–3): 281–299, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901679.

Vardeman:1983:AEF

[VM83] Stephen Vardeman and Glen Meeden. Admissible estimators in finite population sampling employing various types of prior information. Journal of Statistical Planning and Inference, 7(4): 329–341, June 1983. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375883900046.

Wang:1987:EDF

[Wan87] Jane-Ling Wang. Estimators of a distribution function with increasing failure rate average. Journal of Statistical Planning and Inference, 16(??):415–427, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375887900929.

Watson:1982:BRP

[Wat82] G. S. Watson. Book review: Practical experiences with modelling and forecasting time series: E. M. Jenkins (Time Series Library, a GJP publication, 1979) 146 pp. Journal of Statistical Planning and Inference, 6(2):203, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900908.

Watson:1983:LST

[Wat83] Geoffrey S. Watson. Large sample theory of the Langevin distribution. Journal of Statistical Planning and Inference, 8(3):245– 256, December 1983. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375883900435.

Watanabe:1986:GPB

[Wat86] Toshihiro Watanabe. On a generalization of polynomials in the ballot problem. Journal of Statistical Planning and Inference, 14(1): 143–152, May 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900170.

Weier:1980:IKM

[WB80] D. R. Weier and A. P. Basu. An investigation of Kendall’s τ modified for censored data with applications. Journal of Statistical Planning and Inference, 4(4):381–390, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375880900233.

Wegman:1984:ONF

[Weg84] Edward J. Wegman. Optimal nonparametric function estimation. Journal of Statistical Planning and Inference, 9(3):375– 387, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375884900120.

Wegman:1988:RDC

[Weg88] Edward J. Wegman. On randomness, determinism and computability. Journal of Statistical Planning and Inference, 20(3): 279–294, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888900936.

Wehrhahn:1986:PSD

[Weh86] K. Wehrhahn. Properties of some distributions in randomization theory. Journal of Statistical Planning and Inference, 13(??):231– 237, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901357.

Whitney:1987:CLS

[Whi87] Paul Whitney. Consistent least squares nonparametric regression. Journal of Statistical Planning and Inference, 17(??):137– 148, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887901078.

Whittinghill:1989:BBD

[Whi89] Dexter C. Whittinghill. Balanced block designs robust against the loss of a single observation. Journal of Statistical Planning and Inference, 22(1):71–80, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900669.

Wierich:1985:ODU

[Wie85] W. Wierich. Optimum designs under experimental constraints for a covariate model and an intra-class regression model. Journal of Statistical Planning and Inference, 12(??):27–40, ???? 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375885900503.

Wiens:1988:NND

[Wie88a] Douglas Wiens. The non-null distribution of the beta statistic in the test of the univariate general linear hypothesis, when the error distribution is spherically symmetric. Journal of Statistical Planning and Inference, 19(2):245–252, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375888900766.

Wierich:1988:ODM

[Wie88b] W. Wierich. A-optimal design measures for one-way layouts with additive regression. Journal of Statistical Planning and Inference, 18(1):57–68, January 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588890016X.

Wierich:1989:AKW

[Wie89] W. Wierich. An application of the Kiefer–Wolfowitz equivalence theorem to ANOVA models with additive regression. Journal of Statistical Planning and Inference, 21(3):277–283, March 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900463.

Wijesinha:1987:COD

[WK87] M. Cooray Wijesinha and AndréI. Khuri. Construction of optimal designs to increase the power of the multiresponse lack of fit test. Journal of Statistical Planning and Inference, 16(??): 179–192, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588790067X.

Wehrhahn:1982:RCB

[WO82] K. H. Wehrhahn and J. Ogawa. Randomization in a complete block design with few blocks. Journal of Statistical Planning and Inference, 6(1):33–45, ???? 1982. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375882900544.

Wood:1984:TNE

[Woo84] Constance L. Wood. On tests for normality of experimental error in ridge regression. Journal of Statistical Planning and Inference, 9(3):367–374, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375884900119.

Weideman:1987:SON

[WR87] C. A. Weideman and D. Raghavarao. Some optimum non-adaptive hypergeometric group testing designs for identifying two defectives. Journal of Statistical Planning and Inference, 16(??):55– 61, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887900553.

Wolfe:1984:NSP

[WS84] Douglas A. Wolfe and Edna Schechtman. Nonparametric statistical procedures for the changepoint problem. Journal of Statistical Planning and Inference, 9(3):389–396, May 1984. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375884900132.

Wegman:1988:IAU

[WS88] Edward J. Wegman and Herbert Solomon. Introduction to assessing uncertainty. Journal of Statistical Planning and Inference, 20(3):241–243, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375888900912.

Wu:1987:RWC

[Wu87] Tiee-Jian Wu. Rates of weak convergence and large deviation probabilities for linear rank tests with type I censored data. Journal of Statistical Planning and Inference, 17(??):67– 79, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887901017.

Wu:1989:NCL

[Wu89] Tiee-Jian Wu. L _{1} nonuniform central limit bounds for generalized rank statistics. Journal of Statistical Planning and Inference, 22(2):243–250, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889901146.

Yanagisawa:1988:DDB

[Yan88] Yukio Yanagisawa. Designs for discrimination between binary response models. Journal of Statistical Planning and Inference, 19 (1):31–41, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837588890050X.

Yanagawa:1980:MPR

[YC80] T. Yanagawa and Shan-Huo Chen. The mg-procedure in ranked set sampling. Journal of Statistical Planning and Inference, 4(1): 33–44, ???? 1980. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375880900312.

Yeh:1988:CUO

[Yeh88] Ching-Ming Yeh. A class of universally optimal binary block designs. Journal of Statistical Planning and Inference, 18(3):355– 361, March 1988. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375888901115.

Ylvisaker:1981:BRE

[Ylv81] D. Ylvisaker. Book review: Exploratory and explanatory statistical analysis of spatial data: Cornelis P. A. Bartels and Ronald H. Ketellapper, editors (Martinus Nijhoff Publishing, Boston, 1979). Journal of Statistical Planning and Inference, 5(4):413, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375881900434.

Yamato:1989:EUI

[YM89] Hajime Yamato and Yoshihiko Maesono. Estimation under invariant distributions. Journal of Statistical Planning and Inference, 22(1):55–61, May 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900645.

Young:1987:QDS

[YMO87] Dean M. Young, Virgil R. Marco, and Patrick L. Odell. Quadratic discrimination: Some results on optimal low-dimensional representation. Journal of Statistical Planning and Inference, 17(??): 307–319, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375887901224.

Suen:1989:COM

[yS89a] Chung yi Suen. A class of orthogonal main-effect plans. Journal of Statistical Planning and Inference, 21(3):391–394, March 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375889900554.

Suen:1989:SRD

[yS89b] Chung yi Suen. Some rectangular designs constructed by the method of differences. Journal of Statistical Planning and Inference, 21(2):273–276, February 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375889900116.

Zabell:1988:PAT

[Zab88] S. L. Zabell. The probabilistic analysis of testimony. Journal of Statistical Planning and Inference, 20(3):327–354, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900961.

Zahedi:1985:SNC

[Zah85] Hassan Zahedi. Some new classes of multivariate survival distribution functions. Journal of Statistical Planning and Inference, 11 (2):171–188, February 1985. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375885900047.

Zacks:1981:BPD

[ZB81] S. Zacks and Z. Barzily. Bayes procedures for detecting a shift in the probability of success in a series of Bernoulli trials. Journal of Statistical Planning and Inference, 5(2):107– 119, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375881900215.

Zijlstra:1981:SOR

[ZdK81] M. Zijlstra and J. P. M. de Kroon. Stochastic orderings related to ranking in an n th order record process. Journal of Statistical Planning and Inference, 5(1):79–91, ???? 1981. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375881900537.

Zelterman:1988:RET

[Zel88] Daniel Zelterman. Robust estimation in truncated discrete distributions with application to capture-recapture experiments. Journal of Statistical Planning and Inference, 18(2):225–237, February 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375888900079.

Zidek:1988:GDA

[Zid88] James V. Zidek. Group decision analysis and its application to combining opinions. Journal of Statistical Planning and Inference, 20(3):307–325, ???? 1988. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837588890095X.